A Complete Bibliography of ACM Transactions on Modeling and Computer Simulation

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

16 May 2024 Version 1.90

Title word cross-reference

1, 2, 3 [SMDS11]. 3 [JBH^{+}22, Pac08]. b [Joh96]. m [MK96, Mat98]. O(1) [TGT05]. q [GDB14].

-Gaussian [GDB14]. -sequence [Mat98]. -sequences [MK96].

1 [JB22a].

2 [JB22b]. 2005 [AGG^{+}07]. 2011 [LCK11]. 2017 [BB19]. 2019 [LK21, PW21]. 2020 [GC22]. 2021 [AM23, BNSS24, DT22].

623-dimensionally [MN98]. 64-bit [Nis00].

Abstraction [Lor19, MHS19, LW97a]. Accelerated [MJV^{+}15, HD07, SLCP01]. Accelerating [And99]. acceleration [PF11]. Accelerators [RAGN19]. acceptance [Bel05]. acceptance-rejection [Bel05]. Access [CTF^{+}19, AZLT10, KHJ^{+}08]. access/modification [Mat05]. accessibility [YJ96]. accreditation [PCT97]. Accurate [CMM^{+}16, KPG15]. Achieving [LBL01]. Active [LW97a, WG04]. active-idle [WG04]. activities [DOD93]. activity [CKL^{+}13]. Actor [PBB16]. Adaptation [HERU15, PBB16]. Adaption [Di 23, WWH^{+}23]. Adaptive [Ald18, Bee18, Bha05, Bha07, BBCD22, BCZ14, DHK15, DF97, ESZH21, FHG16, JV23, LCT^{+}15, Lüc16, SFM13, SK23, TL18, VAB^{+}18, WYT^{+}20, HD07, Kaw10, MKPR98, MY08]. add [TLC93]. add-with-carry [TLC93]. address [DJS94]. admission [Lim12]. Adomian [Tur17]. Advanced [Ano18, JW19, MST17, QTP20, Wai15, PCT97]. Advancement [BN22]. Advances [BSV16]. Adversarial [FBCS22]. AES [HW03]. aesthetic [FDD05]. affect [FA06]. Agenda [RSG21]. Agent [And22, KH19, LCT17, LCL16, Mar22, PE11, RWU22, XCA^{+}17, EK04, LLT07, NCV06, RD10]. Agent-Based [And22, KH19, LCT17, Mar22, PE11, XCA^{+}17, RWU22, LLT07, RD10]. Agents [NB93]. aggregation [KK00]. ahead [MWMD07]. AI [MFFR92]. AIR [WAGP15]. aircraft [RFA00]. Airlift [PBAB^{+}11, WPW09]. Alarms [BDK^{+}19]. algebraic [PB96]. Algorithm [BBMK16, LF13, TDR^{+}11, WDYR16, WYT^{+}20, CO98, EK04, EK07, KCK08, Kra96, KT10, LL91b, PTCL11, RTY05, SG91, XNH10]. algorithmic [BKM09]. Algorithms [CTC^{+}05, GDB14, HERU15, HWMU17, Hil17, LT14, PPT14, PBB16, Sch13, SMI15, SSZ^{+}13, Bha05, Bha07, BHM11, BCZ14, HN07, MWM91, NH95, Nut08, PS09, RR93, RA97]. Allocation [CWGZ24, MRB^{+}18, YX17, HLC^{+}10, Kaw10, LBL01, ZK10, ZG94]. Allocations [FH18]. Alternative [KW15, CTLZ05, Owe03]. Alternatives [ÇTI13]. among [WM99]. analogue [Tez93, TT94]. Analysis [BN09, BBCD22, CHA^{+}22, De 06, DNRD96, GK95, GH91, JBH^{+}22, KV23, Kra96, LCK11, NY12, PH21, SDZ^{+}15, Van19, VLN^{+}19, XNB16, XLZ17, ZC18, ZL17, ZH19, AQVA10, BL02, BCL^{+}97, BG93, Buc98, CGN06, Cal07, Cal09, DJS94, GBA^{+}14, KSZ11, LSW91, MR02, PF11, RRW00, Sch10, SLW^{+}05, TFR07, Vor10, WG04, WCLG10]. Analytics [GB19]. analyzers [Lin92].

Annealing [HZF14]. annotations [DKVR09]. antithetic [YL96]. any [HLC12]. Application [DTCU19, RSG21, SPYG24, WAGP15, YP18, CKL^{+}13, HT99, PW95, PRO13]. Application-Level [WAGP15]. Applications [BMLY19, CFL12, Bal01, BCZ14, EGLW93, KT10, Lim12, MWMD07, PTCL11]. applied [Tuf97]. Applying [Nak14, CN12]. appreciation [AGG^{+}07]. Approach [And21, BHH21, CM21, KCS20, SALS18, Van18, WCS16, WCCY19, YX17, ZZC18, BHG10, BÖ96, BKV04, BB94, Buc98, FHD09, Fis92, HD07, LL02, MPK06, MMRC^{+}08, QFL^{+}10, RAF^{+}04, SV97, TTSM12, Uhr01]. Approaches [MKT21, SSY21]. Approximate [CERT15, JKE14, Ros08]. Approximation [CPRV23, FHG16, HSS24, JYE24, LF13, Lüc16, PH21, ZS17, BFMW03, Bha05, BHM11, BCZ14, KT10, LCT07, Lim12, PS09]. Approximation-based [PH21]. approximations [FK91, MR02]. Architectural [KCS20]. Architecture [CAT22, CHIW98, SQ12, SB01]. Architectures [PCGM18, CG02, TAO08]. Area [JBH^{+}22]. arithmetic [Tez93, TT94]. Arrival [MJ15, WCF23, WPN98]. Article [Pic24]. artificial [Fis92]. ASAP3 [SLW^{+}05]. ascent [MSK10]. Assessment [LB15]. assignment [VaAE02]. Assimilation [HW19, XGH12]. assisted [XYZ21]. astrophysical [Pac08]. Asymptotic [GJ13, LBTG10]. Asymptotically [Kaw10, RW93]. Asymptotics [JKS07]. Asynchronous [CB24, MWM91, BP94, EK04]. ATM [KW93, LC01, SLCP01, UXC^{+}00]. attacks [CFS08]. audio [ABGR01]. Automata [BSV16, BHH21, DWYM16, GB19, DSR23, TDR^{+}11, TKS16, Mat98]. Automated [AGMW17, CH23, RDSJ18, SSDW18, ZIC06, MV02]. Automatic [FHG16, GLC17, HERU15, LZ20, Lüc16, Di 23, WPN98, WWH^{+}23]. Automating [Yau99]. autonomic [MY08]. Autonomous [BLG^{+}21]. autoregressive [BN03]. available [HD07]. averages [FA06, KSW07]. Averaging [HZF14]. Avoiding [HWdF13]. Aware [LZ20, JV23].

Background [LL15, NY04]. Bad [Ent98]. Balanced [CERT15]. Balancing [WYT^{+}20]. Bandwidth [MRB^{+}18, FMN00]. bandwidths [FDL99]. base [ZLK91]. Based [And22, CDS16, CG13, FDP15, GJ13, HYJ21, HERU15, HZF14, JN15, KH19, KW15, LF13, LCT^{+}15, LL15, LCT17, Mar22, MJV^{+}15, PE11, RL15, SMI15, SP11, SU16, WCZ16, WCCY19, XLZ17, XCA^{+}17, ZS17, ZL17, AZK10, Bel05, Bha05, BÖ96, CAT22, CTF^{+}19, CTC^{+}05, KCS20, KLF02, LS92, LLT07, LCT07, LL02, LSW91, MK96, PBF^{+}00, PTCL11, PF11, PH21, RTY05, RS94, RRW00, RWU22, RD10, TTSM12, TB98, Vor10, WXC^{+}23, XZY23, ZMM^{+}11, vBBR03, Bha07, RFA00]. Batch [SPYG24, AG04, AAAG06, SLW^{+}05]. Batches [LB15]. Batching [SK23]. Bayesian [AG16, GK19, NY12, SCW13, UPB22, UB24, WCS16, XLZ17, YN15, ZS17]. BDI [LSJ10]. Behavior [GH03, GGH^{+}23, LZ20, ZZC18, CFW99, CKP95, GH06, GH09]. behavioral [MMRC^{+}08, YS92]. behaviors [KZ11]. Behavioural [OT24]. benchmark [JC11]. Bernoulli [KO94]. Bernstein [GS12]. best [ICC99, NS06, Oso09, YN93]. Better [Owe13]. between [BHG10, ZIC06]. Bézier [WW95]. BGP [CK08]. Bi [FH18]. Bi-objective [FH18]. Bias [BCM18, YKA^{+}21, AG07, Cal09, HIG04]. Bias-corrected [YKA^{+}21]. bias-reducing [HIG04]. biased [CK14]. biasing [Nak94]. Bifurcation [ACL15]. Binary [Sch13]. Binomial [FFSF13]. Biochemical [RL15]. Biological [DWYM16]. Birth [BK20].

Birth-and-Death [BK20]. Bisection [RL20]. bit [Nis00]. BitTorrent [LPPP13]. BitTorrent-like [LPPP13]. bivariate [Ros08, WW95]. black [FHD09]. blending [QFL^{+}10]. Block [LF13]. blocking [AO95, KC10, RRP00, SW96, VaAE02]. boiler [LS92]. Bootstrap [CLL99]. Border [CK08]. borrow [TLC93]. boundary [LM94]. Bounded [HSN94, DNRD96, JZTB06]. Bounds [FK91, Nic91]. Box [LLCC13, FHD09]. Bridging [TTSM12]. broadband [GMOB01]. Brownian [BCM18, IFPM12]. Bryant [SG91]. Budget [CWGZ24, MH19, YX17, HLC^{+}10]. buffer [CHS95, HHY11, JN05, KM01]. buffers [KW93]. built [Mat98]. built-in [Mat98]. bulk [HVA09]. burst [WG04]. bursty [GMOB01]. Business [BDGP20, RD10].

Cache [ANSW23, TKS16, JSC01].

Calculation [CH04]. calendar [ELL00]. Calibration [WXC^{+}23, YN15, YN20]. cancer [RWK^{+}07, TRK^{+}09]. Capabilities [CN16]. care [MBGF11]. Carlo [DR13, Pel21, CB24, DJLZ17, FS21, FSS95, HHL14b, LDT07, LV00, LG03, XGH12]. Carma [Lor18, GZWG18]. carry [GK03, TLC93]. carrying [GMOB01]. case [CFS08, PCT97, SY95]. CDIF [Fla02]. CDNsim [SPV^{+}10]. cell [LC01]. cell-loss [LC01]. Cellular [BSV16, GB19, TDR^{+}11, TKS16, FSS95, Mat98, VaAE02]. Center [WXC^{+}23, GBA^{+}14]. Centers [LHJS17, HAFDP11]. central [SS05]. certification [Bal01]. CFTP [DJ11]. Chain [MKT21, RK20, AHO93]. Chains [BDK^{+}19, Buc98, HPA07, NH95, RJ04]. Challenges [Fuj16]. Chandy [SG91]. Chandy-Misra-Bryant [SG91]. changing [RR93]. channel [VaAE02]. chaos [SS08]. Characteristic [CFL12]. Characteristics [RK20]. characterization [Nak94]. Characterizing [PRO13]. Checking [AP18, JSD19, LLCC13]. checkpoint [PLM94]. Chief [Qua20]. Chip [CMM^{+}16, CG02]. chip-multiprocessor [CG02]. ChunkedTejas [KCS20]. Chunking [KCS20]. Chunking-based [KCS20]. Circuit [GLC17, EGLW93, SS08]. circuit-switched [EGLW93]. Class [DQZ18, MZ91, DSR23, GS12, HVA09, Vak92]. classes [LPPP13]. classical [BN09]. Cloning [HF01, LCT17, YP18, CTC^{+}05]. closed [CS08, CO98, SMG09]. Closure [FHG16, Lüc16]. Cloud [CAT22, YP15, VSCL13]. Cloud-based [CAT22]. Cloud/Virtual [YP15]. Clouds [SALS18, Van18]. Cluster [LL15]. Cluster-Based [LL15]. Co [TFR07]. Co-Plot [TFR07]. Codes [CSRE21]. Coefficients [DC22]. Coevolution [FDMS16]. Collaborative [SALS18, Van18]. Collective [Ald18, Bee18, FHG16, Lüc16, TL18, VAB^{+}18]. collisions [PP13]. colorectal [RWK^{+}07, TRK^{+}09]. combination [HLC12]. Combined [WYT^{+}20, PN03, TL91]. combines [MBGF11]. Combining [HYJ^{+}18, JB22b, RJ04, YL96, Buc98]. Common [GK19, MWKA07, TKS16, Joh96, Nel93, CAN12]. Communication [KPG15, ZL17, AO95, DG10, LM94]. communications [CHS95]. Community [FDP15]. Community-Based [FDP15]. comparative [FL09, NH95, RA97]. Comparing [ABGR01]. Comparison [Kim05, AG07, DN99, DNRD96, SJY03, TRK^{+}09]. comparisons [HE12, Nel93, YN93]. COMPASS [XNH10]. completion [GH91]. Complex [CPRV23, CDS16, HWdF13, SSZ^{+}13, DKVR09, LDL04, RBDH97, SS14]. Component [HERU15, LL02]. Component-Based [HERU15, LL02]. composite [SS05]. Compositional [CPRV23]. compound [BL11, Lev01]. comprehensive [XNH10]. compression [MM07]. Computation [GLC17, CPF99, MH92]. Computational [And21, Ano21, Lüc16, Par18, Pel21, Qua19, Di 23, WJ22, YJ96]. Computations [Ale17, Bee18, Hil17, KH18, Lor18, Lor19, Nel17, Van18]. Computer [YN20, CHIW98, FW97, HD98, MV02, RBDH97]. Computing [AG16, CWGZ24, DHK15, FH97, KV23, LCK11, LG03, RRP00, ZC18, BCD^{+}14, FDD05, HLC^{+}10, KFL00]. concave [Ley98]. Conceptual [GDP14]. concerning [HW03]. ConceVE [GDP14]. Conditional [HHL14b, YKA^{+}21, LG03]. conditioning [LG03]. conditions [PT00]. Conference [LCK11]. Confidence [CN12, EH21, FG99, Nak14, Sin14, SPYG24, CH04, CLL99]. congestion [SJSM10]. congruential [EHG92, EHN94b, Ent98, LW97b]. conjoint [HD98]. Conjugacy [ZS17]. Conservation [BBCD22, HAFDP11]. Conservative [JB22b, BP94]. Consistency [RNS97, ZCLT04]. constant [RB08]. Constants [DJLZ17]. Constrained [FDP15, HKP21, PHP^{+}15, UB24, BHM11, MSK10, SF10]. constraint [GH91]. Constraints [HAK14, BK10]. Constructed [SPYG24]. Constructing [HLC12, Nut08]. Construction [HPS^{+}21, DSR23]. Contact [WCL^{+}19]. Contagion [XZY23]. container [ZIC06]. containment [HN09]. content [SPV^{+}10]. Contextual [CWGZ24]. Continuity [ÇVS15]. Continuous [BDK^{+}19, HL03, Buc98, LX14, NH95]. Continuous-time [BDK^{+}19]. Control [NS06, XZY23, AHO93, CKP95, DF97, Lim12, RJ04, SJSM10, YL96]. Controlling [BCM18]. Convergence [LF13, SFM13, Tur17, And99, And06]. convergent [HN07]. Conversion [Doo07, SQ12]. Copula [BLST16]. copulas [HE12]. Cores [PPT14]. corrected [YKA^{+}21]. Correlated [CMZ18, HAK14, GH03, GH06, GH09]. correlation [LCT07, Ros08]. correlations [WM99]. corresponding [QFL^{+}10]. Corrigendum [GH06, GH09]. Cost [PBAB^{+}11, FW97, MKPR98, TRK^{+}09]. cost-effectiveness [TRK^{+}09]. cost/performance [FW97]. Could [EH18, KH18]. countably [And06]. Coupled [KSL^{+}16]. Couplings [SU16]. Covariance [JFST24]. creating [NCV06]. Credit [ANSW23]. Critic [PBB16]. Cross [AZK10, Rub02, SF10, WZ15, DG10, HLC^{+}10, WCLG10]. Cross-Entropy [WZ15, Rub02, HLC^{+}10]. Cross-layer [AZK10, SF10, DG10, WCLG10]. Crowd [LZ20, LCL16, XZY23, ZZC18, ZLH^{+}22, ZCC^{+}10]. crowded [KZ11]. Crowds [HW21]. cryo [HAFDP11]. cryo-conservation [HAFDP11]. CTMC [BK20]. CUDA [SM12]. Cumulative [DHK15]. CURAND [SM12]. Curves [HHH^{+}19]. customization [RD10]. cut [Rub02]. Cyber [Ano21, BDH21, HYJ21]. Cyber-Physical [Ano21, HYJ21, BDH21]. Cycle [CMM^{+}16, CKL^{+}13, DX03]. Cycle-Accurate [CMM^{+}16].

D [JBH^{+}22, Pac08]. DAE [vBBR03]. DAE-based [vBBR03]. Data [BMLY19, CTF^{+}19, EH18, FBS20, HT20, HSS24, HW19, KH19, KW15, KH18, LL20, LHJS17, MD20, NCN^{+}22, San20, SS14, WXC^{+}23, XGH12, ZZC18, ZLH^{+}22, ZLZ23, BCD^{+}14, DOD93, FLV01, GBA^{+}14, HBE95, Mat05]. Data-Driven [KH19, ZZC18, CTF^{+}19, NCN^{+}22, SS14, ZLH^{+}22]. data-intensive [BCD^{+}14]. Database [FS21, Pel21, SSH97]. DDM [PTCL11, RTY05]. Death [BK20]. Debugging [GRK^{+}15, VVB^{+}20]. Decision [HHH^{+}19, LJS22, PTE^{+}11, SCW13, Kiv91, LSJ10, MY08]. Decision-Making [LJS22]. Decisions [PBAB^{+}11]. Declaration [vBBR03]. Decomposition [Tur17, AD92]. decoupling [FDL99]. Deep [GGH^{+}23]. defects [MWKA07]. defense [Pag93].

Defined [JN15]. Delay [CMZ18, FLV01]. delayed [JS02]. Deletion [WG16]. Demand [WCF23, ZK10]. denial [CFS08]. denotational [TB98]. densities [Dev97, HLD07]. Density [CPRV23, YKA^{+}21, ZZC18, DHL10, HLD07]. departments [ZMM^{+}11]. dependability [HD98]. dependable [HSN94]. Dependencies [BV22, WJ22]. Dependent [ZZC18, GMOB01, MSM10]. Deployment [ÇTI13]. Depth [JBH^{+}22]. Derivative [LN18]. Derivatives [BG93]. Deriving [ÇTI13, NNB11]. Design [Ald18, Bee18, FG98, NY12, RL15, AZK10, CHIW98, DHM93, GBA^{+}14, RRW00, RFA00, SB01, WCLG10]. Design-time [FG98]. designs [SS05]. Detection [BK20, CTF^{+}19, PTE^{+}11, AGT92, EK04, EK07, RB08]. Determination [SMI15, SSY21]. Deterministic [RB08, BFMW03]. Development [ÇVS15, RWK^{+}07]. Deviation [WCZ16, SM12]. Deviations [GJ13, MR02, MK96]. Device [KKTM17]. Devices [PTE^{+}11, CF11]. DEVS [CHA^{+}22, SU16]. DEVStone [CHA^{+}22]. DGHPSim [GP11]. diagram [CKL^{+}13]. diaphony [HN98]. difference [RJ04]. different [Ros08, Vak92]. Differentiable [And22]. Differential [HSL^{+}19, PB96]. differential-algebraic [PB96]. Differentiation [RLDH16, HVAPFY10]. Diffserv [LBL01]. Diffusion [RMWLP21]. Diffusions [DC22, LTM^{+}17]. Digital [EHN94a, WXC^{+}23, ZBTT24, Owe03, SG91]. dimension [GH03, GH06, GH09]. Dimensional [SNS16, DX03, Owe98]. Dimensionally [LZ20, MN98]. direct [HT99]. Disaster [LJS22]. discarding [WM99]. Discontinuous [DC22]. Discovery [FBS20]. discrepancy [BFN92, Hic96, RGTL12]. Discrete [Ano18, BBMK16, ÇVS15, HSL^{+}19, HPA07, HW19, JB22a, JW19, Mar22, MJV^{+}15, MST17, NY04, PPT14, PCGM18, PTD^{+}20, QTP20, RAGN19, RWU22, RMWLP21, SP11, SJY03, VWD22, VXE^{+}22, Wai15, WYT^{+}20, WMC^{+}18, WZCJ22, YP15, And99, BKV04, GLM96, HVAPFY10, HG01, HN07, HD96, Lim12, Lin92, MBGF11, MCC11, NOP99, Nic91, Nut06, Nut08, Pag93, RS94, RR93, TGT05, Vak92, YJ96, LG03]. Discrete-Event [Mar22, PCGM18, RMWLP21, VWD22, WMC^{+}18, WZCJ22, RWU22, SJY03, VXE^{+}22, HVAPFY10, HG01, MBGF11, MCC11, Nic91, Pag93]. Discrete-time [HPA07]. Discretization [BCM18]. Disease [PE11]. Dispersion [ACL15]. Displacement [SMI15]. Distributed [ÇTI13, CAT22, DTCU19, FHG16, Fuj16, HYJ21, LLT07, LT14, Lüc16, MKT21, PE11, Pic24, PTD^{+}20, WKC^{+}24, BCL91, BCL^{+}97, CTLZ05, CTC^{+}05, FK91, FG98, LLHL00, MH92, MWMD07, PCT97, RAF^{+}04, SSH97, SJSM10, SKR97, SB01, ST13, TTSM12, Vak92, ZCLT04]. Distribution [MD20, WZCJ22, LG03, SPV^{+}10]. Distributionally [LL20]. Distributions [DQZ18, GDB14, Hof11, MJ15, QDZ21, WCF23, Dev09, FA06, HD96, Ley98, RR93, SZ99, WW95]. Divergence [CB24]. division [LL91b]. domain [EU14]. domain-specific [EU14]. Donald [GH15a]. Doping [Ano21, BDH21]. double [DJ11]. down [CK08]. downlink [AZLT10]. Drawing [Gou22]. Driven [ÇVS15, JKE14, KCS20, KH19, ZZC18, CTF^{+}19, CSK10, DJS94, MWM91, MH92, NCN^{+}22, SS14, UNMS97, ZLH^{+}22]. Driving [BLG^{+}21, OLAM08]. Drug [XVN14]. DSMC [GGH^{+}23]. duration [NNB11]. dust [CFW99]. Dynamic [BBMK16, Bar03, CKM23, NCN^{+}22, Uhr01, VVB^{+}20, Bar97, FSS95, PTCL11, QFL^{+}10, VaAE02]. Dynamical [FDMS16, GHS18, Par18, BB94, MWM91]. Dynamical-Related [FDMS16].

Dynamics [HWdF13, HW21, MJV^{+}15, PH21, MMRC^{+}08].

easy [SMDS11]. Ecosystem [HT20].

Editor [BSV16, GH15a, CY10, CL98, DG10, HHL14a, TR08, Qua20, Wil07]. Editor-in-Chief [Qua20]. Editorial [Ano18, BSV16, Hei97, JW19, MST17, Nic97, Nic04, Qua20, QTP20, TL18, Wai15, FN03, MV02, Bal97]. Effect [PBAB^{+}11, RLDH16, LM94]. effective [FDL99]. effectiveness [TRK^{+}09]. Effects [ACL15, PLM94, CAN12]. Efficiency [GJ13, VAVA06, And06]. Efficient [AK18, BBMK16, BL11, BGL12, CPF99, Den05, FSS95, FFSF13, GK03, HWMU17, Hil17, JN05, KN02, LJS22, LBEJ19, LX14, MDH^{+}23, Nic08, NZ07, RDSJ18, SW13, TL91, VXE^{+}22, WW03, WCL^{+}19, YX17, YP15, AZK10, DX03, Kra96, MM07, SMG09]. EIA [Fla02]. EIA/CDIF [Fla02]. Elastic [SR98, PP13]. electronic [SS08]. elements [SLCP01]. Eliminating [LM94]. embedded [LDNA03]. Emergence [ST15, XVN14]. Emergency [OT24, ZMM^{+}11]. Emotional [XZY23]. Empirical [BP94, HW03, HIG04, FDD05, ICC99, Joh96, LW97b]. empirically [SS03]. Emulation [BN22, ERL15, HYJ^{+}18, JN15, KKTM17, LBN^{+}18, CFS08]. enabled [CSRE21]. end [FHD09]. end-to-end [FHD09]. Energy [SFM13]. Engineering [VAB^{+}18, Fis92, FZ92]. Engines [CHA^{+}22]. Enhanced [WDYR16]. Enhancing [WhN20, WNFM04]. Entropy [WZ15, HLC^{+}10, PRO13, Rub02]. Enumeration [WPS13]. environment [CHIW98, SB01]. Environments [LT14, VVB^{+}20, CKP95, ZCLT04]. epidemic [BCD^{+}14]. equations [BC93, BHL13]. Equi [SFM13]. Equi-Energy [SFM13]. equidistributed [MN98]. Equine [XVN14]. Equivalence [ZBTT24, YS92]. equivalent [FMN00]. Error [WG16, WG04, HSN94]. Estimate [AMD23, KSW07, SW96]. estimates [CK14, NNB11]. Estimating [CMZ18, JFST24, JYE24, LC01, WCF23, HSN94]. Estimation [AGMW17, BLST16, JSD19, LN18, Mat05, SPYG24, VaAE02, WCZ16, YKA^{+}21, AK11, BKM09, DHM93, GAG14, HVA09, HVAPFY10, LCT07, NS06, Owe13, Raa93]. estimator [GK95]. Estimators [BC13, CN15, CERT15, AAAG06, AAGM10, AG07, Cal09, HIG04, LBTG10]. evacuation [LSJ10]. Evacuations [OT24]. Evaluating [CDS16, ZG94]. Evaluation [DTCU19, GGH^{+}23, HYJ^{+}18, KWU22, MRB^{+}18, TL18, ZH19, HD98, HD07, ICC99, PT00, SG91]. Event [BBMK16, BC13, ÇVS15, HSL^{+}19, HW19, JB22a, KSL^{+}16, Mar22, MJV^{+}15, PPT14, PCGM18, PTD^{+}20, RAGN19, RMWLP21, SP11, VWD22, VXE^{+}22, WYT^{+}20, WMC^{+}18, WZCJ22, YP15, AK11, BHLZ22, BL11, BHL13, BKV04, EK04, EK07, GLM96, HT99, HVAPFY10, HG01, LBTG10, Lin92, MWM91, MH92, MBGF11, MCC11, NOP99, Nic91, NY04, Nut06, Nut08, Pag93, PB96, RS94, RWU22, SJY03, TGT05, Vak92, YJ96]. Event-Based [MJV^{+}15]. event-driven [MWM91, MH92]. Events [RH19, GL05, Hei95, JB00, LDT07, LDF91, Rub02]. everyone [GDP14]. evidence [HW03]. evolution [PBF^{+}00, SC08]. Evolutionary [RGTL12, JC11]. Exact [BW15, DQZ18, DLQ20, HSL^{+}19]. Exact-Differential [HSL^{+}19]. Execution [DJS94, KPG15, PPT14, Di 23, SALS18, Van18, WWH^{+}23, NH96]. Execution-driven [DJS94]. Expanded [KSL^{+}16]. Expectation [LF13, STHL13, YKA^{+}21, LG03]. Expectations [AK18, CLL99]. Experiences [NCV06]. Experiment [RL15]. Experimental [Vig16, DHM93]. Experiments [FS17, Nel17, Di 23, WWH^{+}23, EU14, MKPR98, SWL09, YL96].

Explicit [HW21]. Explicitly [VVB^{+}20]. Exploiting [CN16, KSW07]. Exploration [SU16, Vig16]. Exponential [MJ15, QDZ21, BB99, LX14]. Exponentially [Hof11, Dev09]. Exposing [LBEJ19]. Expressive [HWMU17, Hil17]. Extended [GGH^{+}23]. Extending [Tuz95, VVB^{+}20]. extensions [Joh96]. Extrapolated [QF14]. Extreme [AGMW17, LHJS17]. Extreme-Scale [LHJS17]. Extrinsic [RLDH16].

Fabrics [ZL17]. Factor [XLZ17].

Factor-Based [XLZ17]. factorial [SS05]. factory [KO94]. failure [Nak94]. failure-biasing [Nak94]. fair [LBL01]. Falsification [ESZH21]. families [BB99]. Farming [San20]. Fast [AXE^{+}20, CHS95, DHK15, DM06, FDL99, GMOB01, Hei95, HD02, IFPM12, KWU22, Lem19, Qua19, RR93, SLF14, CFW99, HL03, JKS07, MR02]. Fast-forwarding [AXE^{+}20]. FastSlim [JSC01]. Fat [LHJS17]. Fat-Tree [LHJS17]. Feasibility [SSY21]. Feasible [ÇTI13, And06, BK10]. Feature [PBB16]. federated [RAF^{+}04]. federation [TAO08]. feed [SW13]. feed-forward [SW13]. feedback [JN05]. feedforward [SKR97]. few [JKS07]. Fidelity [AXE^{+}20, CFS08, KKTM17]. Field [SSDW18]. Fields [LCL16, SMI15, LX14]. figure [GCB95]. file [Mat05, WPN98]. file-access [Mat05]. file-access/modification [Mat05]. filtered [AQVA10]. filtering [BCL^{+}97]. Financial [CFL12]. Finding [BK10, Oso09, RL20, PS09, PK11]. Fine [PQ17]. Fine-Grain [PQ17]. Finite [GH15b, HHY11, KSW07]. Finite-State [GH15b]. Firings [HPS^{+}21]. first [CY10, DHM93]. first- [DHM93]. FISTE [FHD09]. Fitting [Che13]. Fixed [AK11, EH21]. Fixed-Confidence [EH21]. Fixed-Tolerance [EH21]. Flattening [BBMK16]. Flexible [KSW03]. Floating [Gou22, Doo07]. Floating-point [Gou22]. Flow [WCZ16, LBL01, PG14, VSCL13]. flow-level [VSCL13]. Flowpipe [DSR23]. Fluid [FDMS16, PH21, KW93, KM01, LPM^{+}04, MR02, NY04]. Fly [WMC^{+}18]. FNM [WDYR16]. folded [AAGM10]. FORECAST [TL18]. Forest [BHLZ22]. Form [MRB^{+}18, CO98, FSS95, RW93, Tuf97]. formal [ABGR01, GDP14, TL18]. formalisms [Bar97]. Formalization [ST15]. Formulation [SP11, SS08]. forward [SW13]. Forwarding [CF11, AXE^{+}20]. Fostering [GGH^{+}23]. foundation [BÖ96, RS94]. foundations [Bal97]. FPGAs [RAGN19]. fractional [IFPM12, SS05]. Framework [CDS16, ÇVS15, DC22, HW19, JBH^{+}22, LJS22, OT24, WFH12, XLZ17, BCL91, BCD^{+}14, CKP95, HN07, JC11, LSJ10, MBGF11, MY08, OLAM08, SC08, WCLG10]. Frequentist [JSD19]. fully [KN01, Kim05]. Function [CPRV23, HKP21, LG03]. Functional [GDB14, Bha07]. Functionals [SPYG24]. Functions [CFL12]. Future [San20]. fuzzy [BB94, MPK06].

gambler [KCK08]. Game [CN16, TKS16]. Games [JYE24, Vor10]. Gamma [QDZ21, Ros08]. gap [TTSM12]. Gate [GLC17]. Gate-Level [GLC17]. Gateway [CK08]. Gaussian [CWGZ24, DM06, GDB14, HE12, KDV^{+}20, LX14, WCCY19, WhN20, YN15]. GDCSim [GBA^{+}14]. Gene [FDP15]. General [DC22, RDSJ18, KSZ11, WS04]. Generalized [FL09, KC10, RL20, SSZ^{+}13, ZH19]. generate [BHG10, HD96]. Generated [ZLZ23, CFW99, FA06, Hör94]. Generating [ES94, KWU22, BN03, RR93, SS03]. Generation [CH23, EH95, GLC17, Lem19, LL15, QDZ21, Qua19, CL98, DHL10, Dev97, Dev09, GH03, GH06, GH09, HD02, HL03, Nie94, PG14, Wu01]. Generative [CH23, FBCS22]. Generator [LZW16, Bel05, EHG92, MN98, Pet91, Ros08, SM12]. Generators [Bre04, MZ91, MZ93, Vig16, DX03, Den05, Ent98, GK03, Joh96, LBC93, LW97b, MK92, MK94, Mat98, MWKA07, PL05, PW95, PJ10, SLF14, TL91, TLC93]. Generic [GP11]. Genetic [LZ20]. geometric [JC11]. Getting [WM99]. GFSR [MK92, MK94]. Gibbs [AQVA10]. Global [PE11, XYZ21, FH97]. Global-local [XYZ21]. Global-Scale [PE11]. Go [EH18, KH18]. good [LBC93]. GPGPU [SM12]. GPU [CB24, PF11, YP18]. GPU-based [PF11]. GPUs [LLCC13]. Gradient [HVA09, HVAPFY10, QF14]. Grain [PQ17]. Graph [DKVR09]. Graphical [WW95]. Graphs [MDH^{+}23, IMW00]. Green [FS17, FS21, Nel17, Pel21, GBA^{+}14]. Grid [HYJ^{+}18, VSCL13, ZK10]. Grids [YP18]. Guarantees [EH21, SJSM10]. Guest [Ano18, Bal97, CY10, CL98, DG10, FN03, GH15a, HHL14a, JW19, L’E03, MV02, MST17, TR08, TL18]. Guests [BSV16]. guided [NCN^{+}22]. GVT [PPT14].

Half [AK18]. Half-Spaces [AK18]. Halton [FL09, Tez93, TT94]. Hamming [WJ22, BV22]. Hamming-Weight [WJ22, BV22]. Hard [NH15, Kra96]. hard-sphere [Kra96]. Hardware [NAT^{+}21, PF11, SV97]. HAVEGE [SS03]. hazard [JS02]. HCSM [CKP95]. Healthcare [RSG21, EY11]. Heap [RH19]. Heavier [MJ15]. heavy [BL11, BHL13, FA06, HPA07, HS12, JS02, WW03]. heavy-tailed [BL11, BHL13, FA06, HPA07]. Hedging [AMD23]. Heterogeneous [NAT^{+}21]. Heteroscedastic [WCCY19]. heuristic [SS03]. heuristics [NZ07]. Hierarchical [BBMK16, KDV^{+}20, LJS22, CHIW98, KK00, SSRT91]. High [KKTM17, LCK11, SNS16, ZZC18, AZLT10, BHG10, BCD^{+}14, DX03, Doo07, Owe98, SQ12, Tuz95]. High-Density [ZZC18]. High-Dimensional [SNS16, DX03, Owe98]. High-Fidelity [KKTM17]. high-level [BHG10, SQ12, Tuz95]. high-performance [BCD^{+}14]. high-period [Doo07]. high-speed [AZLT10]. higher [BHG10]. Highly [RDSJ18, HSN94, HD07, Nak94]. Histograms [STHL13]. hit [KSZ11]. hit-and-run [KSZ11]. HIV [MCC11]. HLA [CTC^{+}05, LLT07, LCT^{+}15, LLHL00, PTCL11, RTY05]. HLA-Based [LCT^{+}15, CTC^{+}05]. HNS [MPW04]. Hölder [LX14]. Holistic [SALS18, Van18, BKV04]. Honoring [GH15a, Wil07]. hospital [GP11]. household [MCC11]. HPC [LHJS17]. HSL [SSRT91]. Hub [HHFS16]. hubs [KFL00]. human [GCB95, LSJ10]. Hybrid [ESZH21, HPS^{+}21, DSR23, BL02, EK04, EK07, LL02, SLCP01, VSS^{+}14, ZJTB04, vBBR03, MPW04]. hypercubes [HLC12].

I-Sim [BNSS24]. I/O [JSC01]. identification [HAFDP11]. Identify [GB19, RK20]. idle [WG04]. Iglehart [GH15a]. II [Cal09, MK94, UNMS97]. IID [DjWS19]. III [JB22a, JB22b]. Illustration [SFM13, WPW09]. Image [SMI15]. Image-Based [SMI15]. Impact [CKM23, YX17, ZK10]. Impacts [HAFDP11]. Implementation [BFN92, IMW00]. implementations [NCV06]. Implemented [RAGN19]. Importance [BGL12, DHN22, DLW07, RDSJ18, AK11, De 06, GK95, HS12, LC01, LV00, MSM10, NZ07, RJ04, RW93, SW13]. importance-sampling [De 06]. Improved [HKP21, HW21, JFST24, KDV^{+}20]. Improving [JZTB06, LCT^{+}15, RFA00, WS04]. IMSAT [NB93]. In-Depth [JBH^{+}22]. inaccuracies [JZTB06]. Incorporating [MCC11, NNB11]. increases [GH03, GH06, GH09]. incremental [BKV04]. Indemics [BCD^{+}14]. independence [EHN94b, Emm98, Lev01]. Independent [HAK14, De 06]. indices [Owe13]. indifference [KN01]. indifference-zone [KN01]. indirect [Mat05]. Industrial [XNH10]. Inequalities [BGL12]. Inference [FDP15, JKE14, RL15, SSZ^{+}13, WCS16, WCCY19]. Inference-Based [WCCY19]. Infinite [DjWS19, And06]. Information [LBEJ19, RS10]. INFORMS [HHL14a, BNSS24, CY10]. infrastructure [AK02]. Inhibition [RLDH16]. Inhomogeneous [BK20]. Initial [WG16, AAAG06, AGT92]. initialization [MWKA07]. initiating [FK91, Nic91]. inland [ZIC06]. innovations [BHL13]. Input [HSS24, UPB22, XNB16, YX17, ZLZ20, BN03, DM06, WW95]. Inputs [CH23, MR02]. insider [MMRC^{+}08]. insider-threat [MMRC^{+}08]. Insiders [CTF^{+}19]. Instability [SKR97]. instruction [MM07]. Integer [HWdF13, Lem19, Qua19, WPS13]. Integer-Ordered [WPS13]. integrals [LX14]. Integrated [HN09, YN15, YN20, Cal07, Cal09, Fis92, LDNA03, LSJ10, SB01]. Integrating [LCL16, ZH19, ZJTB04]. Integration [LBN^{+}18, EK04]. intelligence [Fis92]. Intelligent [ZBTT24, NB93]. Intensional [SU16]. Intensity [DHN22]. intensive [BCD^{+}14]. Inter [LBEJ19]. Inter-process [LBEJ19]. interaction [CS92, WCLG10]. interactions [BHG10, DG10, SF10]. interactive [BCL^{+}97, BCD^{+}14, MWMD07, SSH97, WW95]. interactively [QFL^{+}10]. Interest [LT14]. Interference [WAGP15]. International [LCK11]. Internet [ABGR01, CK08, KHJ^{+}08, Mat05, Nic08]. interoperability [SSH97]. Interpolation [WPS13]. interruptions [DOD93]. Intersection [LLCC13]. Interval [Gou22, HHH^{+}19, Lem19, Qua19, Sin14, PLM94]. Intervals [Nak14, SPYG24, CH04, CLL99, CN12, FG99, IMW00]. Intractable [JKE14]. Introduction [AM23, BNSS24, BB19, DT22, DR13, EY11, GC22, GH15a, HT20, HAA^{+}19, LK21, PW21, CY10, CL98, DG10, HHL14a, L’E03, TR08, Wil07]. Intrusion [PTE^{+}11]. invalidates [PJ10]. inventory [Lim12]. Inverse [HLD07]. inversion [DHL10, HD96, HL03]. Inversive [LW97b, EHG92, EHN94a, Emm98, Nie94]. invoked [LDF91]. IP [LPM^{+}04]. Irreducible [GH15b]. Issue [AM23, Ano18, BSV16, BNSS24, BB19, DR13, GH15a, HT20, JW19, LK21, MST17, PW21, QTP20, TL18, CY10, CL98, DG10, EY11, HHL14a, MV02, TR08, Wil07]. issues [SSH97, YJ96].

Jackson [JN05, KN02, MSM10, NZ07]. Joint [SJSM10, WhN20]. Jointly [NCN^{+}22].

Keddah [DTCU19]. Key [WZCJ22]. know [MFFR92]. Knowledge [FBS20, WXC^{+}23, ZBTT24, BÖ96]. Knowledge-based [WXC^{+}23, BÖ96]. known [DHL10, Ent98]. Kolmogorov [KW15]. Kriging [NY12, QF14, CAN12, CK14].

L [GH15a]. Ladder [RH19, TGT05]. Language [HWMU17, Hil17, Mar22, RWU22, EU14, SSRT91, TB98]. Large [CMZ18, CK08, GJ13, HSL^{+}19, KV23, LLCC13, PTE^{+}11, WCZ16, WMC^{+}18, WCL^{+}19, YP18, Buc98, Den05, FG98, LM94, LPM^{+}04, LLHL00, MR02, SS05, TGT05, UXC^{+}00, ZCLT04]. Large-Deviation-Based [WCZ16]. Large-Scale [HSL^{+}19, KV23, LLCC13, PTE^{+}11, WMC^{+}18, YP18, CK08, WCL^{+}19, FG98, LM94, LPM^{+}04, LLHL00, TGT05, ZCLT04].

Lateral [RLDH16]. Latin [Owe98, HLC12]. lattice [TLC93]. Laws [BBCD22]. layer [AZK10, BHG10, DG10, SF10, WCLG10, BHG10]. Leader [JFST24]. Learning [BLG^{+}21, CSRE21, GGH^{+}23, SCW13, ZLZ23, KT10]. Learning-enabled [CSRE21]. Least [SNS16]. lengths [SW96]. Level [GLC17, Hil17, WAGP15, BHG10, DOD93, SQ12, SS03, Tuz95, VSCL13, WPN98, WG04]. Lévy [CFL12, DLQ20]. like [LPPP13]. likelihood [BG93]. Likelihoods [JKE14]. Limit [CG13]. Limited [LL20]. limits [SKR97]. Lindley [KC10]. Lindley-type [KC10]. Linear [SSZ^{+}13, WPS13, Ent98, Ent99, GAG14, LW97b]. Linked [Mar22, RWU22]. links [KHJ^{+}08]. Linux [LBN^{+}18]. Lives [Mar22, RWU22]. Load [WYT^{+}20]. Local [BBCD22, DC22, HW21, XYZ21]. locally [HN07]. location [PB96]. log [Ley98]. log-concave [Ley98]. logic [RS94, SG91, Tuz95]. logic-based [RS94]. Logit [FFSF13]. Logs [CTF^{+}19, CPQ17, TFR07]. Long [NCN^{+}22, DX03, GMOB01]. long-cycle [DX03]. long-range [GMOB01]. lookahead [FK91, JB00]. loss [AO95, CHS95, LC01, LV00]. low [AG07, BFN92, DOD93, RGTL12, Cal09]. low-bias [AG07]. low-discrepancy [BFN92, RGTL12]. low-level [DOD93]. Lyapunov [BGL12].

MAC [BHG10]. MAC-layer [BHG10]. Machine [CSRE21, JN15, YP15, RBDH97]. MaD0 [LZW16]. maintenance [RNS97]. Major [HHFS16]. Making [LJS22, LSJ10]. Malicious [CTF^{+}19]. Management [LJS22, LT14, MD20, Pic24, WKC^{+}24, CTLZ05, DF97, FHD09, KM01, LP91, SQ12, WNFM04, ZLK91]. Managing [RH19]. Manufacturing [NB93]. Manufacturing-Simulation [NB93]. many [LPPP13, MR02]. Marginal [WG16, Ros08].

Marine [HHFS16]. Markov [AHO93, BDK^{+}19, Buc98, BHH21, GL05, HHH^{+}19, KW93, MR02, MBGF11, NH95, RK20, RJ04, SCW13]. Markov-reward [GL05]. Markovian [DHN22, HSN94, Nak94, RDSJ18, WCF23]. Marsaglia [Bre04, PW95, Vig16]. Marshall [BLST16]. Massive [SSZ^{+}13]. Massively [PCGM18, Vak92, HD98]. Matching [LT14, ZS17, PTCL11, RTY05]. Mathematical [CS08]. Max [Ale17, CS17, KT10]. max-norm [KT10]. maximal [GK03, Rub02]. Maximization [LF13]. Maximum [AGMW17, JKS07]. MAYA [ZJTB04]. MCMC [FFSF13]. Mean [BDK^{+}19, Hic96]. Mean-payoff [BDK^{+}19]. Means [AG16, AAAG06, Raa93, SLW^{+}05]. measure [HVAPFY10, WCLG10]. measure-valued [HVAPFY10]. measurements [BP94, CF11, LH02]. measures [BK10, De 06, HSN94]. Mechanism [LCT^{+}15, CTLZ05]. Mechanisms [BN22, LDF91, ABGR01, LL91a, MH92]. Memoization [SSDW18]. Memory [HKP21, NCN^{+}22, PTD^{+}20, TKS16, DF97, FH97, LW97a, LP91, MD20, UNMS97, UXC^{+}00, ZG94]. Mersenne [MN98, Nis00]. Mesoscopic [GZWG18, Lor18]. Message [SDZ^{+}15, WDYR16]. meta [Fla02]. meta-metamodel [Fla02]. Metamodel [XYZ21, TAO08, Fla02]. Metamodel-assisted [XYZ21]. Metamodeling [Fla02, KDV^{+}20, SNS16, WCCY19]. Metamodels [YN15, CAN12, DHM93, Fla02]. Method [FBCS22, LCL16, Tur17, YN20, CGN06, DJ11, GH03, GH06, GH09, HLC^{+}10, Hör94, KT10, Nak94, Nie94, Nut06, FDD05]. methodologies [Fis92, TR08]. Methodology [KPG15, Bal01, FZ92, LDNA03, LF99].

Methods [BMLY19, DR13, EH95, HHL14b, RL15, San20, WG16, ABGR01, And99, HDM03, ICC99, TL18, XGH12]. Metric [CHA^{+}22]. Metropolitan [CKM23]. Metropolitan-scale [CKM23]. microarchitecture [WWFH06]. Microscopic [AXE^{+}20, NCN^{+}22]. Middleware [PTD^{+}20]. Military [PBAB^{+}11, WPW09]. Minimum [MKPR98]. Minority [CN16]. Misra [SG91]. mission [SB01]. Mitigation [FDMS16]. mixed [LL02, QFL^{+}10]. mixed-signal [LL02]. Mixing [CPQ17]. Mixtures [WZ15, HS12]. MNO [Ale17, CS17]. Mobile [KH19, CSK10]. Mode [PH21]. Mode-switching [PH21]. Model [AP18, ÇVS15, CTF^{+}19, DHN22, FDD05, GLC17, HZF14, JFST24, JSD19, KKTM17, KPG15, MRB^{+}18, PCGM18, SP11, SSZ^{+}13, WhN20, WXC^{+}23, XZY23, ZLK91, EK07, FZ92, FSS95, KHJ^{+}08, LH02, LS92, LSJ10, MCC11, NOP99, RWK^{+}07, SF10]. Model-Based [HZF14, CTF^{+}19, LS92]. Model-Driven [ÇVS15]. Modeling [BSV16, Bar97, BL02, BHG10, BMLY19, BN03, BKV04, BDGP20, CH23, DWYM16, FW97, HWMU17, HW21, Hil17, HHY11, HM08, KH19, KZ11, LDNA03, LZ20, LPPP13, LHJS17, LDL04, Mar22, RWU22, RMWLP21, TKS16, WMC^{+}18, ZL17, ZZC18, ZLH^{+}22, Bal01, Bar03, BCD^{+}14, CSK10, DOD93, DG10, DKVR09, EY11, Fis92, GDP14, HPA07, KLF02, LL02, MBGF11, MV02, NY04, NCV06, RS94, RFA00, Sch10, TR08, Uhr01, WW95, WPN98, WG04, ZJTB04, ZCC^{+}10]. Modelled [VVB^{+}20]. Modelling [GZWG18, Lor18, OT24]. Models [BBMK16, BK20, ÇVS15, Che13, FFSF13, HT20, JKE14, KDV^{+}20, Nut20, PE11, SABF15, SU16, WhN20, YN15, YN20, BÖ96, BB94, BN09, CS08, FLV01, Hei95, LPM^{+}04, MPK06, MBGF11, MT06, Pac08, PB96, QFL^{+}10, RS10, RB08, SY95, TFR07, VSCL13, YS92, ZMM^{+}11, ZG94]. Modest [BHH21]. modification [CS92, Mat05]. modulus [EHG92]. Moment [FHG16, Lüc16, RL15, ZS17]. Moment-Based [RL15]. Moment-Closure [FHG16, Lüc16]. Moment-Matching-Based [ZS17]. Monkey [MZ93, PW95]. monotone [HD96, HLD07]. Monte [DR13, Pel21, CB24, DJLZ17, FS21, FSS95, HHL14b, LDT07, LV00, LG03, XGH12]. Monte-Carlo [FSS95]. Morphological [FDMS16]. Motion [BCM18, GCB95, IFPM12]. Movement [GZWG18, Lor18]. movements [LDL04]. Moving [SNS16]. MS [TTSM12]. MTSS [HHFS16]. Multi [And21, Con20, CM21, HSS24, Hil17, LZ20, Lor19, MHS19, EK04, MV02]. multi-agent [EK04]. Multi-Level [Hil17]. Multi-Objective [And21, LZ20, CM21]. multi-paradigm [MV02]. Multi-period [HSS24]. Multi-scale [Lor19, MHS19]. Multi-server [Con20]. Multiagent [ST15, ST13]. Multiclass [WCF23, KW93, RRP00, Tuf97]. Multicore [MKG^{+}17, TKS16, WAGP15, WDYR16]. Multicores [LBEJ19]. Multidimensional [BCZ14, Lim12, PS09, SS14, VAVA06]. multifacetted [ZLK91]. Multifractal [JFST24]. multihop [NNB11, SF10]. Multilevel [DJLZ17, HWMU17, SU16]. Multilevel-DEVS [SU16]. multimodel [FZ92]. multimodeling [LF99]. multinomial [VSS^{+}14]. Multiobjective [HAA^{+}19, MSK10]. multiparadigm [Bar03]. Multiple [HHH^{+}19, HAK14, HPS^{+}21, YN93, BK10, DN99, DOD93, Den05, KK00, LBC93, Nel93, PT00, SJY03]. multiple-comparison [DN99]. Multiplex [RMWLP21]. multiply [GK03]. multiply-with-carry [GK03]. multiprocessor [CG02, SY95].

Multiprocessors

[LBN^{+}18, MD20, DJS94, FH97].

multiresolution [RNS97]. Multiscale [DWYM16]. multiserver [KC10]. multisimulation [MY08]. Multistep [MWMD07]. Multistep-ahead [MWMD07]. Multitasking [LS92]. Multithreaded [KV23, LTM^{+}17]. Multivariate [SDLH12, XNB16, Bha05, Bha07, BN03, Dev97, HBE95, Ley98].

Nash [JYE24]. Nearly [LV00, HLC12]. need [MFFR92]. Neighborhood [WPS13]. NeMo [PCGM18]. Nested [DK22, YKA^{+}21]. Nets [HPS^{+}21, ZH19, BC93, BKV04, Hic96, Owe03]. Network [BN22, BHLZ22, BLST16, CERT15, CMM^{+}16, DTCU19, ERL15, FDP15, KKTM17, KPG15, LBN^{+}18, LL15, MPW04, NAT^{+}21, SABF15, VXE^{+}22, WNFM04, WZCJ22, CFS08, DKVR09, HPA07, JZTB06, KFL00, KN02, LM94, LALGSG^{+}00, MWMD07, MSM10, MT06, PF11, PRO13, RRW00, RAF^{+}04, SLCP01, SW13, SV97, VSCL13, ZJTB04]. network-computing [KFL00]. Networking [LCK11]. Networks [CDS16, CH23, FBCS22, JN15, KWU22, LJS22, LHJS17, MJ15, Pic24, PTE^{+}11, RL15, RMWLP21, WCS16, WMC^{+}18, WCL^{+}19, WKC^{+}24, ZLZ23, AZLT10, AO95, CS08, CO98, CSK10, DG10, EGLW93, FDL99, FLV01, GMOB01, JN05, KK00, Lim12, LPM^{+}04, LDL04, NZ07, RRP00, RW93, SLCP01, SJSM10, SKR97, SMG09, SF10, SPV^{+}10, Tuf97, UXC^{+}00, VaAE02]. Neural [BHLZ22, CH23, NAT^{+}21, ZLZ23, MWMD07]. neural-network [MWMD07]. Neuromorphic [PCGM18]. Neurons [LTM^{+}17]. Neutron [CB24]. Newton [Bha07]. Newton-based [Bha07]. NIM [CH23]. NoCs [JBH^{+}22]. node [De 06]. Noise [RLDH16]. Noisy [SCW13]. non [HSN94]. non-Markovian [HSN94].

Nonhomogeneous [SDLH12]. Nonlinear [EH95, LZW16, EHN94b]. Nonnegativity [Ale17, CS17]. nonsaturated [HLC12]. nonstationary [BN09]. nonuniform [Bel05]. norm [KT10]. Normalizing [DJLZ17]. NORTA [GH03, GH06, GH09]. Note [BSV16, Bre04, CHIW98, Hör94, TT94]. Novel [SSY21, KM01]. November [LCK11]. Noxim [CMM^{+}16]. Null [WDYR16]. Null-Message [WDYR16]. Number [Bre04, EH95, LZW16, MZ91, MZ93, Pet91, AK11, CL98, DX03, EHG92, Ent98, GK03, Joh96, LBC93, LW97b, MN98, MWKA07, PL05, PW95, PJ10, SM12, SLF14, TL91, TLC93, Wu01]. Numbers [GK19, Gou22, Pet91, Doo07, EHN94a, EHN94b, Ent99, Lev01, Nel93, SMDS11, SS03, WM99, CAN12]. Numerical [ZH19, DHL10, HL03]. numerically [EK07].

O [JSC01]. object [FG98]. Objective [And21, LZ20, CM21, FH18]. Objectives [HHH^{+}19]. Observation [JKE14]. Observation-Driven [JKE14]. off [KW93]. Olkin [BLST16]. Omnithermal [Con20]. On-GPU [CB24]. on-off [KW93]. One [PPT14]. One-Sided [PPT14]. Online [LF13, PBB16]. only [DHL10]. OOPM [LF99]. OOPM/RT [LF99]. Open [JBH^{+}22, PK11]. Open-Source [JBH^{+}22]. Operational [ZMM^{+}11]. Operations [PBAB^{+}11, RSG21]. Opportunities [San20]. Optimal [AZLT10, BKM09, CWGZ24, LP91, HLC^{+}10, Kaw10, LV00, PG14, RW93]. Optimisation [UPB22, UB24]. optimism [DF97]. Optimistic [CPQ17, JB22b, CPF99, Nut08, SQ12]. Optimization [And21, BDK^{+}19, CDS16, CG13, CM21, FBCS22, GDB14, HKP21, HSS24, HAA^{+}19, JBH^{+}22, LL20, Sch13, SPYG24, WPS13, And99, And06, BL02, Bha05, Bha07, BHM11, CSK10, HLC^{+}10, HDM03, HN07, HN09, MSK10, PG14, PN03, RGTL12, SJY03, XNH10]. Optimization-Based [CDS16]. Optimizations [DK22]. Optimizing [ELL00, LLCC13, DSR23, WPW09]. optimizing-simulator [WPW09]. Optimum [Tur17]. OR/MS [TTSM12]. order [Den05, DHM93, HD02]. Ordered [WPS13]. Ordering [Ale17, CS17]. Ordering-Piecewise-Quadratic [Ale17, CS17]. Organogenesis [SMI15]. oriented [KK00, SSRT91]. orthogonal [HLC12]. orthonormally [FG99]. Output [FS17, Nel17, XNB16, ZC18, CGN06, Cal07, Cal09, CH04]. overflow [DM06, JN05, NZ07]. overheads [BP94]. Overlapping [LB15, SPYG24]. Overview [PK11].

Packet [FLV01, AZLT10, CHS95]. PADS [Ano18, JW19, MST17, QTP20, DT22, GC22, LK21]. Pairwise [LLCC13]. PAM [DWYM16]. paradigm [MV02]. Parallel [ANSW23, BC93, BMLY19, BBCD22, ÇTI13, CG02, Ent99, Fuj16, JB22a, JN15, KSL^{+}16, MKG^{+}17, MD20, NH96, PCGM18, PTD^{+}20, RAGN19, RH19, SMDS11, SP11, UXC^{+}00, WDYR16, WYT^{+}20, WMC^{+}18, WCL^{+}19, WZCJ22, WKC^{+}24, XCA^{+}17, YP15, ZC18, AO95, CPF99, EGLW93, FW97, GH91, GLM96, HD98, HF01, LP91, LL91b, Lin92, MWM91, Nic91, NH95, RA97, TFR07, Vak92, Yau99, Pic24]. Parallelism [Lin92, SY95]. Parallelization [SSZ^{+}13]. Parallelizing [KCS20]. Parameter [RL15, SSDW18, WCS16, YN20, BKM09, NC06]. parameterization [LH02]. Parameterized [CKL^{+}13, BKM09]. parameters [KK00]. Parametric [BDK^{+}19, LL20]. Parametrized [Tur17]. Parasites [XVN14]. ParaSol [MKPR98]. Pareto [HHH^{+}19]. part [Lev01, JB22a, JB22b]. ParTejas [MKG^{+}17]. Partial [KWU22]. Particle [DWYM16, LF13, Sch13]. Particle-Based [LF13]. partition [Rub02]. parts [Emm98]. pass [MM07]. Passing [SDZ^{+}15]. patchwork [SZ99]. path [NNB11, RDSJ18]. Path-ZVA [RDSJ18]. paths [Cal07, Cal09]. patient [MBGF11]. Patterns [GB19, Di 23, WWH^{+}23]. Paved [STHL13]. payoff [BDK^{+}19]. PDES [CPQ17, GLC17, LTM^{+}17, LBEJ19, RAGN19, WAGP15]. PDES-A [RAGN19]. Pedestrian [GZWG18, Lor18, KZ11]. Penalty [HKP21]. Pending [RH19]. pentanomials [Wu01]. per-application [PRO13]. per-flow [LBL01]. Perfect [Con20, MT06]. Performance [AAGM10, AXE^{+}20, BCL^{+}97, BMLY19, BBCD22, CHA^{+}22, CSRE21, HD98, JBH^{+}22, KV23, KM01, LCK11, LCT^{+}15, LN18, MRB^{+}18, MJV^{+}15, Nic91, PT00, BK10, BCD^{+}14, FW97, GP11, HIG04, SKR97, UXC^{+}00, WS04]. period [Doo07, Emm98, GK03, HSS24, Lev01]. permutations [CN98]. Persistence [WYT^{+}20]. persistent [IFPM12]. Personal [HW21, LM94]. Personality [XZY23]. Personality-based [XZY23]. perspective [Vak92]. perturbation [BFMW03, BG93, MSK10]. Perwez [AGG^{+}07, Wil07]. Petri [BC93, BKV04, HPS^{+}21, ZH19]. phase [SWL09]. phenomena [QFL^{+}10]. Physical [Ano21, CTF^{+}19, HYJ21, BDH21, Pac08, QFL^{+}10, ZJTB04]. Physics [NCN^{+}22]. Physics-guided [NCN^{+}22]. Piecewise [Ale17, CS17, WPS13]. Piecewise-Linear [WPS13]. Placement [PTE^{+}11]. places [KZ11]. Planning [HYJ^{+}18]. Platform [PE11]. Platforms [YP15, YP18]. playback [GCB95]. Plot [TFR07]. Point [MDH^{+}23, Doo07, Gou22]. Poisson [SDLH12]. Polynomial [Tez93, CO98, LCT07, SS08, TT94]. polynomial-time [CO98]. polynomially [Dev09]. polynomials [GS12]. Pool [TKS16]. Population [PH21, NZ07]. Port [HHFS16, ZIC06]. portable [DX03, Den05, TL91]. Possibly [EH18, KH18]. Posterior [STHL13]. Potential [LCL16]. Power [JBH^{+}22, KW15, WCZ16, EHG92, PG14]. Powered [DTCU19]. PQRS [Ale17, CS17]. Practical [And21, CM21]. preattentive [HBE95]. Precise [BN22]. Predicting [MH19]. Prediction [CSRE21, MJV^{+}15, YN15]. Predictions [WhN20]. Predictors [BHLZ22, MWMD07]. prefetch [JSC01]. prefetch-safe [JSC01]. Prescriptive [OT24]. presence [AAAG06, BK10]. Present [San20]. prespecified [Ros08]. Prevention [PTE^{+}11]. Price [JYE24]. primitive [Wu01]. Principles [Wai15, Ano18, JW19, MST17, QTP20]. priority [RA97, TGT05]. Probabilistic [ESZH21, GHS18, Par18, RL20, TRK^{+}09, Vor10]. Probabilities [BC13, CMZ18, DSR23, DM06, JN05, JKS07, RRP00, SW96, VaAE02]. Probability [CPRV23, HT99]. probably [Oso09]. Problem [CERT15, AGT92, BG93, HVA09, Kra96, PS09, PK11, QC02, WPW09]. Problems [UB24, YKA^{+}21, Rub02]. Procedure [GK19, WFH12, BKM09, DHM93, KN01, PN03, Raa93, SWL09, SLW^{+}05, VSS^{+}14]. Procedures [EH21, HAK14, MH19, Sin14, DN99, Kim05, SJY03]. Proceedings [LCK11]. Process [BDGP20, KDV^{+}20, WCCY19, WhN20, YN15, CS92, KLF02, Kiv91, LBEJ19, RFA00, RD10, SSRT91, TB98]. process-based [TB98]. process-oriented [SSRT91]. Processes [CWGZ24, CFL12, HHH^{+}19, MDH^{+}23, RMWLP21, SDLH12, SCW13, WCF23, BN03, GL05, JS02, LALGSG^{+}00, WPN98]. Processing [BMLY19, HSL^{+}19, HBE95, HM08]. Processor [PPT14, QC02]. Processors [MKG^{+}17]. Product [MRB^{+}18, CO98, FSS95, RW93, Tuf97]. Product-Form [MRB^{+}18, CO98, FSS95, RW93]. Production [GCB95]. professional [AGG^{+}07]. Profile [CSK10]. Profile-driven [CSK10]. Programming [GHS18, LZ20, Par18, XYZ21, CS08, HE12]. Programs [LB15]. Projections [SDLH12, KT10]. Propagation [SP11]. Properties [JFST24, JSD19, RK20, Van19, VLN^{+}19, Emm98, HPA07]. ProPPA [Par18, GHS18]. Protocol [CK08, VXE^{+}22]. protocols [JB00, NNB11]. Provenance [Di 23, WWH^{+}23]. Provisioning [LCT^{+}15]. pseudo [MN98]. pseudo-random [MN98]. Pseudorandom [EH95, LZW16, Nie94, EHG92, EHN94a, EHN94b, Emm98, Ent98, Joh96, LW97b, Lev01, Mat98, MWKA07, PW95, SM12, SLF14, WM99]. public [HVA09]. Purdue [KFL00].

Q [KT10]. Q-learning [KT10]. QEST [AM23, PW21, BB19]. QoS [ABGR01, FHD09, KK00]. QRF [CDS16]. Quadratic [Ale17, CS17]. qualitative [BB94, FZ92, IMW00, LS92]. Quality [LB15, Hör94]. Quantification [ZLZ20]. Quantifying [YX17]. Quantiles [AGMW17, Nak14, SPYG24, CN12]. quantitative [TL18]. Quantum [Pic24, WZCJ22, WKC^{+}24]. quasi [LDT07]. quasi-Monte [LDT07]. queries [ST13]. questions [PK11]. Queue [RH19, De 06, DM06, MR02, RA97, TGT05]. Queueing [MJ15, NH15, WCS16, CS08, Hei95, Lim12, RS10, SKR97, SMG09, SF10]. Queues [AMD23, BW15, CMZ18, Con20, AO95, ELL00, FDL99, GK95, KC10, MT06, WW03]. Queuing [XZY23, DOD93, PF11, RW93, Tuf97]. Quick [KW93]. quickly [Oso09].

R [WPS13]. R-SPLINE [WPS13]. Radio [SP11, HAFDP11]. radio-identification [HAFDP11]. Radix [Joh96]. Radix- [Joh96]. rail [LDL04]. Railway [DK22]. Random [BHLZ22, Bre04, CAN12, Che13, CG13, DHL10, Dev97, Dev09, GK19, Gou22, HWdF13, HZF14, Lem19, MZ91, MZ93, Pet91, QDZ21, Qua19, STHL13, Wu01, YN15, And99, Bel05, CL98, DX03, Doo07, DLW07, Ent99, ES94, GH03, GH06, GH09, GK03, HN07, Hör94, HL03, HS12, JKS07, LBC93, LX14, MN98, Nel93, PL05, PJ10, RR93, RB08, SMDS11, SS03, TL91, TLC93]. random-number [Pet91]. random-search [HN07]. randomization [Buc98]. randomized [CO98, Hic96]. randomly [KHJ^{+}08]. randomness [KCK08, MK96]. range [GMOB01, ST13]. Ranking [ANSW23, CWGZ24, EH18, EH21, FH18, GK19, KH18, MH19, PHP^{+}15, ZS17, SJY03]. Ranking-and-Selection [EH21]. Rapid [LH02]. Rare [BHLZ22, BHL13, BC13, LDT07, AK11, BL11, GL05, HT99, Hei95, LBTG10, Rub02]. Rare-event [BHLZ22, BHL13, LBTG10]. Ratatoskr [JBH^{+}22]. Rate [Ale17, CS17, JS02, LBL01]. rates [CHS95, Mat05]. ratio [Hör94, LCT07]. ratios [BG93, CLL99, LC01]. RCR [Ale17, And21, Bee18, Hil17, KH18, Lor18, Lor19, Lüc16, Nel17, Par18, Qua19, Van18, WJ22, Mar22, Van19]. Re [PJ10, XVN14]. Re-Emergence [XVN14]. Re-seeding [PJ10]. Reachability [DSR23]. Reaction [KWU22, RL15]. Reactions [LTM^{+}17]. Real [CFW99, LCL16, HBE95, LF99, MY08, WNFM04]. Real-Time [LCL16, CFW99, HBE95, LF99, MY08, WNFM04]. Realistic [OT24, SABF15]. reality [QFL^{+}10]. really [MFFR92]. Rearchitecting [AK02]. reasoning [LS92]. reconfigurable [SV97]. reconstruction [Pac08]. Recovery [CPQ17]. Recurrence [GH15b, BC93, BHL13]. Recursive [CERT15, Den05, KC10, LBC93]. redistribution [HT99]. reduce [CN98]. Reducing [NC06, HIG04]. Reduction [CB24, Nak14, SMG09, SK23, UPB22, WXC^{+}23, AHO93, CN12, JSC01, KSW03, Kaw10, MWMD07, Tuf97]. Redundancy [NOP99]. reference [KSW03]. Reflective [DK22, Uhr01]. Regenerative [CN15, HG01, MJ15, CN98, CGN06, HIG04, KSW07]. regions [And06, KSZ11]. Regression [SNS16, CSK10, GAG14]. Regularly [STHL13, DLW07]. Regulatory [FDP15]. Reinforcement [BLG^{+}21, GGH^{+}23]. Rejection [HD96, Bel05, HLD07, Ley98, SZ99]. Rejection-inversion [HD96]. Related [DQZ18, FDMS16]. relative [HSN94]. relaxation [EGLW93]. Relaxing [XCA^{+}17]. relevance [BCL^{+}97]. Reliability [BDGP20, BLST16, CERT15, WCZ16, BCL^{+}97, Hei95]. Reliable [RDSJ18, Nak94]. renewable [PG14]. Repast [NCV06]. Repeated [FS17, Nel17]. Replicated [AAAG06, Ale17, And21, Lüc16, Nel17, Qua19, WJ22, GH91, Bee18, Hil17, KH18, Lor18, Lor19, Par18, Van18]. Replication [Ano21, Pel21, Di 23]. Report [Ale17, And21, Ano21, Bee18, Hil17, KH18, Lor18, Lor19, Lüc16, Mar22, Nel17, Par18, Pel21, Pic24, Qua19, Di 23, Van18, Van19, WJ22]. representation [FDD05]. representations [KC10]. Reproducibility [Pic24]. Rerouting [CKM23]. Resampled [CN15]. Research [Fuj16, HHL14a, RSG21, CY10]. Resilience [WAGP15]. Resilient [VAB^{+}18]. Resistance [XVN14]. Resource [LCT^{+}15, TKS16, AZK10, FSS95, ZK10]. Resources [BDGP20]. Response [CPRV23, WhN20]. RESTART [VAVA06]. Restricted [DSR23, VSS^{+}14]. Results [Ale17, And21, Ano21, Bee18, Hil17, KH18, Lor18, Lor19, Lüc16, Nel17, Par18, Pel21, Qua19, Di 23, Van18, WJ22]. retraction [LDF91]. Retrospective [PS09, WPS13]. Retrospective-approximation [PS09].

Reuse [Di 23, WWH^{+}23]. Reusing [EH18, FS17, KH18, Nel17]. Reverse [GLC17, CPF99]. Reversed [BW15]. Reversibility [CPQ17]. Reversible [PP13, SP11]. Review [HHL14b, MKT21, RSG21]. revolution [PBF^{+}00]. reward [GL05]. Rewards [DHK15]. rid [WM99]. Risk [FDMS16, HHL14b, XLZ17, ZLZ20]. risks [MMRC^{+}08]. RNGs [Mar03]. Road [XCA^{+}17]. Robust [GGH^{+}23, HHH^{+}19, LL20, Nel93, PBAB^{+}11]. Robustness [FBCS22, LBTG10]. Role [ZZC18]. Role-Dependent [ZZC18]. rollback [LL91a, LSW91]. rollback-based [LSW91]. Root [RL20, PS09, PK11]. root-finding [PS09, PK11]. routing [BG93, RRP00]. RT [LF99]. ruin [KCK08]. Rule [WG16]. Rules [Sin14]. run [HLC12, KSZ11, SW96]. run-variable [HLC12]. running [KFL00]. Runtime [HERU15, CSK10].

Safe [CPRV23, GGH^{+}23, JSC01]. Sample [LCT07, CK14, HDM03]. Sample-based [LCT07]. Sampler [SFM13]. Samplers [DJLZ17, AQVA10]. Samples [DjWS19]. Sampling [BGL12, BW15, DHN22, HAK14, Hof11, RDSJ18, De 06, DLW07, GK95, HS12, Kaw10, KSZ11, LC01, LV00, Ley98, MSM10, MT06, NZ07, Owe98, RJ04, RW93, SW13, SZ99, WWFH06]. saturate [KHJ^{+}08]. SC’11 [LCK11]. scalability [JZTB06]. Scalable [CSRE21, LPM^{+}04, WZCJ22, YP18, BCL^{+}97, HD98]. Scale [HSL^{+}19, KV23, LHJS17, LLCC13, PE11, PTE^{+}11, WMC^{+}18, YP18, AD92, CK08, CKM23, FG98, LM94, LPM^{+}04, Lor19, LLHL00, MHS19, PT00, TGT05, WS04, WCL^{+}19, ZCLT04]. scale-down [CK08]. scanning [KHJ^{+}08, RB08]. Scenario [HHFS16, LL20, CKP95]. scenarios [BHG10, LSJ10]. Scheduling [AMD23, CB24, AZLT10, HM08, QC02, SJSM10].

Scheme [WZ15]. Schemes [JSD19, SW13]. Scientific [CSRE21]. SCORE [FH18, PHP^{+}15]. Scrambled [Vig16]. scramblings [Owe03]. Screening [ACL15, NS06, SWL09, TRK^{+}09]. SDEs [BKM09]. Search [Che13, CG13, EH18, ESZH21, HZF14, KH18, WPS13, And99, HN07, LBC93]. Seattle [LCK11]. second [DHM93]. second-order [DHM93]. Section [DT22, GC22]. Sectioning [Nak14]. seeding [PJ10]. segmentation [AO95]. SEH [AMD23]. select [ICC99]. Selecting [Sin14, WFH12]. Selection [And21, ANSW23, CWGZ24, CM21, EH18, EH21, FH18, GK19, HAK14, KH18, MH19, PHP^{+}15, WFH12, YN20, ZS17, KN01, NS06, SJY03, VSS^{+}14]. Self [HWdF13, VAB^{+}18, FK91, FMN00, LALGSG^{+}00, Mat98, Nic91, PT00]. Self-Avoiding [HWdF13]. self-initiating [FK91, Nic91]. self-similar [FMN00, LALGSG^{+}00, PT00]. Self-Stabilisation [VAB^{+}18]. self-test [Mat98]. Semantics [HWMU17, Hil17, TB98]. semi [CGN06]. semi-regenerative [CGN06]. Semiautomatic [SDZ^{+}15]. semidefinite [HE12]. sensitivity [BL02, Owe13, WCLG10]. sensor [SF10]. sequence [Mat98]. sequences [BFMW03, BFN92, FL09, MK96, RGTL12, Tez93, TT94]. Sequential [ACL15, DJLZ17, DK22, GK19, JSD19, RH19, DHM93, GAG14, KN01, Kim05, Raa93, RA97, SY95, XGH12]. Sequentially [ZLZ23]. Serial [SSZ^{+}13, NH96]. Series [JKE14, SPYG24, BN03, BN09, FG99, SS14]. server [Con20, HHY11]. Service [RSG21, WCF23, CFS08, LM94, ZK10]. services [HVA09, HD07]. SESSL [EU14]. set [MPK06]. sets [Lim12]. Setwise [AQVA10]. several [ICC99, Raa93]. Shahabuddin [AGG^{+}07, Wil07]. shapes [Ros08]. Shared [MD20, PTD^{+}20, CHS95, FH97, KM01, UXC^{+}00]. Shared-memory [MD20, FH97, UXC^{+}00]. Sharing [PQ17, FSS95]. Sharpening [HE12]. Short [NCN^{+}22]. Short-Term [NCN^{+}22]. Sided [PPT14]. Signal [SP11, LL02]. Sim [BNSS24]. similar [FMN00, LALGSG^{+}00, PT00]. SimOS [RBDH97]. Simple [DHN22, Mat98, Nak94]. Simplifying [DOD93]. Simulate [BDGP20, DC22, ZLZ23, RJ04]. Simulated [HW21]. Simulating [CKM23, CFL12, DTCU19, GL05, JS02, SDLH12, SMI15, TDR^{+}11, EK04, EK07, GS12, LL02, NH95, XVN14]. Simulation [AK18, And21, And22, And06, AG16, Ano18, BHLZ22, BB99, BNSS24, Cal07, Cal09, CHA^{+}22, CMM^{+}16, ÇTI13, CH23, ÇVS15, CAT22, Che13, CG13, Con20, CM21, DQZ18, DLQ20, ERL15, FBS20, FS17, FS21, Fuj16, GJ13, HHL14a, HT20, HKP21, HSL^{+}19, HYJ^{+}18, HYJ21, HLC^{+}10, HSS24, HERU15, HWMU17, Hil17, HHFS16, HAA^{+}19, JB22a, JV23, JN15, JW19, KH19, Kiv91, KPG15, KSL^{+}16, LL15, LCT17, LHJS17, LCL16, MH19, Mar22, MDH^{+}23, MJ15, MST17, MKT21, NB93, NCN^{+}22, Nel17, Nut20, OT24, Pel21, Pic24, PCGM18, PTD^{+}20, QTP20, RAGN19, RK20, RWU22, RSG21, RMWLP21, SNS16, Di 23, Sch10, SABF15, SW96, SSDW18, VVB^{+}20, VWD22, Wai15, WPS13, WDYR16, WCCY19, WhN20, WYT^{+}20, WWH^{+}23, WMC^{+}18, WCL^{+}19, WZCJ22, WKC^{+}24, XNB16, XYZ21, XCA^{+}17, YKA^{+}21, YX17, YP15, YN15, ZMM^{+}11]. Simulation [ZC18, ZLZ20, ZH19, AAGM10, AD92, AO95, BC93, BCL91, Bal01, Bar03, BL02, BCL^{+}97, Bha05, Bha07, BHM11, BÖ96, BL11, BHL13, BB94, Buc98, CGN06, CHS95, CFW99, CTC^{+}05, CH04, CFS08, CY10, CG02, CHIW98, DG10, DM06, DHM93, DJS94, EY11, EU14, FDL99, FK91, FA06, Fis92, FSS95, FG98, GMOB01, GCB95, GP11, HT99, Hei95, HD98, HG01, HN07, HHY11, HN09, HM08, IMW00, JB00, JZTB06, JSC01, JN05, JKS07, KSW07, KFL00, KW93, KN01, KLF02, KZ11, KN02, LBTG10, LV00, LW97a, LDNA03, LS92, LF99, LLT07, LP91, LL91b, Lin92, LM94, LALGSG^{+}00, LLHL00, LSW91, MWM91, MR02, MPK06, MBGF11, MCC11, MY08, NOP99, Nic08, NZ07, Nut06, Nut08, OLAM08, Pag93, PCT97, PBF^{+}00, PF11, PN03, RS94, RFA00, RNS97, RAF^{+}04]. simulation [RWK^{+}07, RD10, RS10, SWL09, SSRT91, SSH97, SLCP01, SS14, SY95, SMG09, SG91, SPV^{+}10, SLW^{+}05, SV97, SC08, SS08, SJY03, TGT05, TR08, TTSM12, TB98, UNMS97, Uhr01, Vak92, Vor10, WW95, WS04, WW03, WNFM04, WWFH06, XNH10, XGH12, YL96, Yau99, YN93, YS92, YJ96, ZCC^{+}10, Bal97]. Simulation-Based [CG13, ZMM^{+}11, Vor10]. simulation-generated [FA06]. simulationists [MFFR92]. Simulations [AXE^{+}20, DK22, GB19, GRK^{+}15, HSL^{+}19, HAK14, HW19, LCT^{+}15, LLCC13, NY12, NH15, RH19, VXE^{+}22, XLZ17, XZY23, YP18, AHO93, BP94, BN09, CTLZ05, CN98, CPF99, CF11, DN99, EGLW93, GH91, GLM96, GAG14, HIG04, HF01, KSW03, KM01, LPM^{+}04, LX14, Nak94, Nic91, Oso09, Owe98, PP13, ST13, Tuz95, VSCL13]. Simulator [KCS20, MKG^{+}17, FW97, GBA^{+}14, RBDH97, UXC^{+}00, WPW09, MPW04]. Simulators [DK22, KWU22, LBN^{+}18, NAT^{+}21, NH96, OLAM08, SKR97]. Simulink [ZL17]. Simultaneous [JB00, YN20, BFMW03, MSK10, Raa93]. single [MM07]. single-pass [MM07]. singularities [EK07]. SIP [HHY11]. Site [SABF15]. Site-Specific [SABF15]. Size [AMD23]. Skeletons [SDZ^{+}15]. Slim [WMC^{+}18]. Small [BC13, LC01, Owe13]. Smart [HYJ^{+}18]. Smirnov [KW15]. Smoothed [GDB14, Bha07, BG93].

Smoothing [Ale17, CS17, AHO93]. Sobol’ [Owe13]. Social [CN16, LJS22, WCL^{+}19]. Society [BNSS24, HHL14a, CY10]. Software [CPQ17, JN15, KPG15, SDZ^{+}15, Fis92, SS03, SC08, XNH10]. Software-Defined [JN15]. Solution [LB15, YN15]. solutions [PK11]. Solved [BK20]. some [Joh96]. sort [PTCL11, RTY05]. sort-based [PTCL11, RTY05]. Source [JBH^{+}22, KK00]. Source-oriented [KK00]. sources [FMN00, KW93, WG04]. Space [HW21, LT14, HPS^{+}21, PLM94, ZCLT04]. Space-Time [LT14]. Spaces [AK18, Sch13, LG03]. Sparse [MDH^{+}23]. Spatial [LBEJ19]. Spatially [FHG16, Lüc16]. Spatio [Lor19, MHS19, VLN^{+}19]. Spatio-temporal [Lor19, MHS19, VLN^{+}19]. Spatiotemporal [LL15, Van19]. Special [AM23, Ano18, BSV16, BNSS24, BB19, DT22, GC22, GH15a, HT20, JW19, LK21, MST17, PW21, QTP20, TL18, CY10, CL98, EY11, HHL14a, TR08, DG10, DR13, MV02, Wil07]. Specialized [KWU22]. Specific [SABF15, EU14]. specification [Nut08, vBBR03]. specifications [NOP99]. spectral [Ent99, HN98]. Speculative [BBCD22, PTD^{+}20]. speed [AZLT10]. sphere [Kra96]. SPICE [SS08]. SPICE-type [SS08]. Spiking [NAT^{+}21]. SPLINE [WPS13]. splittable [SLF14]. Splitting [WCZ16, AK11, LDT07]. Spread [TDR^{+}11, KHJ^{+}08, XGH12]. square [Hic96]. Squares [SNS16]. Stabilisation [VAB^{+}18]. Stability [JYE24]. Stable [DQZ18, Hof11, QDZ21, Dev09]. staffing [ZMM^{+}11]. stage [DN99, KLF02, PG14]. Staged [WS04]. Stages [GGH^{+}23]. Standard [SM12, WG16, Kim05]. standardized [FG99]. standards [TTSM12]. standards-based [TTSM12]. State [AGMW17, AXE^{+}20, BK20, CPQ17, GH15b, HPS^{+}21, MSM10, NH15, PB96, Pic24, WKC^{+}24, AG07, DN99, De 06, EK07, GAG14, HG01, HIG04, SLW^{+}05, VAVA06, YN93]. State-dependent [MSM10]. state-independent [De 06]. State-space [HPS^{+}21]. Static [BLST16, SDZ^{+}15, ELL00]. Stationarity [AGT92]. Stationary [BW15]. Statistical [AP18, Che13, Emm98, GRK^{+}15, JSD19, Lor19, MHS19, SPYG24, WWFH06, ZC18, EHN94b, JC11, Lev01]. statistics [HD02, DR13]. Steady [AGMW17, BK20, NH15, AG07, DN99, GAG14, HG01, HIG04, SLW^{+}05, YN93]. Steady-State [AGMW17, BK20, NH15, AG07, DN99, GAG14, HG01, HIG04, SLW^{+}05, YN93]. Stealing [KV23, WYT^{+}20]. Steepest [MSK10]. Steepest-ascent [MSK10]. Stepped [YP18]. Stochastic [BHM11, CDS16, CK14, GHS18, GDB14, GH15b, HSS24, HZF14, HPS^{+}21, JYE24, KWU22, Lim12, LTM^{+}17, LBEJ19, LB15, NY12, Par18, QF14, RL20, RL15, SNS16, DSR23, SS08, Van19, VLN^{+}19, WhN20, XNB16, XLZ17, XYZ21, YX17, ZLZ20, ZH19, And99, BC93, BFMW03, Bha05, BHL13, BN09, BCZ14, CAN12, HDM03, KT10, NC06, PS09, PK11, PG14, RB08]. Stochastically [HKP21, PHP^{+}15]. Stopping [Sin14, GAG14]. Storage [LCK11]. strata [Kaw10]. strategic [ZMM^{+}11]. Strategies [HHH^{+}19, KV23, TRK^{+}09, ZK10]. Strategy [MRB^{+}18]. stratified [Kaw10]. Streaming [HSS24]. streamlined [MPW04]. streams [Ent99, MM07, Yau99]. strength [XNH10]. stroke [MBGF11]. Strong [MK96, SS03]. Structural [YS92, SC08]. Structure [BBMK16, SU16, VVB^{+}20, Bar97, Bar03, DOD93, KSW07, MCC11, TGT05, TLC93]. structures [Uhr01]. Studies [Che13, HHFS16, SSDW18]. Study [RK20, CFS08, FL09, FDD05, LL91a, NH95, PCT97, RA97, RBDH97, SY95].

Subordinator [DLQ20]. subsequences [Ent98]. Subset [And21, CM21]. subsetting [JC11]. SubsetTrio [JC11]. Subsolutions [BGL12, DHN22]. subtract [TLC93]. subtract-with-borrow [TLC93]. successes [AK11, TR08]. sums [BL11, DLW07, JKS07]. supercube [Owe98]. Superdense [Nut20]. Superfast [GLM96]. superior [Pet91]. supplies [Pet91]. Supply [MKT21]. Support [PTE^{+}11, MY08, RD10, Tuz95]. Supporting [DK22, LLHL00]. Supremum [BCM18]. surrounding [OLAM08]. Survey [AP18, ZLH^{+}22, RD10, SJY03]. Swapping [DjWS19]. switch [CHS95]. switched [EGLW93, HM08]. switches [LC01]. switching [PH21]. Symbiotic [ERL15, MY08]. Synchronised [ST13]. Synchronization [HYJ21, JB22a, JB22b, MH92, XCA^{+}17, MKPR98, QC02, SQ12]. Synchronous [EGLW93]. Synthesis [SDZ^{+}15, Fis92, IFPM12]. System [HHFS16, PQ17, PTE^{+}11, VWD22, DX03, Fis92, FSS95, FG98, ICC99, KM01, LW97a, LS92, MMRC^{+}08, MKPR98, Nut08, RS10, SB01, WPN98, ZIC06, ZK10, vBBR03]. systematic [BHG10]. Systems [Ald18, Ano21, Bee18, BDH21, ÇTI13, CKM23, DWYM16, ESZH21, FHG16, GHS18, GH15b, HWdF13, HYJ^{+}18, HYJ21, JV23, KH19, KSL^{+}16, LBEJ19, LHJS17, Lor19, Lüc16, MHS19, Par18, RDSJ18, ST15, Van19, VAB^{+}18, VLN^{+}19, WAGP15, WDYR16, ZBTT24, Bar97, BL02, BK10, BKV04, EK04, EK07, HSN94, HVA09, HVAPFY10, HD98, HG01, HM08, LV00, LDNA03, LLT07, LPPP13, Lim12, LL02, MWM91, NC06, Oso09, RBDH97, ST13, Vak92, VAVA06, ZLK91, TL18, Nak94].

Tables [Nis00]. tactical [ZMM^{+}11].

TADSim [MJV^{+}15]. Tail [MJ15, JKS07]. tailed [BL11, BHL13, FA06, HPA07, JS02]. tails [DLW07, HS12]. tandem [CS08, De 06, GK95, HHY11, KC10, KN02, MSM10]. Tapeworm [UNMS97]. targeted [CFS08]. Tausworthe [TL91]. TCP [CFS08, NY04, PT00, VSCL13]. TCP-targeted [CFS08]. Technical [CHIW98]. technique [BN03, Ley98, MM07, SLCP01, SZ99, WS04]. Techniques [Nak14, SDZ^{+}15, ZLH^{+}22, Bal97, CN12]. technologies [ZCC^{+}10]. technology [Kiv91]. telecommunications [GMOB01]. teletraffic [AQVA10]. Temperature [MJV^{+}15]. Temperature-Accelerated [MJV^{+}15]. Tempered [DQZ18]. tempering [WM99]. Temporal [GB19, LBN^{+}18, VXE^{+}22, IMW00, Lor19, MHS19, RJ04, Tuz95, VLN^{+}19]. Term [NCN^{+}22]. terminals [ZIC06]. Terrain [SSH97]. Test [BV22, WJ22, Ent99, HN98, Mat98, PW95]. Testbed [WZCJ22]. Testing [VXE^{+}22, WG16, CK08]. Tests [Ano21, BDH21, KCK08, KW15, MZ93, BFN92, Joh96, LW97b, PJ10]. Their [CFL12, HPA07]. Theorems [CG13]. theoretic [MPK06]. theoretical [AG07, WCLG10]. Theory [Nut20, PW95, HT99, MMRC^{+}08, Pet91]. Third [HHL14a]. threat [MMRC^{+}08, SB01]. Three [RH19, Bha05, NCV06]. Three-tier [RH19]. three-timescale [Bha05]. throughput [SJSM10]. tier [RH19]. Tightly [KSL^{+}16]. Tilted [Hof11, QDZ21, Dev09]. Time [AD92, AO95, BN22, BCM18, BW15, CPRV23, HYJ21, JKE14, JB22a, JB22b, LT14, LCL16, Nut20, PQ17, SPYG24, YP18, ZCLT04, BDK^{+}19, BN03, BN09, Buc98, CTLZ05, CO98, CFW99, DF97, DNRD96, FA06, FG99, FG98, FH97, GH91, HBE95, HPA07, KSW07, LF99, LP91, LL91a, LL91b, LDF91, MY08, NH95, PT00, PLM94, QC02, SQ12, SS14, SR98, WNFM04, Yau99]. Time-Based [HYJ21]. time-division [LL91b]. time-management [SQ12]. Time-Reversed [BW15]. Time-segmentation [AO95]. time-series [BN03, BN09]. Time-Sharing [PQ17]. Time-space [ZCLT04]. Time-Stepped [YP18]. Times [DC22]. timescale [BFMW03, Bha05]. Timestepping [BBCD22]. timing [DJS94]. TLM [SP11]. TLM-Based [SP11]. Tolerance [EH21]. TOMACS [Ano18, JW19, MST17]. Tool [NB93, ZL17, SSRT91, SPV^{+}10]. toolkit [NCV06]. Tools [GZWG18, Lor18, KFL00, RD10]. topological [CK08]. topologies [DKVR09]. topology [KK00]. Trace [KCS20, JSC01, KSW03, MM07]. Trace-Driven [KCS20]. Traffic [AXE^{+}20, CKM23, DTCU19, DK22, HHFS16, LL15, NCN^{+}22, XCA^{+}17, GMOB01, HPA07, LH02, MWMD07, NY04, PT00, PRO13, WW03]. train [LDL04]. Trained [NCN^{+}22]. Training [ZLZ23, Bal97, SSH97, SB01]. trajectory [BKM09]. Transfer [BLG^{+}21]. Transformation [AGMW17]. Transformations [KW15]. transformed [HLD07]. Transience [GH15b]. Transient [WG16, AAAG06, AGT92, HSN94, MR02]. Transitioning [NAT^{+}21]. Transmission [PE11]. Transparent [SQ12]. Transparently [CPQ17]. Transport [CB24, ZIC06]. transportation [HVA09]. Trap [UNMS97]. Trap-driven [UNMS97]. traveling [CFW99]. Tree [LHJS17]. triangulations [ES94]. trinomials [MK96]. Truncated [DLQ20]. Trusted [Ald18, Bee18]. TSTL [Van19, VLN^{+}19]. tuberculosis [MCC11]. Twins [WXC^{+}23, ZBTT24]. Twisted [MK92, MK94]. Twister [MN98]. twisters [Nis00]. twisting [JS02]. Two [BFMW03, CMZ18, DN99, PPT14, PG14, RH19, SWL09, De 06, EHG92, KLF02, WPN98]. two-level [WPN98]. two-node [De 06]. Two-phase [SWL09]. Two-Sided [PPT14].

Two-stage [DN99, PG14, KLF02]. Two-tier [RH19]. Two-timescale [BFMW03]. type [KC10, SS08].

Ultrafast [LZW16]. UML [AK02]. unbounded [HLD07]. Uncertain [VXE^{+}22]. Uncertainty [JV23, PBAB^{+}11, UPB22, VWD22, XNB16, YX17, ZLZ20, MY08, NC06, PG14]. Uncertainty-aware [JV23]. Undo [CPQ17]. Unified [JB22a]. uniform [CL98, DX03, KSZ11, MN98]. Uniformization [BK20, DHK15]. uniforms [Hör94]. unifying [BCL91]. unimodal [Dev97, SZ99]. Union [AK18]. Universal [Bel05]. University [KFL00]. unknowns [vBBR03]. unmodified [KFL00]. Use [GK19, LALGSG^{+}00]. user [LDF91, SS03]. user-invoked [LDF91]. user-level [SS03]. users [LPPP13]. Using [AG16, ANSW23, CN98, DHK15, DHN22, ESZH21, FBCS22, GZWG18, GDB14, Nak14, RH19, RBDH97, SDLH12, TKS16, Van19, VLN^{+}19, VXE^{+}22, WCF23, WMC^{+}18, WZCJ22, AD92, BC93, BFMW03, BN03, BKV04, BN09, Cal07, Cal09, CPF99, DjWS19, Fis92, FG99, GAG14, HLC^{+}10, HBE95, JS02, LS92, Lor18, LLHL00, NCN^{+}22, Pac08, PF11, PRO13, RFA00, SJY03, WPS13, WPN98, WPW09, XGH12, Yau99]. Utilization [TKS16, AZK10]. utilizing [MM07]. UWB [AZK10]. UWB-based [AZK10].

validate [MPK06]. Validation [YN15, BÖ96, GDP14, PCT97]. validity [VSCL13]. Value [HHL14b]. Value-at-Risk [HHL14b]. valued [HVAPFY10]. VANET [NNB11]. Variability [GB19, ZK10]. Variable [HDM03, SU16, HLC12]. Variable-sample [HDM03]. Variable-Structure [SU16]. variables [DLW07, JKS07]. Variance [AHO93, BC13, CERT15, GAG14, LN18, Nak14, Owe03, SK23, Tuf97, AAAG06, AAGM10, CN98, CN12, Kaw10]. Variance-Reduction [Nak14, CN12]. variant [AK11]. Variate [QDZ21, DHL10, Dev97, Dev09, HL03]. variates [AHO93, Hör94, HD96, NS06, RR93, RJ04, YL96]. variation [KSZ11]. Variational [WCCY19]. varying [DLW07]. vector [Bel05, BN03, GH03, GH06, GH09, Nie94]. vectors [Emm98]. VEEs [LCT^{+}15]. vehicle [CFW99]. vehicles [OLAM08]. Verification [Ald18, Bee18, PCT97]. versatile [SSRT91]. Version [GGH^{+}23]. versus [WM99]. Very [SS05, Owe98]. via [ABGR01, AGMW17, And21, BHM11, CK08, CTF^{+}19, CG13, CM21, HKP21, HE12, HN07, KSW07, KFL00, Kim05, KWU22, LC01, LG03, Oso09, PHP^{+}15, PN03, SQ12, Di 23, DSR23, WXC^{+}23, WWH^{+}23, XYZ21, XNH10, ZLZ23]. View [LL20, CS92]. Virtual [BN22, HYJ21, JB22a, JB22b, JN15, KKTM17, LN18, LT14, YP15, CKP95, FH97, ZCLT04]. Virtual-Machine-Based [JN15]. Visual [GB19, GCB95]. visualization [Pac08]. Visualizing [HBE95]. VM [KSW03]. volumes [Pac08]. Volunteer [SALS18, Van18]. vs [UPB22].

WA [LCK11]. walks [HS12]. Warp [PQ17, AD92, DF97, DNRD96, LP91, LL91a, LDF91, PLM94, QC02]. Wasserstein [ZLZ23]. wave [Nut06]. wavelength [RRP00]. Wavelet [JFST24]. Waves [RLDH16]. Weak [ST15]. Web [KLF02, PBF^{+}00, RRW00, RFA00]. Web-based [RFA00, KLF02, PBF^{+}00, RRW00]. Weight [BV22, WJ22]. weighted [FG99, HN98]. well [Ent98]. well-known [Ent98]. Wildfire [TDR^{+}11, HN09, XGH12]. wimedia [AZK10]. wind [Pac08]. Wireless [KKTM17, SABF15, JZTB06, SJSM10, SF10]. WiseMove [BLG^{+}21]. WiseSim [BLG^{+}21]. within [DK22]. without [FK91]. WLAN [KKTM17]. Work [KV23, WYT^{+}20]. Work-stealing [WYT^{+}20]. Workflow [CAT22]. Workflows [CPRV23]. Workload [SALS18, Van18]. workloads [TFR07, WPN98]. Workshop [BNSS24, CY10, HHL14a]. world [CS92, ZJTB04]. worms [KHJ^{+}08, Nic08, RB08]. WPANs [AZK10]. Wrong [EH18, KH18]. WSNs [MRB^{+}18]. WWW [KFL00].

xMAS [ZL17]. xMAS-Based [ZL17]. Xorshift [Bre04, Mar03, PL05, Vig16].

YAWNS [DNRD96].

Zero [CERT15]. Zero-Variance [CERT15]. zone [KN01]. ZVA [RDSJ18].

References

Alexopoulos:2006:RBM

[AAAG06] Christos Alexopoulos, Sigrún Andradóttir, Nilay Tanik Argon, and David Goldsman. Replicated batch means variance estimators in the presence of an initial transient. ACM Transactions on Modeling and Computer Simulation, 16(4):317–328, October 2006. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Alexopoulos:2010:PFV

[AAGM10] Christos Alexopoulos, Claudia Antonini, David Goldsman, and Melike Meterelliyoz. Performance of folded variance estimators for simulation. ACM Transactions on Modeling and Computer Simulation, 20(3): 11:1–11:??, September 2010. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Aldini:2001:CQI

[ABGR01] Alessandro Aldini, Marco Bernardo, Roberto Gorrieri, and Marco Roccetti. Comparing the QoS of Internet audio mechanisms via formal methods. ACM Transactions on Modeling and Computer Simulation, 11(1):1–42, January 2001. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Ankenman:2015:SDE

[ACL15] Bruce E. Ankenman, Russell C. H. Cheng, and Susan M. Lewis. Screening for dispersion effects by sequential bifurcation. ACM Transactions on Modeling and Computer Simulation, 25(1):2:1–2:??, January 2015. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Ammar:1992:TWS

[AD92] Hany H. Ammar and Su Deng. Time warp simulation using time scale decomposition. ACM Transactions on Modeling and Computer Simulation, 2(2):158–177, April 1992. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Alexopoulos:2004:BB

[AG04] Christos Alexopoulos and David Goldsman. To batch or not to batch? ACM Transactions on Modeling and Computer Simulation, 14(1): 76–114, January 2004. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Awad:2007:TCL

[AG07] Hernan P. Awad and Peter W. Glynn. On the theoretical comparison of lowbias steady-state estimators. ACM Transactions on Modeling and Computer Simulation, 17(1):??, January 2007. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Andradottir:2016:CBM

[AG16] Sigrún Andradóttir and Peter W. Glynn. Computing Bayesian means using simulation. ACM Transactions on Modeling and Computer Simulation, 26(2):10:1–10:??, January 2016. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Andradottir:2007:PSP

[AGG^{+}07] Sigrún Andradóttir, Paul Glasserman, Peter W. Glynn, Philip Heidelberger, and Sandeep Juneja. Perwez Shahabuddin, 1962–2005: a professional appreciation. ACM Transactions on Modeling and Computer Simulation, 17(2): ??, April 2007. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Alexopoulos:2017:AEE

[AGMW17] Christos Alexopoulos, David Goldsman, Anup C. Mokashi, and James R. Wilson. Automated estimation of extreme steady-state quantiles via the maximum transformation. ACM Transactions on Modeling and Computer Simulation, 27(4):22:1–22:??, December 2017. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Asmussen:1992:SDI

[AGT92] Søren Asmussen, Peter W. Glynn, and Hermann Thorisson. Stationarity detection in the initial transient problem. ACM Transactions on Modeling and Computer Simulation, 2(2):130–157, April 1992. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Andradottir:1993:VRT

[AHO93] Sigrún Andradóttir, Daniel P. Heyman, and Teunis J. Ott. Variance reduction through smoothing and control variates for Markov chain simulations. ACM Transactions on Modeling and Computer Simulation, 3(3):167–189, July 1993. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Atkinson:2002:RUI

[AK02] Colin Atkinson and Thomas Kühne. Rearchitecting the UML infrastructure. ACM Transactions on Modeling and Computer Simulation, 12(4): 290–321, October 2002. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Amrein:2011:VIS

[AK11] Michael Amrein and Hans R.

Künsch. A variant of importance splitting for rare event estimation: Fixed number of successes. ACM Transactions on Modeling and Computer Simulation, 21(2):13:1– 13:??, February 2011. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Ahn:2018:ESE

[AK18] Dohyun Ahn and Kyoung-Kuk Kim. Efficient simulation for expectations over the union of half-spaces. ACM Transactions on Modeling and Computer Simulation, 28(3):23:1– 23:??, August 2018. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Aldini:2018:DVT

[Ald18] Alessandro Aldini. Design and verification of trusted collective adaptive systems. ACM Transactions on Modeling and Computer Simulation, 28(2):

9:1–9:??, April 2018. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Alexopoulos:2017:RCR

[Ale17] Christos Alexopoulos. Replicated computations results (RCR) report for “MNO– PQRS: Max Nonnegativity Ordering-Piecewise-Quadratic Rate Smoothing”. ACM Transactions on Modeling and Computer Simulation, 27(3): 18:1–18:??, September 2017. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). See [CS17].

Abate:2023:ISI

[AM23] Alessandro Abate and Andrea Marin. Introduction to the special issue on QEST 2021. ACM Transactions on Modeling and Computer Simulation, 33(4):13:1–13:??, October 2023. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL

https://dl.acm.org/doi/
10.1145/3631707.

Akbari-Moghaddam:2023:SSE

[AMD23] Maryam Akbari-Moghaddam and Douglas G. Down. SEH: Size estimate hedging scheduling of queues. ACM Transactions on Modeling and Computer Simulation, 33(4): 14:1–14:??, October 2023. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3580491.

Andradottir:1999:ACR

[And99] Sigrún Andradóttir. Accelerating the convergence of random search methods for discrete stochastic optimization. ACM Transactions on Modeling and Computer Simulation, 9(4):349–380, October 1999. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Andradottir:2006:SOC

[And06] Sigrún Andradóttir. Simulation optimization with countably infinite feasible regions: Efficiency and convergence. ACM Transactions on Modeling and Computer Simulation, 16(4):357–374, October 2006. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Andelfinger:2021:RCR

[And21] Philipp Andelfinger. Replicated computational results (RCR) report for “A Practical Approach to Subset Selection for Multi-Objective Optimization via Simulation”. ACM Transactions on Modeling and Computer Simulation, 31 (4):21:1–21:2, October 2021. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3453987.

Andelfinger:2022:TDA

[And22]  Philipp Andelfinger.  To-

wards differentiable agentbased simulation. ACM Transactions on Modeling and Computer Simulation, 32(4): 27:1–27:??, October 2022. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3565810.

Anonymous:2018:GET

[Ano18] Anonymous. Guest editorial for the TOMACS special issue on the Principles of Advanced Discrete Simulation (PADS). ACM Transactions on Modeling and Computer Simulation, 28(4):25:1–25:??, October 2018. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Anonymous:2021:RCR

[Ano21] Anonymous. Replication of computational results report for “Doping Tests for Cyber-Physical Systems”. ACM Transactions on Modeling and Computer Simulation, 31(3):17:1–17:2, July 2021. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3459667.

Avci:2023:UCC

[ANSW23] Harun Avci, Barry L. Nelson, Eunhye Song, and Andreas Wächter. Using cache or credit for parallel ranking and selection. ACM Transactions on Modeling and Computer Simulation, 33(4): 12:1–12:??, October 2023. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618299.

Andradottir:1995:TSP

[AO95] Sigrún Andradóttir and Teunis J. Ott. Time-segmentation parallel simulation of networks of queues with loss or communication blocking. ACM Transactions on Modeling and Computer Simulation, 5(4): 269–305, October 1995. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Agha:2018:SSM

[AP18] Gul Agha and Karl Palmskog. A survey of statistical model checking. ACM Transactions on Modeling and Computer Simulation, 28(1): 6:1–6:??, January 2018. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Andrew:2010:SFG

[AQVA10] Lachlan L. H. Andrew, Guoqi Qian, and Felisa J. Vázquez-Abad. Setwise and filtered Gibbs samplers for teletraffic analysis. ACM Transactions on Modeling and Computer Simulation, 20(2):7:1–7:??, April 2010. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Andelfinger:2020:FPS

[AXE^{+}20] Philipp Andelfinger, Yadong Xu, David Eckhoff, Wentong Cai, and Alois Knoll. Fidelity and performance of state fast-forwarding in microscopic traffic simulations. ACM Transactions on Modeling and Computer Simulation, 30(2):10:1–10:26, April 2020. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:

                                        //dl.acm.org/doi/abs/10.
                                        1145/3366019.

Al-Zubi:2010:CLD

[AZK10] Raed Al-Zubi and Marwan Krunz. Cross-layer design for efficient resource utilization in wimedia UWB-based WPANs. ACM Transactions on Modeling and Computer Simulation, 21(1):8:1–8:??, December 2010. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Al-Zubaidy:2010:OSH

[AZLT10] Hussein Al-Zubaidy, Ioannis Lambadaris, and Jerome Talim. Optimal scheduling in high-speed downlink packet access networks. ACM Transactions on Modeling and Computer Simulation, 21(1):3:1– 3:??, December 2010. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Balci:1997:GES

[Bal97] Osman Balci. Guest editorial— Simulation for training: foundations and techniques. ACM Transactions on Modeling and Computer Simulation, 7(3): 291–292, July 1997. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Balci:2001:MCM

[Bal01] Osman Balci. A methodology for certification of modeling and simulation applications. ACM Transactions on Modeling and Computer Simulation, 11(4):352–377, October 2001. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Barros:1997:MFD

[Bar97] Fernando J. Barros. Modeling formalisms for dynamic structure systems. ACM Transactions on Modeling and Computer Simulation, 7(4):501– 515, October 1997. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Barros:2003:DSM

[Bar03] Fernando J. Barros. Dynamic structure multiparadigm modeling and simulation. ACM Transactions on Modeling and Computer Simulation, 13(3):

259–275, July 2003. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Bonarini:1994:QSA

[BB94] Andrea Bonarini and Gianluca Bontempi. A qualitative simulation approach for fuzzy dynamical models. ACM Transactions on Modeling and Computer Simulation, 4(4): 285–313, October 1994. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Barbe:1999:SEF

[BB99] Philippe Barbe and Michel Broniatowski. Simulation in exponential families. ACM Transactions on Modeling and Computer Simulation, 9(3): 203–223, July 1999. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Bortolussi:2019:ISI

[BB19] Luca Bortolussi and Nathalie Bertrand. Introduction to the special issue on Qest 2017. ACM Transactions on Modeling and Computer Simulation, 29(4):19:1–19:??, December 2019. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3363784.

Bremer:2022:PAS

[BBCD22] Maximilian Bremer, John Bachan, Cy Chan, and Clint Dawson. Performance analysis of speculative parallel adaptive local timestepping for conservation laws. ACM Transactions on Modeling and Computer Simulation, 32(4): 26:1–26:??, October 2022. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3545996.

Bae:2016:EFA

[BBMK16] Jang Won Bae, Sang Won Bae, Il-Chul Moon, and Tag Gon Kim. Efficient flattening algorithm for hierarchical and dynamic structure discrete event models. ACM Transactions on Modeling and Computer Simulation, 26(4): 25:1–25:??, May 2016. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Baccelli:1993:PSS

[BC93] François Baccelli and Miguel Canales. Parallel simulation of stochastic Petri nets using recurrence equations. ACM Transactions on Modeling and Computer Simulation, 3(1): 20–41, January 1993. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Broniatowski:2013:SVE

[BC13] Michel Broniatowski and Virgile Caron. Small variance estimators for rare event probabilities. ACM Transactions on Modeling and Computer Simulation, 23(1):7:1–7:??, January 2013. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Bisset:2014:IIH

[BCD^{+}14] Keith R. Bisset, Jiangzhuo Chen, Suruchi Deodhar, Xizhou Feng, Yifei Ma, and Madhav V. Marathe. Indemics: an interactive high-performance computing framework for data-intensive epidemic modeling. ACM Transactions on Modeling and Computer Simulation, 24(1): 4:1–4:??, January 2014. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Bagrodia:1991:UFD

[BCL91] R. Bagrodia, K. M. Chandy, and Wen Toh Liao. A unifying framework for distributed simulation. ACM Transactions on Modeling and Computer Simulation, 1(4):348–385, October 1991. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Bassiouni:1997:PRA

[BCL^{+}97] Mostafa A. Bassiouni, Ming-Hsing Chiu, Margaret Loper, Michael Garnsey, and Jim Williams. Performance and reliability analysis of relevance filtering for scalable distributed interactive simulation. ACM Transactions on Modeling and Computer Simulation, 7(3):293–331, July 1997. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Bisewski:2018:CTD

[BCM18] Krzysztof Bisewski, Daan Crommelin, and Michel Mandjes. Controlling the time discretization bias for the supremum of Brownian motion. ACM Transactions on Modeling and Computer Simulation, 28(3):24:1–24:??, August 2018. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Broadie:2014:MSA

[BCZ14] Mark Broadie, Deniz M. Cicek, and Assaf Zeevi. Multidimensional stochastic approximation: Adaptive algorithms and applications. ACM Transactions on Modeling and Computer Simulation, 24(1): 6:1–6:??, January 2014. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Bocciarelli:2020:MRS

[BDGP20] Paolo Bocciarelli, Andrea D’Ambrogio, Andrea Giglio, and Emiliano Paglia. Modeling resources to simulate business process reliability. ACM Transactions on Modeling and Computer Simulation, 30(3): 14:1–14:25, July 2020. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3381453.

Biewer:2021:DTC

[BDH21] Sebastian Biewer, Pedro R. D’argenio, and Holger Hermanns. Doping tests for cyber-physical systems. ACM Transactions on Modeling and Computer Simulation, 31(3):16:1–16:27, July 2021. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3449354.

Baier:2019:MPO

[BDK^{+}19] Christel Baier, Clemens Dubslaff, Lúbos Korenciak, Antonín Kucera, and Vojtech Rehák. Mean-payoff optimization in continuous-time Markov chains with parametric alarms. ACM Transactions on Modeling and Computer Simulation, 29(4):28:1– 28:??, December 2019. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Beek:2018:RCR

[Bee18] Maurice H. Ter Beek. Replicated Computations Results (RCR) report for “Design and Verification of Trusted Collective Adaptive Systems”. ACM Transactions on Modeling and Computer Simulation, 28(2): 10:1–10:??, April 2018. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Beliakov:2005:UNR

[Bel05]  Gleb Beliakov.  Universal

nonuniform random vector generator based on acceptancerejection. ACM Transactions on Modeling and Computer Simulation, 15(3):205–232, July 2005. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Bhatnagar:2003:TTS

[BFMW03] Shalabh Bhatnagar, Michael C.

Fu, Steven I. Marcus, and I-Jeng Wang. Two-timescale simultaneous perturbation stochastic approximation using deterministic perturbation sequences. ACM Transactions on Modeling and Computer Simulation, 13(2):180–209, April 2003. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Bratley:1992:ITL

[BFN92] Paul Bratley, Bennett L. Fox, and Harald Niederreiter. Implementation and tests of lowdiscrepancy sequences. ACM Transactions on Modeling and Computer Simulation, 2(3): 195–213, July 1992. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Bremaud:1993:DLR

[BG93] P. Brémaud and W.-B. Gong. Derivatives of likelihood ratios and smoothed perturbation analysis for the routing problem. ACM Transactions on Modeling and Computer Simulation, 3(2):134–161, April 1993. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Blanchet:2012:LIS

[BGL12] Jose Blanchet, Peter Glynn, and Kevin Leder. On Lyapunov inequalities and subsolutions for efficient importance sampling. ACM Transactions on Modeling and Computer Simulation, 22(3):13:1– 13:??, August 2012. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Bhatnagar:2005:AMT

[Bha05] Shalabh Bhatnagar. Adaptive multivariate three-timescale stochastic approximation algorithms for simulation based optimization. ACM Transactions on Modeling and Computer Simulation, 15(1):74– 107, January 2005. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Bhatnagar:2007:ANB

[Bha07] Shalabh Bhatnagar. Adaptive Newton-based multivariate smoothed functional algorithms for simulation optimization. ACM Transactions on Modeling and Computer Simulation, 18(1):2:1– 2:35, December 2007. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Begum:2010:MIB

[BHG10] Shamim Begum, Ahmed Helmy, and Sandeep Gupta. Modeling the interactions between MAC and higher layer: a systematic approach to generate high-level scenarios from MAC-layer scenarios. ACM Transactions on Modeling and Computer Simulation, 21(1): 7:1–7:??, December 2010. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Butkova:2021:MAM

[BHH21] Yuliya Butkova, Arnd Hartmanns, and Holger Hermanns. A modest approach to Markov automata. ACM Transactions on Modeling and Computer Simulation, 31(3):14:1–14:34, July 2021. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL

https://dl.acm.org/doi/
10.1145/3449355.

Blanchet:2013:RES

[BHL13] Jose Blanchet, Henrik Hult, and Kevin Leder. Rareevent simulation for stochastic recurrence equations with heavy-tailed innovations. ACM Transactions on Modeling and Computer Simulation, 23(4): 22:1–22:??, October 2013. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Bai:2022:RES

[BHLZ22] Yuanlu Bai, Zhiyuan Huang, Henry Lam, and Ding Zhao. Rare-event simulation for neural network and random forest predictors. ACM Transactions on Modeling and Computer Simulation, 32(3):18:1–18:33, July 2022. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3519385.

Bhatnagar:2011:SAA

[BHM11] Shalabh Bhatnagar, N. Hemachandra, and Vivek Kumar Mishra. Stochastic approximation algorithms for constrained optimization via simulation. ACM Transactions on Modeling and Computer Simulation, 21(3): 15:1–15:??, March 2011. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Batur:2010:FFS

[BK10] Demet Batur and Seong-Hee Kim. Finding feasible systems in the presence of constraints on multiple performance measures. ACM Transactions on Modeling and Computer Simulation, 20(3):13:1– 13:??, September 2010. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Burak:2020:ICB

[BK20] Maciej Rafal Burak and Przemyslaw Korytkowski. Inhomogeneous CTMC birth-and-death models solved by uniformization with steady-state detection. ACM Transactions on Modeling and Computer Simulation, 30(3):18:1– 18:18, July 2020. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3373758.

Bhatnagar:2009:OPT

[BKM09] Shalabh Bhatnagar, Karmeshu, and Vivek Kumar Mishra. Optimal parameter trajectory estimation in parameterized SDEs: an algorithmic procedure. ACM Transactions on Modeling and Computer Simulation, 19(2):8:1–8:??, March 2009. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Bobeanu:2004:MDE

[BKV04] Carmen-Veronica Bobeanu, Eugene J. H. Kerckhoffs, and Hendrik Van Landeghem. Modeling of discrete event systems: a holistic and incremental approach using Petri nets. ACM Transactions on Modeling and Computer Simulation, 14(4):389–423, October 2004. CODEN ATMCEZ.

ISSN 1049-3301 (print), 1558-1195 (electronic).

Barton:2002:MSS

[BL02] Paul I. Barton and Cha Kun Lee. Modeling, simulation, sensitivity analysis, and optimization of hybrid systems. ACM Transactions on Modeling and Computer Simulation, 12(4):256–289, October 2002. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Blanchet:2011:ERE

[BL11] Jose Blanchet and Chenxin Li. Efficient rare event simulation for heavy-tailed compound sums. ACM Transactions on Modeling and Computer Simulation, 21(2):9:1– 9:??, February 2011. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Balakrishnan:2021:TRL

[BLG^{+}21] Aravind Balakrishnan, Jaeyoung Lee, Ashish Gaurav, Krzysztof Czarnecki, and Sean Sedwards. Transfer reinforcement learning for autonomous driving: From WiseMove to WiseSim. ACM Transactions on Modeling and Computer Simulation, 31(3):15:1–15:26, July 2021. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3449356.

Botev:2016:SNR

[BLST16] Zdravko I. Botev, Pierre L’Écuyer, Richard Simard, and Bruno Tuffin. Static network reliability estimation under the Marshall–Olkin copula. ACM Transactions on Modeling and Computer Simulation, 26(2):14:1–14:??, January 2016. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Bhimani:2019:NPM

[BMLY19] Janki Bhimani, Ningfang Mi, Miriam Leeser, and Zhengyu Yang. New performance modeling methods for parallel data processing applications. ACM Transactions on Modeling and Computer Simulation, 29(3):15:1–15:??, July 2019. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3309684.

Biller:2003:MGM

[BN03] Bahar Biller and Barry L. Nelson. Modeling and generating multivariate time-series input processes using a vector autoregressive technique. ACM Transactions on Modeling and Computer Simulation, 13(3):211–237, July 2003. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Brandao:2009:ANS

[BN09] Rita Marques Brandão and Acácio M. O. Porta Nova. Analysis of nonstationary stochastic simulations using classical time-series models. ACM Transactions on Modeling and Computer Simulation, 19(2):9:1–9:??, March 2009. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Babu:2022:MPV

[BN22] Vignesh Babu and David Nicol. Mechanisms for precise virtual time advancement in network emulation. ACM Transactions on Modeling and Computer Simulation, 32(2):9:1–9:26, April 2022. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478867.

Barton:2024:ISI

[BNSS24] Russell R. Barton, Marvin K. Nakayama, Uday V. Shanbhag, and Eunhye Song. Introduction to the special issue for INFORMS Simulation Society (I-Sim) Workshop, 2021. ACM Transactions on Modeling and Computer Simulation, 34(2):5:1– 5:??, April 2024. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3655711.

Birta:1996:KBA

[BÖ96] Louis G. Birta and F. Nur Özmizrak. A knowledge-based approach for the validation of simulation models: the foundation. ACM Transactions on Modeling and Computer Simulation, 6(1):76–98, January 1996. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Bailey:1994:EMO

[BP94] Mary L. Bailey and Michael A.

Pagels. Empirical measurements of overheads in conservative asynchronous simulations. ACM Transactions on Modeling and Computer Simulation, 4(4):350–367, October 1994. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Brent:2004:NMX

[Bre04] Richard P. Brent. Note on Marsaglia’s xorshift random number generators. Journal of Statistical Software, 11 (5):1–5, 2004. CODEN JS-SOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/
counter.php?id=101&url=
v11/i05/v11i05.pdf&ct=1.

See [Mar03, PL05, Vig16]. This article shows the equivalence of xorshift generators and the well-understood linear feedback shift register generators.

Bandini:2016:GEE

[BSV16] Stefania Bandini, Georgios Ch.

Sirakoulis, and Giuseppe Vizzari. Guests editors’ editorial note on special issue of advances in cellular automata modeling. ACM Transactions on Modeling and Computer Simulation, 26(3):17:1– 17:??, February 2016. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Buchholz:1998:NAC

[Buc98] Peter Buchholz. A new approach combining simulation and randomization for the analysis of large continuous time Markov chains. ACM Transactions on Modeling and Computer Simulation, 8(2): 194–222, April 1998. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Blackman:2022:NTH

[BV22] David Blackman and Sebastiano Vigna. A new test for Hamming-weight dependencies. ACM Transactions on Modeling and Computer Simulation, 32(3):19:1–19:13, July 2022. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3527582.

Blanchet:2015:ESS

[BW15] Jose Blanchet and Aya Wallwater. Exact sampling of stationary and time-reversed queues. ACM Transactions on Modeling and Computer Simulation, 25(4):26:1– 26:??, November 2015. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Calvin:2007:SOA

[Cal07]  James M. Calvin.  Sim-

ulation output analysis using integrated paths. ACM Transactions on Modeling and Computer Simulation, 17(3): 13:1–13:??, July 2007. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Calvin:2009:SOA

[Cal09] James M. Calvin. Simulation output analysis using integrated paths II: Low bias estimators. ACM Transactions on Modeling and Computer Simulation, 19(3):11:1–11:??, June 2009. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Chen:2012:ECR

[CAN12] Xi Chen, Bruce E. Ankenman, and Barry L. Nelson. The effects of Common Random Numbers on stochastic kriging metamodels. ACM Transactions on Modeling and Computer Simulation, 22(2): 7:1–7:20, March 2012. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Chaudhry:2022:WAC

[CAT22] Nauman Riaz Chaudhry, Anastasia Anagnostou, and Simon J. E. Taylor. A workflow architecture for cloudbased distributed simulation. ACM Transactions on Modeling and Computer Simulation, 32(2):15:1–15:26, April 2022. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3503510.

Cuneo:2024:DRM

[CB24] Braxton Cuneo and Mike Bailey. Divergence reduction in Monte Carlo neutron transport with on-GPU asynchronous scheduling. ACM Transactions on Modeling and Computer Simulation, 34 (1):2:1–2:??, January 2024. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3626957.

Casale:2016:QOB

[CDS16] Giuliano Casale, Vittoria De Nitto Personé, and Evgenia Smirni. QRF: an optimization-based framework for evaluating complex stochastic networks. ACM Transactions on Modeling and Computer Simulation, 26(3): 15:1–15:??, February 2016. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Cancela:2015:BAZ

[CERT15] Hector Cancela, Mohamed El Khadiri, Gerardo Rubino, and Bruno Tuffin. Balanced and approximate zero-variance recursive estimators for the network reliability problem. ACM Transactions on Modeling and Computer Simulation, 25(1):5:1–5:??, January 2015. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Chertov:2011:FDM

[CF11] Roman Chertov and Sonia Fahmy. Forwarding devices: From measurements to simulations. ACM Transactions on Modeling and Computer Simulation, 21(2):12:1– 12:??, February 2011. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Chen:2012:SLP

[CFL12] Zisheng Chen, Liming Feng, and Xiong Lin. Simulating Lévy processes from their characteristic functions and financial applications. ACM Transactions on Modeling and Computer Simulation, 22(3): 14:1–14:??, August 2012. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Chertov:2008:FNS

[CFS08] Roman Chertov, Sonia Fahmy, and Ness B. Shroff. Fidelity of network simulation and emulation: a case study of TCPtargeted denial of service attacks. ACM Transactions on Modeling and Computer Simulation, 19(1):4:1–4:??, December 2008. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Chen:1999:RTS

[CFW99] Jim X. Chen, Xiadong Fu, and J. Wegman. Real-time simulation of dust behavior generated by a fast traveling vehicle. ACM Transactions on Modeling and Computer Simulation, 9(2):81–104, April 1999. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Chidester:2002:PSC

[CG02] Matthew Chidester and Alan George. Parallel simulation of chip-multiprocessor architectures. ACM Transactions on Modeling and Computer Simulation, 12(3):176–200, July 2002. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Chia:2013:LTS

[CG13] Yen Lin Chia and Peter W. Glynn. Limit theorems for simulation-based optimization via random search. ACM Transactions on Modeling and Computer Simulation, 23(3): 16:1–16:??, July 2013. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Calvin:2006:SRM

[CGN06] James M. Calvin, Peter W.

Glynn, and Marvin K. Nakayama. The semi-regenerative method of simulation output analysis. ACM Transactions on Modeling and Computer Simulation, 16(3):280–315, July 2006. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Cheng:2004:CCI

[CH04] R. C. H. Cheng and W. Holland. Calculation of confidence intervals for simulation output. ACM Transactions on Modeling and Computer Simulation, 14(4):344–362, October 2004. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Cen:2023:NGN

[CH23] Wang Cen and Peter J.

Haas. NIM: Generative neural networks for automated modeling and generation of simulation inputs. ACM Transactions on Modeling and Computer Simulation, 33(3):10:1–10:??, July 2023. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592790.

Cardenas:2022:DMP

[CHA^{+}22] Román Cárdenas, Kevin Henares, Patricia Arroba, José L. Risco-Martín, and Gabriel A. Wainer. The DE-VStone metric: Performance analysis of DEVS simulation engines. ACM Transactions on Modeling and Computer Simulation, 32(3):21:1–21:20, July 2022. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL

https://dl.acm.org/doi/
10.1145/3543849.

Cheng:2013:FSM

[Che13] Russell C. H. Cheng. Fitting statistical models of random search in simulation studies. ACM Transactions on Modeling and Computer Simulation, 23(3):15:1–15:??, July 2013. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Coe:1998:TNH

[CHIW98] P. S. Coe, F. W. Howell, R. N. Ibbett, and L. M. Williams. Technical note: a hierarchical computer architecture design and simulation environment. ACM Transactions on Modeling and Computer Simulation, 8(4):431–446, October 1998. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Chang:1995:FSP

[CHS95] Cheng-Shang Chang, Philip Heidelberger, and Perwez Shahabuddin. Fast simulation of packet loss rates in a shared buffer communications switch. ACM Transactions on Modeling and Computer Simulation, 5(4):306–325, October 1995. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Carl:2008:LST

[CK08] Glenn Carl and George Kesidis. Large-scale testing of the Internet’s Border Gateway Protocol (BGP) via topological scale-down. ACM Transactions on Modeling and Computer Simulation, 18(3): 11:1–11:??, July 2008. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Chen:2014:SKB

[CK14] Xi Chen and Kyoung-Kuk Kim. Stochastic kriging with biased sample estimates. ACM Transactions on Modeling and Computer Simulation, 24(2): 8:1–8:??, February 2014. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Choi:2013:PAC

[CKL^{+}13] Byoung K. Choi, Donghun Kang, Taesik Lee, Arwa A. Jamjoom, and Maysoon F. Abulkhair. Parameterized activity cycle diagram and its application. ACM Transactions on Modeling and Computer Simulation, 23(4):24:1– 24:??, October 2013. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Chan:2023:SID

[CKM23] Cy Chan, Anu Kuncheria, and Jane Macfarlane. Simulating the impact of dynamic rerouting on metropolitanscale traffic systems. ACM Transactions on Modeling and Computer Simulation, 33 (1–2):7:1–7:??, April 2023. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3579842.

Cremer:1995:HFB

[CKP95] James Cremer, Joseph Kearney, and Yiannis Papelis. HCSM: a framework for behavior and scenario control in virtual environments. ACM Transactions on Modeling and Computer Simulation, 5(3): 242–267, July 1995. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Couture:1998:GEI

[CL98] Raymond Couture and Pierre L’Ecuyer. Guest editors’ introduction: special issue on uniform random number generation. ACM Transactions on Modeling and Computer Simulation, 8(1):1–2, January 1998. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Choquet:1999:BCI

[CLL99] Denis Choquet, Pierre L’Ecuyer, and Christian Léger. Bootstrap confidence intervals for ratios of expectations. ACM Transactions on Modeling and Computer Simulation, 9(4): 326–348, October 1999. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Currie:2021:PAS

[CM21] Christine S. M. Currie and Thomas Monks. A practical approach to subset selection for multi-objective optimization via simulation. ACM Transactions on Modeling and Computer Simulation, 31(4): 20:1–20:15, October 2021. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3462187.

Catania:2016:CAN

[CMM^{+}16] Vincenzo Catania, Andrea Mineo, Salvatore Monteleone, Maurizio Palesi, and Davide Patti. Cycle-accurate network on chip simulation with Noxim. ACM Transactions on Modeling and Computer Simulation, 27(1):4:1–4:??, November 2016. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Cahen:2018:ELD

[CMZ18] Ewan Jacov Cahen, Michel Mandjes, and Bert Zwart. Estimating large delay probabilities in two correlated queues. ACM Transactions on Modeling and Computer Simulation, 28(1):2:1–2:??, January 2018. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Calvin:1998:UPR

[CN98] James M. Calvin and Marvin K. Nakayama. Using permutations in regenerative simulations to reduce variance. ACM Transactions on Modeling and Computer Simulation, 8(2):153–193, April 1998. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Chu:2012:CIQ

[CN12] Fang Chu and Marvin K. Nakayama. Confidence intervals for quantiles when applying variance-reduction techniques. ACM Transactions on Modeling and Computer Simulation, 22(2):10:1–10:25, March 2012. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Calvin:2015:RRE

[CN15] James M. Calvin and Marvin K. Nakayama. Resampled regenerative estimators. ACM Transactions on Modeling and Computer Simulation, 25(4):23:1–23:??, November 2015. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Cicirelli:2016:ESC

[CN16] Franco Cicirelli and Libero Nigro. Exploiting social capabilities in the minority game. ACM Transactions on Modeling and Computer Simulation, 27(1):6:1–6:??, November 2016. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Chen:1998:TPT

[CO98] Wu-Lin Chen and Colm Art O’Cinneide. Towards a polynomial-time randomized algorithm for closed productform networks. ACM Transactions on Modeling and Computer Simulation, 8(3): 227–253, July 1998. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Connor:2020:OPS

[Con20] Stephen Connor. Omnithermal perfect simulation for multi-server queues. ACM Transactions on Modeling and Computer Simulation, 30(1): 6:1–6:15, February 2020. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3361743.

Carothers:1999:EOP

[CPF99] Christopher D. Carothers, Kalyan S. Perumalla, and Richard M. Fujimoto. Efficient optimistic parallel simulations using reverse computation.

ACM Transactions on Modeling and Computer Simulation, 9(3):224–253, July 1999. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Cingolani:2017:TMU

[CPQ17] Davide Cingolani, Alessandro Pellegrini, and Francesco Quaglia. Transparently mixing undo logs and software reversibility for state recovery in optimistic PDES. ACM Transactions on Modeling and Computer Simulation, 27(2): 11:1–11:??, July 2017. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Carnevali:2023:CSA

[CPRV23] Laura Carnevali, Marco Paolieri, Riccardo Reali, and Enrico Vicario. Compositional safe approximation of response time probability density function of complex workflows. ACM Transactions on Modeling and Computer Simulation, 33(4): 16:1–16:??, October 2023. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3591205.

Cota:1992:MPI

[CS92] Bruce A. Cota and Robert G. Sargent. A modification of the process interaction world view. ACM Transactions on Modeling and Computer Simulation, 2(2):109–129, April 1992. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Chan:2008:MPM

[CS08] Wai Kin Victor Chan and Lee W. Schruben. Mathematical programming models of closed tandem queueing networks. ACM Transactions on Modeling and Computer Simulation, 19(1):3:1–3:??, December 2008. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Chen:2017:MPM

[CS17] Huifen Chen and Bruce W.

Schmeiser. MNO–PQRS: Max nonnegativity orderingpiecewise-quadratic rate smoothing. ACM Transactions on Modeling and Computer Simulation, 27(3):18:1–18:??, September 2017. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). See replication report [Ale17].

Clary:2010:PDR

[CSK10] Daniel W. Mc Clary, Violet R. Syrotiuk, and Murat Kulahci. Profile-driven regression for modeling and runtime optimization of mobile networks. ACM Transactions on Modeling and Computer Simulation, 20(3):17:1– 17:??, September 2010. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Chennupati:2021:MLE

[CSRE21] Gopinath Chennupati, Nandakishore Santhi, Phill Romero, and Stephan Eidenbenz. Machine learning-enabled scalable performance prediction of scientific codes. ACM Transactions on Modeling and Computer Simulation, 31 (2):11:1–11:28, April 2021. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450264.

Chen:2005:AHB

[CTC^{+}05] Dan Chen, Stephen J. Turner, Wentong Cai, Boon Ping Gan, and Malcolm Yoke Hean Low. Algorithms for HLA-based distributed simulation cloning. ACM Transactions on Modeling and Computer Simulation, 15(4):316–345, October 2005. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Cheh:2019:DDM

[CTF^{+}19] Carmen Cheh, Uttam Thakore, Ahmed Fawaz, Binbin Chen, William G. Temple, and William H. Sanders. Datadriven model-based detection of malicious insiders via physical access logs. ACM Transactions on Modeling and Computer Simulation, 29(4): 26:1–26:??, December 2019. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3309540.

Celik:2013:DFD

[ÇTI13] Turgay Çeli̇k, Bedi̇r Teki̇nerdogan, and Kayhan M. Imre. Deriving feasible deployment alternatives for parallel and distributed simulation systems. ACM Transactions on Modeling and Computer Simulation, 23(3):18:1–18:??, July 2013. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Cai:2005:ATM

[CTLZ05] Wentong Cai, Stephen J.

Turner, Bu-Sung Lee, and Junlan Zhou. An alternative time management mechanism for distributed simulations. ACM Transactions on Modeling and Computer Simulation, 15(2):109–137, April 2005. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Cetinkaya:2015:MCD

[ÇVS15] Deniz Çetinkaya, Alexander Verbraeck, and Mamadou D. Seck. Model continuity in discrete event simulation: a framework for model-driven development of simulation models. ACM Transactions on Modeling and Computer Simulation, 25(3):17:1–17:??, April 2015. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Cakmak:2024:CRS

[CWGZ24] Sait Cakmak, Yuhao Wang, Siyang Gao, and Enlu Zhou. Contextual ranking and selection with Gaussian processes and optimal computing budget allocation. ACM Transactions on Modeling and Computer Simulation, 34(2):8:1–8:??, April 2024. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3633456.

Chick:2010:GEI

[CY10] Stephen E. Chick and Enver Yücesan. Guest editors’ introduction to special issue on the first INFORMS simulation society research workshop. ACM Transactions on Modeling and Computer Simulation, 20(1): 1:1–1:3, January 2010. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Ding:2022:GFS

[DC22] Kailin Ding and Zhenyu Cui. A general framework to simulate diffusions with discontinuous coefficients and local times. ACM Transactions on Modeling and Computer Simulation, 32(4):22:1– 22:??, October 2022. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3559541.

DeBoer:2006:ASI

[De 06] Pieter Tjerk De Boer. Analysis of state-independent importance-sampling measures for the two-node tandem queue. ACM Transactions on Modeling and Computer Simulation, 16(3):225–250, July 2006. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Deng:2005:EPM

[Den05]  Lih-Yuan Deng.  Efficient

and portable multiple recursive generators of large order. ACM Transactions on Modeling and Computer Simulation, 15(1):1–13, January 2005. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Devroye:1997:RVG

[Dev97] Luc Devroye. Random variate generation for multivariate unimodal densities. ACM Transactions on Modeling and Computer Simulation, 7(4): 447–477, October 1997. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Devroye:2009:RVG

[Dev09] Luc Devroye. Random variate generation for exponentially and polynomially tilted stable distributions. ACM Transactions on Modeling and Computer Simulation, 19(4): 18:1–18:??, October 2009. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Das:1997:AMM

[DF97] Samir R. Das and Richard M. Fujimoto. Adaptive memory management and optimism control in time warp. ACM Transactions on Modeling and Computer Simulation, 7(2):239–271, April 1997. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Devetsikiotis:2010:GEI

[DG10] Michael Devetsikiotis and Fabrizio Granelli. Guest editors’ introduction: Special issue on modeling and simulation of cross-layer interactions in communication networks. ACM Transactions on Modeling and Computer Simulation, 21(1):1:1–1:??, December 2010. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Dannenberg:2015:CCR

[DHK15] Frits Dannenberg, Ernst Moritz Hahn, and Marta Kwiatkowska. Computing cumulative rewards using fast adaptive uniformization. ACM Transactions on Modeling and Computer Simulation, 25(2):9:1– 9:??, February 2015. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Derflinger:2010:RVG

[DHL10] Gerhard Derflinger, Wolfgang Hörmann, and Josef Leydold. Random variate generation by numerical inversion when only the density is known. ACM Transactions on Modeling and Computer Simulation, 20(4): 18:1–18:??, October 2010. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Donohue:1993:SED

[DHM93] Joan M. Donohue, Ernest C.

Houck, and Raymond H. Myers. A sequential experimental design procedure for the estimation of first- and secondorder simulation metamodels. ACM Transactions on Modeling and Computer Simulation, 3(3):190–224, July 1993. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Djehiche:2022:ISS

[DHN22] Boualem Djehiche, Henrik Hult, and Pierre Nyquist. Importance sampling for a simple Markovian intensity model using subsolutions. ACM Transactions on Modeling and Computer Simulation, 32(2):14:1–14:25, April 2022. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3502432.

Sanzo:2023:RCR

[Di 23] Pierangelo Di Sanzo. Replication of computational results report for “Automatic Reuse, Adaption, and Execution of Simulation Experiments via Provenance Patterns”. ACM Transactions on Modeling and Computer Simulation, 33 (1–2):5:1–5:??, April 2023. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3577007.

Devroye:2011:DCM

[DJ11] Luc Devroye and Lancelot F. James. The double CFTP method. ACM Transactions on Modeling and Computer Simulation, 21(2):10:1– 10:??, February 2011. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

DelMoral:2017:MSM

[DJLZ17] Pierre Del Moral, Ajay Jasra, Kody J. H. Law, and Yan Zhou. Multilevel sequential Monte Carlo samplers for normalizing constants. ACM Transactions on Modeling and Computer Simulation, 27(3): 20:1–20:??, September 2017. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Dwarkadas:1994:EDS

[DJS94] S. Dwarkadas, J. R. Jump, and J. B. Sinclair. Executiondriven simulation of multiprocessors: address and timing analysis. ACM Transactions on Modeling and Computer Simulation, 4(4):314–338, October 1994. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Dupuis:2019:ISU

[DjWS19] Paul Dupuis, Guo jhen Wu, and Michael Snarski. Infinite swapping using IID samples. ACM Transactions on Modeling and Computer Simulation, 29(3):13:1–13:??, July 2019. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3317605.

Divis:2022:RNS

[DK22] Roman Divis and Antonín Kavicka. Reflective nested simulations supporting optimizations within sequential railway traffic simulators. ACM Transactions on Modeling and Computer Simulation, 32(1):1:1–1:34, January 2022. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3467965.

Dimitropoulos:2009:GAM

[DKVR09] Xenofontas Dimitropoulos, Dmitri Krioukov, Amin Vahdat, and George Riley. Graph annotations in modeling complex network topologies. ACM Transactions on Modeling and Computer Simulation, 19(4): 17:1–17:??, October 2009. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Dassios:2020:EST

[DLQ20] Angelos Dassios, Jia Wei Lim, and Yan Qu. Exact simulation of a truncated Lévy subordinator. ACM Transactions on Modeling and Computer Simulation, 30(3): 17:1–17:17, July 2020. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3368088.

Dupuis:2007:ISS

[DLW07] Paul Dupuis, Kevin Leder, and Hui Wang. Importance sampling for sums of random variables with regularly varying tails. ACM Transactions on Modeling and Computer Simulation, 17(3):14:1–14:21, July 2007. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Dieker:2006:FSO

[DM06] A. B. Dieker and M. Mandjes. Fast simulation of overflow probabilities in a queue with Gaussian input. ACM Transactions on Modeling and Computer Simulation, 16(2): 119–151, April 2006. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Damerdji:1999:TSM

[DN99] Halim Damerdji and Marvin K. Nakayama. Two-stage multiple-comparison procedures for steady-state simulations. ACM Transactions on Modeling and Computer Simulation, 9(1):1–30, January 1999. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Dickens:1996:ABT

[DNRD96] Phillip M. Dickens, David M.

Nicol, Paul F. Reynolds, Jr., and J. M. Duva. Analysis of bounded time warp and comparison with YAWNS. ACM Transactions on Modeling and Computer Simulation, 6(4): 297–320, October 1996. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Davies:1993:SMM

[DOD93] Ruth M. Davies, Robert M.

O’Keefe, and Huw T. O. Davies. Simplifying the modeling of multiple activities, multiple queuing, and interruptions: a new low-level data structure. ACM Transactions on Modeling and Computer Simulation, 3(4):332–346, October 1993. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Doornik:2007:CHP

[Doo07] Jurgen A. Doornik. Conversion of high-period random numbers to floating point. ACM Transactions on Modeling and Computer Simulation, 17(1):3:1–3:5, January 2007. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Dassios:2018:ESC

[DQZ18] Angelos Dassios, Yan Qu, and Hongbiao Zhao. Exact simulation for a class of tempered stable and related distributions. ACM Transactions on Modeling and Computer Simulation, 28(3):20:1–20:??, August 2018. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Doucet:2013:ISI

[DR13] Arnaud Doucet and Christian P. Robert. Introduction to Special Issue on Monte Carlo Methods in Statistics. ACM Transactions on Modeling and Computer Simulation, 23(1):1:1–1:??, January 2013. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Silva:2023:ORP

[DSR23] Carina Da Silva, Stefan Schupp, and Anne Remke. Optimizing reachability probabilities for a restricted class of stochastic hybrid automata via flowpipe construction. ACM Transactions on Modeling and Computer Simulation, 33(4):18:1–18:??, October 2023. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL

https://dl.acm.org/doi/
10.1145/3607197.

Diallo:2022:ISS

[DT22] Saikou Y. Diallo and Andreas Tolk. Introduction to the special section on PADS 2021. ACM Transactions on Modeling and Computer Simulation, 32(4):25:1–25:??, October 2022. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL

https://dl.acm.org/doi/
10.1145/3579840.

Deng:2019:KNE

[DTCU19] Jie Deng, Gareth Tyson, Felix Cuadrado, and Steve Uhlig. Keddah: Network evaluation powered by simulating distributed application traffic. ACM Transactions on Modeling and Computer Simulation, 29(3):16:1–16:??, July 2019. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3301503.

Dzwinel:2016:PPA

[DWYM16] Witold Dzwinel, Rafal Wcisło, David A. Yuen, and Shea Miller. PAM: Particle automata in modeling of multiscale biological systems. ACM Transactions on Modeling and Computer Simulation, 26(3): 20:1–20:??, February 2016. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Deng:2003:SHD

[DX03] Lih-Yuan Deng and Hongquan Xu. A system of highdimensional, efficient, longcycle and portable uniform random number generators. ACM Transactions on Modeling and Computer Simulation, 13(4):299–309, October 2003. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Eick:1993:SRP

[EGLW93] Stephen G. Eick, Albert G. Greenberg, Boris D. Lubachevsky, and Alan Weiss. Synchronous relaxation for parallel simulations with applications to circuit-switched networks. ACM Transactions on Modeling and Computer Simulation, 3(4):287–314, October 1993. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Eichenauer-Herrmann:1995:PNG

[EH95] Jürgen Eichenauer-Herrmann. Pseudorandom number generation by nonlinear methods. International Statistical Review = Revue Internationale de Statistique, 63(2): 247–255, August 1995. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/1403620.

Eckman:2018:RSD

[EH18] David J. Eckman and Shane G.

Henderson. Reusing search data in ranking and selection: What could possibly go wrong? ACM Transactions on Modeling and Computer Simulation, 28(3):18:1–18:??, August 2018. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Eckman:2021:FCF

[EH21] David J. Eckman and Shane G.

Henderson. Fixed-confidence, fixed-tolerance guarantees for ranking-and-selection procedures. ACM Transactions on Modeling and Computer Simulation, 31(2):7:1–7:33, April 2021. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL

https://dl.acm.org/doi/
10.1145/3432754.

Eichenauer-Herrmann:1992:NIC

[EHG92] Jürgen Eichenauer-Herrmann and Holger Grothe. A new inversive congruential pseudorandom number generator with power of two modulus. ACM Transactions on Modeling and Computer Simulation, 2(1):1–11, January 1992. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Eichenauer-Herrmann:1994:DIP

[EHN94a] Jürgen Eichenauer-Herrmann and Harald Niederreiter. Digital inversive pseudorandom numbers. ACM Transactions on Modeling and Computer Simulation, 4(4):339–349, October 1994. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Eichenauer-Herrmann:1994:SIN

[EHN94b] Jürgen Eichenauer-Herrmann and Harald Niederreiter. On the statistical independence of nonlinear congruential pseudorandom numbers. ACM Transactions on Modeling and Computer Simulation, 4(1): 89–95, January 1994. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Esposito:2004:AIE

[EK04] Joel M. Esposito and Vijay Kumar. An asynchronous integration and event detection algorithm for simulating multi-agent hybrid systems. ACM Transactions on Modeling and Computer Simulation, 14(4):363–388, October 2004. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Esposito:2007:SED

[EK07] Joel M. Esposito and Vijay Kumar. A state event detection algorithm for numerically simulating hybrid systems with model singularities. ACM Transactions on Modeling and Computer Simulation, 17(1):??, January 2007. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Erickson:2000:OSC

[ELL00] K. Bruce Erickson, Richard E.

Ladner, and Anthony Lamarca. Optimizing static calendar queues. ACM Transactions on Modeling and Computer Simulation, 10(3):179–214, July 2000. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Emmerich:1998:SIP

[Emm98] Frank Emmerich. Statistical independence properties of inversive pseudorandom vectors over parts of the period. ACM Transactions on Modeling and Computer Simulation, 8(2):140–152, April 1998. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Entacher:1998:BSW

[Ent98] Karl Entacher. Bad subsequences of well-known linear congruential pseudorandom number generators. ACM Transactions on Modeling and Computer Simulation, 8(1): 61–70, January 1998. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Entacher:1999:PSL

[Ent99]  Karl Entacher.  Parallel

streams of linear random numbers in the spectral test. ACM Transactions on Modeling and Computer Simulation, 9(1): 31–44, January 1999. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Erazo:2015:SNS

[ERL15] Miguel A. Erazo, Rong Rong, and Jason Liu. Symbiotic network simulation and emulation. ACM Transactions on Modeling and Computer Simulation, 26(1):2:1–2:??, December 2015. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Epstein:1994:GTR

[ES94] Peter Epstein and Jörg-Rüdiger Sack. Generating triangulations at random. ACM Transactions on Modeling and Computer Simulation, 4(3): 267–278, July 1994. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Ernst:2021:FHS

[ESZH21] Gidon Ernst, Sean Sedwards, Zhenya Zhang, and Ichiro Hasuo. Falsification of hybrid systems using adaptive probabilistic search. ACM Transactions on Modeling and Computer Simulation, 31(3):18:1–18:22, July 2021. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3459605.

Ewald:2014:SDS

[EU14] Roland Ewald and Adelinde M.

Uhrmacher. SESSL: a domainspecific language for simulation experiments. ACM Transactions on Modeling and Computer Simulation, 24(2): 11:1–11:??, February 2014. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Eldabi:2011:ISI

[EY11] Tillal Eldabi and Terry Young.

Introduction to special issue on healthcare modeling and simulation. ACM Transactions on Modeling and Computer Simulation, 21(4):22:1– 22:??, August 2011. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Fishman:2006:HHT

[FA06] George S. Fishman and Ivo J. B. F. Adan. How heavytailed distributions affect simulation-generated time averages. ACM Transactions on Modeling and Computer Simulation, 16(2):152–173, April 2006. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Feldkamp:2022:MUG

[FBCS22] Niclas Feldkamp, Soeren Bergmann, Florian Conrad, and Steffen Strassburger. A method using generative adversarial networks for robustness optimization. ACM Transactions on Modeling and Computer Simulation, 32 (2):12:1–12:22, April 2022. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3503511.

Feldkamp:2020:KDS

[FBS20] Niclas Feldkamp, Soeren Bergmann, and Steffen Strassburger. Knowledge discovery in simulation data. ACM Transactions on Modeling and Computer Simulation, 30(4): 24:1–24:25, December 2020. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3391299.

Fishwick:2005:MRA

[FDD05] Paul Fishwick, Timothy Davis, and Jane Douglas. Model representation with aesthetic computing: Method and empirical study. ACM Transactions on Modeling and Computer Simulation, 15(3): 254–279, July 2005. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Falkner:1999:FSN

[FDL99] Matthias Falkner, Michael Devetsikiotis, and Ioannis Lambadaris. Fast simulation of networks of queues with effective and decoupling bandwidths. ACM Transactions on Modeling and Computer Simulation, 9(1):45–58, January 1999. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Filippone:2016:MCF

[FDMS16] Giuseppe Filippone, Donato D’ambrosio, Davide Marocco, and William Spataro. Morphological coevolution for fluid dynamical-related risk mitigation. ACM Transactions on Modeling and Computer Simulation, 26(3):18:1– 18:??, February 2016. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Fioretto:2015:CCB

[FDP15] Ferdinando Fioretto, Agostino Dovier, and Enrico Pontelli. Constrained communitybased gene regulatory network inference. ACM Transactions on Modeling and Computer Simulation, 25(2):11:1– 11:??, February 2015. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Fussl:2013:EMB

[FFSF13] Agnes Fussl, Sylvia Frühwirth-Schnatter, and Rudolf Frühwirth. Efficient MCMC for binomial logit models. ACM Transactions on Modeling and Computer Simulation, 23(1):3:1– 3:??, January 2013. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Frolund:1998:DTS

[FG98] Svend Frølund and Pankaj Garg. Design-time simulation of a large-scale, distributed object system. ACM Transactions on Modeling and Computer Simulation, 8(4):374– 400, October 1998. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Foley:1999:CIU

[FG99] Robert D. Foley and David Goldsman. Confidence intervals using orthonormally weighted standardized time series. ACM Transactions on Modeling and Computer Simulation, 9(4):297–325, October 1999. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Fujimoto:1997:CGV

[FH97] Richard M. Fujimoto and Maria Hybinette. Computing global virtual time in shared-memory multiprocessors. ACM Transactions on Modeling and Computer Simulation, 7(4):425–446, October 1997. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Feldman:2018:SAB

[FH18] Guy Feldman and Susan R. Hunter. SCORE allocations for bi-objective ranking and selection. ACM Transactions on Modeling and Computer Simulation, 28(1):7:1– 7:??, January 2018. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Feng:2009:FBB

[FHD09] Benjamin Zhong Ming Feng, Changcheng Huang, and Michael Devetsikiotis. FISTE: a black box approach for end-to-end QoS management. ACM Transactions on Modeling and Computer Simulation, 19(4):16:1–16:??, October 2009. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Feng:2016:AMC

[FHG16] Cheng Feng, Jane Hillston, and Vashti Galpin. Automatic moment-closure approximation of spatially distributed collective adaptive systems. ACM Transactions on Modeling and Computer Simulation, 26(4):26:1–26:??, May 2016. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). See successful replication report [Lüc16].

Fishwick:1992:IAS

[Fis92] Paul A. Fishwick. An integrated approach to system modeling using a synthesis of artificial intelligence, software engineering and simulation methodologies. ACM Transactions on Modeling and Computer Simulation, 2(4):

307–330, October 1992. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Felderman:1991:BAS

[FK91] Robert E. Felderman and Leonard Kleinrock. Bounds and approximations for selfinitiating distributed simulation without lookahead. ACM Transactions on Modeling and Computer Simulation, 1(4): 386–406, October 1991. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Faure:2009:GHS

[FL09] Henri Faure and Christiane Lemieux. Generalized Halton sequences in 2008: a comparative study. ACM Transactions on Modeling and Computer Simulation, 19(4):15:1– 15:??, October 2009. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Flatscher:2002:MEC

[Fla02] Rony G. Flatscher. Metamodeling in EIA/CDIF—metametamodel and metamodels. ACM Transactions on Modeling and Computer Simulation, 12(4):322–342, October 2002. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Fuks:2001:PDM

[FLV01] Henryk Fukś, Anna T.

Lawniczak, and Stanislav Volkov. Packet delay in models of data networks. ACM Transactions on Modeling and Computer Simulation, 11(3): 233–250, July 2001. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Fonseca:2000:EBS

[FMN00] Nelson L. S. Fonseca, Gilberto S.

Mayor, and Cesar A. V. Neto. On the equivalent bandwidth of self-similar sources. ACM Transactions on Modeling and Computer Simulation, 10(2): 104–124, April 2000. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Fu:2003:GE

[FN03] Michael Fu and Barry Nelson. Guest editorial. ACM Transactions on Modeling and Computer Simulation, 13(2): 105–107, April 2003. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Feng:2017:GSR

[FS17] Mingbin Feng and Jeremy Staum. Green simulation: Reusing the output of repeated experiments. ACM Transactions on Modeling and Computer Simulation, 27(4): 23:1–23:??, December 2017.

CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). See replication report [Nel17].

Feng:2021:GSD

[FS21] Mingbin Feng and Jeremy Staum. Green simulation with database Monte Carlo. ACM Transactions on Modeling and Computer Simulation, 31(1):4:1–4:26, February 2021. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/

3429336. See replication report [Pel21].

Fleming:1995:EMC

[FSS95] Philip J. Fleming, Dennis Schaeffer, and Burton Simon. Efficient Monte-Carlo simulation of a product-form model for a cellular system with dynamic resource sharing. ACM Transactions on Modeling and Computer Simulation, 5(1): 3–21, January 1995. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Fujimoto:2016:RCP

[Fuj16] Richard M. Fujimoto. Research challenges in parallel and distributed simulation. ACM Transactions on Modeling and Computer Simulation, 26(4):22:1–22:??, May 2016. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Falsafi:1997:MCP

[FW97] Babak Falsafi and David A.

Wood. Modeling cost/performance of a parallel computer simulator. ACM Transactions on Modeling and Computer Simulation, 7(1):104–130, January 1997. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Fishwick:1992:MMQ

[FZ92] Paul A. Fishwick and Bernard P.

Zeigler. A multimodel methodology for qualitative model engineering. ACM Transactions on Modeling and Computer Simulation, 2(1): 52–81, January 1992. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Gupta:2014:VES

[GAG14] Vivek Gupta, Sigrún Andradóttir, and David Goldsman. Variance estimation and sequential stopping in steadystate simulations using linear regression. ACM Transactions on Modeling and Computer Simulation, 24(2):7:1– 7:??, February 2014. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Giabbanelli:2019:VAI

[GB19] Philippe J. Giabbanelli and Magda Baniukiewicz. Visual analytics to identify temporal patterns and variability in simulations from cellular automata. ACM Transactions on Modeling and Computer Simulation, 29(1):5:1–5:??, February 2019. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Gupta:2014:GSG

[GBA^{+}14] Sandeep K. S. Gupta, Ayan Banerjee, Zahra Abbasi, Georgios Varsamopoulos, Michael Jonas, Joshua Ferguson, Rose Robin Gilbert, and Tridib Mukherjee. GDCSim: a simulator for green data center design and analysis. ACM Transactions on Modeling and Computer Simulation, 24(1): 3:1–3:??, January 2014. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Giabbanelli:2022:ISS

[GC22] Philippe J. Giabbanelli and Christopher D. Carothers. Introduction to the special section on PADS 2020. ACM Transactions on Modeling and Computer Simulation, 32(2):8:1–8:2, April 2022. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3498363.

Granieri:1995:PPH

[GCB95] John P. Granieri, Jonathan Crabtree, and Norman I. Badler. Production and playback of human figure motion for visual simulation. ACM Transactions on Modeling and Computer Simulation, 5(3): 222–241, July 1995. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Ghoshdastidar:2014:SFA

[GDB14] Debarghya Ghoshdastidar, Ambedkar Dukkipati, and Shalabh Bhatnagar. Smoothed functional algorithms for stochastic optimization using q-Gaussian distributions. ACM Transactions on Modeling and Computer Simulation, 24(3):17:1–17:??, June 2014. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Gore:2014:CCM

[GDP14] Ross Gore, Saikou Diallo, and Jose Padilla. ConceVE: Conceptual modeling and formal validation for everyone. ACM Transactions on Modeling and Computer Simulation, 24(2):12:1–12:??, February 2014. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Gros:2023:DES

[GGH^{+}23] Timo P. Gros, Joschka Groß, Daniel Höller, Jörg Hoffmann, Michaela Klauck, Hendrik Meerkamp, Nicola J. Müller, Lukas Schaller, and Verena Wolf. DSMC evaluation stages: Fostering robust and safe behavior in deep reinforcement learning — extended version. ACM Transactions on Modeling and Computer Simulation, 33(4): 17:1–17:??, October 2023. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3607198.

Glynn:1991:APR

[GH91] Peter W. Glynn and Philip Heidelberger. Analysis of parallel replicated simulations under a completion time constraint. ACM Transactions on Modeling and Computer Simulation, 1(1):3–23, January 1991. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Ghosh:2003:BNM

[GH03] Soumyadip Ghosh and Shane G. Henderson. Behavior of the NORTA method for correlated random vector generation as the dimension increases. ACM Transactions on Modeling and Computer Simulation, 13(3): 276–294, July 2003. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Ghosh:2006:CBN

[GH06] Soumyadip Ghosh and Shane G. Henderson. Corrigendum: Behavior of the NORTA method for correlated random vector generation as the dimension increases. ACM Transactions on Modeling and Computer Simulation, 16(1):93–94, January 2006. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Ghosh:2009:CBN

[GH09] Soumyadip Ghosh and Shane G.

Henderson. Corrigendum: Behavior of the NORTA method for correlated random vector generation as the dimension increases. ACM Transactions on Modeling and Computer Simulation, 19(4):20:1–20:??, October 2009. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Glynn:2015:GEI

[GH15a] Peter W. Glynn and Peter J.

Haas. Guest editors’ introduction to special issue honoring Donald L. Iglehart. ACM Transactions on Modeling and Computer Simulation, 25(4): 21:1–21:??, November 2015. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Glynn:2015:TRI

[GH15b] Peter W. Glynn and Peter J.

Haas. On transience and recurrence in irreducible finitestate stochastic systems. ACM Transactions on Modeling and Computer Simulation, 25(4): 25:1–25:??, November 2015. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Georgoulas:2018:PPP

[GHS18] Anastasis Georgoulas, Jane Hillston, and Guid Sanguinetti. ProPPA: Probabilistic programming for stochastic dynamical systems. ACM Transactions on Modeling and Computer Simulation, 28(1): 3:1–3:??, January 2018. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Glynn:2013:ASE

[GJ13] Peter W. Glynn and Sandeep Juneja. Asymptotic simulation efficiency based on large deviations. ACM Transactions on Modeling and Computer Simulation, 23(3):20:1–20:??, July 2013. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Glasserman:1995:AIS

[GK95] Paul Glasserman and Shing-Gang Kou. Analysis of an importance sampling estimator for tandem queues. ACM Transactions on Modeling and Computer Simulation, 5(1): 22–42, January 1995. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Goresky:2003:EMC

[GK03] Mark Goresky and Andrew Klapper. Efficient multiplywith-carry random number generators with maximal period. ACM Transactions on Modeling and Computer Simulation, 13(4):310–321, October 2003. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Gorder:2019:RSN

[GK19] Bjürn Görder and Michael Kolonko. Ranking and selection: a new sequential Bayesian procedure for use with common random numbers. ACM Transactions on Modeling and Computer Simulation, 29(1):2:1–2:24, February 2019. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Grassmann:2005:SMR

[GL05] Winfried K. Grassmann and Jingxiang Luo. Simulating Markov-reward processes with rare events. ACM Transactions on Modeling and Computer Simulation, 15(2): 138–154, April 2005. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Gonsiorowski:2017:AMG

[GLC17] Elsa Gonsiorowski, Justin M.

Lapre, and Christopher D. Carothers. Automatic model generation for gate-level circuit PDES with reverse computation. ACM Transactions on Modeling and Computer Simulation, 27(2):12:1–12:??, July 2017. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Greenberg:1996:SPD

[GLM96] Albert G. Greenberg, Boris D. Lubachevsky, and Isi Mitrani. Superfast parallel discrete event simulations. ACM Transactions on Modeling and Computer Simulation, 6(2): 107–136, April 1996. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Gallardo:2001:FSB

[GMOB01] José R. Gallardo, Dimitrios Makrakis, and Luis Orozco-Barbosa. Fast simulation of broadband telecommunications networks carrying longrange dependent bursty traffic. ACM Transactions on Modeling and Computer Simulation, 11(3):274–293, July 2001. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Goualard:2022:DRF

[Gou22] Frédéric Goualard. Drawing random floating-point numbers from an interval. ACM Transactions on Modeling and Computer Simulation, 32(3):16:1–16:24, July 2022. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3503512.

Gunal:2011:DGS

[GP11] Murat M. Günal and Michael Pidd. DGHPSim: Generic simulation of hospital performance. ACM Transactions on Modeling and Computer Simulation, 21(4):23:1–23:??, August 2011. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Gore:2015:SDS

[GRK^{+}15] Ross Gore, Paul F. Reynolds Jr., David Kamensky, Saikou Diallo, and Jose Padilla. Statistical debugging for simulations. ACM Transactions on Modeling and Computer Simulation, 25(3):16:1–16:??, April 2015. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Goyal:2012:SCB

[GS12] Vineet Goyal and Karl Sigman. On simulating a class of Bernstein polynomials. ACM Transactions on Modeling and Computer Simulation, 22(2): 12:1–12:5, March 2012. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Galpin:2018:MMP

[GZWG18] Vashti Galpin, Natalia Zoń, Pia Wilsdorf, and Stephen Gilmore. Mesoscopic modelling of pedestrian movement using Carma and its tools. ACM Transactions on Modeling and Computer Simulation, 28(2):11:1–11:??, April 2018. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Hunter:2019:IMS

[HAA^{+}19] Susan R. Hunter, Eric A. Applegate, Viplove Arora, Bryan Chong, Kyle Cooper, Oscar Rincón-Guevara, and Carolina Vivas-Valencia. An introduction to multiobjective simulation optimization. ACM Transactions on Modeling and Computer Simulation, 29(1): 7:1–7:??, February 2019. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Housseman:2011:IRI

[HAFDP11] Sylvain Housseman, Nabil Absi, Dominique Feillet, and Stéphane Dauzére-Pèrés. Impacts of radio-identification on cryo-conservation centers. ACM Transactions on Modeling and Computer Simulation, 21(4):27:1–27:??, August 2011. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Healey:2014:SPS

[HAK14] Christopher Healey, Sigrún Andradóttir, and Seong-Hee Kim. Selection procedures for simulations with multiple constraints under independent and correlated sampling. ACM Transactions on Modeling and Computer Simulation, 24(3):14:1–14:??, June 2014. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Healey:1995:VRT

[HBE95] Christopher G. Healey, Kellogg S. Booth, and James T. Enns. Visualizing real-time multivariate data using preattentive processing. ACM Transactions on Modeling and Computer Simulation, 5(3): 190–221, July 1995. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Hormann:1996:RIG

[HD96] W. Hörmann and G. Derflinger. Rejection-inversion to generate variates from monotone discrete distributions. ACM Transactions on Modeling and Computer Simulation, 6(3):169–184, July 1996. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Hein:1998:PDE

[HD98] Axel Hein and Mario Dal Cin.

Performance and dependability evaluation of scalable massively parallel computer systems with conjoint simulation. ACM Transactions on Modeling and Computer Simulation, 8(4):333–373, October 1998. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Hormann:2002:FGO

[HD02] Wolfgang Hörmann and Gerhard Derflinger. Fast generation of order statistics. ACM Transactions on Modeling and Computer Simulation, 12(2): 83–93, April 2002. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Hsu:2007:AAA

[HD07] Chih-Chieh Hsu and Michael Devetsikiotis. An adaptive approach to accelerated evaluation of highly available services. ACM Transactions on Modeling and Computer Simulation, 18(1):1:1–1:26, December 2007. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Homem-De-Mello:2003:VSM

[HDM03] Tito Homem-De-Mello. Variablesample methods for stochastic optimization. ACM Transactions on Modeling and Computer Simulation, 13(2): 108–133, April 2003. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Henderson:2012:SCG

[HE12] Shane G. Henderson and Samuel M. T. Ehrlichman. Sharpening comparisons via Gaussian copulas and semidefinite programming. ACM Transactions on Modeling and Computer Simulation, 22(4): 22:1–22:??, November 2012. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Heidelberger:1995:FSR

[Hei95]  Philip Heidelberger.  Fast

simulation of rare events in queueing and reliability models. ACM Transactions on Modeling and Computer Simulation, 5(1):43–85, January 1995. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Heidelberger:1997:E

[Hei97] Philip Heidelberger. Editorial.

ACM Transactions on Modeling and Computer Simulation, 7(1):3, January 1997. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Helms:2015:ARA

[HERU15] Tobias Helms, Roland Ewald, Stefan Rybacki, and Adelinde M. Uhrmacher. Automatic runtime adaptation for componentbased simulation algorithms. ACM Transactions on Modeling and Computer Simulation, 26(1):7:1–7:??, December 2015. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Hybinette:2001:CPS

[HF01] Maria Hybinette and Richard M.

Fujimoto. Cloning parallel simulations. ACM Transactions on Modeling and Computer Simulation, 11(4):378– 407, October 2001. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Henderson:2001:RSS

[HG01] Shane G. Henderson and Peter W. Glynn. Regenerative steady-state simulation of discrete-event systems. ACM Transactions on Modeling and Computer Simulation, 11(4): 313–345, October 2001. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Huang:2016:MMT

[HHFS16] Shell Ying Huang, Wen Jing Hsu, Hui Fang, and Tiancheng Song. MTSS — a marine traffic simulation system and scenario studies for a major hub port. ACM Transactions on Modeling and Computer Simulation, 27(1):3:1–3:??, November 2016. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Hahn:2019:IMD

[HHH^{+}19] Ernst Moritz Hahn, Vahid Hashemi, Holger Hermanns, Morteza Lahijanian, and Andrea Turrini. Interval Markov decision processes with multiple objectives: From robust strategies to Pareto curves. ACM Transactions on Modeling and Computer Simulation, 29(4):27:1– 27:??, December 2019. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3309683.

Haas:2014:GEI

[HHL14a] Peter J. Haas, Shane G. Henderson, and Pierre L’Ecuyer. Guest editors’ introduction to special issue on the Third INFORMS Simulation Society Research Workshop. ACM Transactions on Modeling and Computer Simulation, 24(1): 1:1–1:??, January 2014. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Hong:2014:MCM

[HHL14b] L. Jeff Hong, Zhaolin Hu, and Guangwu Liu. Monte Carlo methods for value-at-risk and conditional value-at-risk: a review. ACM Transactions on Modeling and Computer Simulation, 24(4):22:1–22:??, August 2014. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Hong:2011:MSS

[HHY11] Yang Hong, Changcheng Huang, and James Yan. Modeling and simulation of SIP tandem server with finite buffer. ACM Transactions on Modeling and Computer Simulation, 21(2):11:1– 11:??, February 2011. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Hickernell:1996:MSD

[Hic96] Fred J. Hickernell. The mean square discrepancy of randomized nets. ACM Transactions on Modeling and Computer Simulation, 6(4):274–296, October 1996. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Hsieh:2004:EPB

[HIG04] Ming-Hua Hsieh, Donald L. Iglehart, and Peter W. Glynn. Empirical performance of bias-reducing estimators for regenerative steady-state simulations. ACM Transactions on Modeling and Computer Simulation, 14(4):325– 343, October 2004. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Hillston:2017:RCR

[Hil17] Jane Hillston. Replicated Computations Results (RCR) report for “Semantics and Efficient Simulation Algorithms for an Expressive Multi-Level Modeling Language”. ACM Transactions on Modeling and Computer Simulation, 27(2): 9:1–9:??, July 2017. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Han:2021:IPF

[HKP21] Jungmin Han, Seong-Hee Kim, and Chuljin Park. Improved penalty function with memory for stochastically constrained optimization via simulation. ACM Transactions on Modeling and Computer Simulation, 31(4):24:1–24:26, October 2021. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL

https://dl.acm.org/doi/
10.1145/3465333.

Hormann:2003:CRV

[HL03] Wolfgang Hörmann and Josef Leydold. Continuous random variate generation by fast numerical inversion. ACM Transactions on Modeling and Computer Simulation, 13 (4):347–362, October 2003. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL http://

statistik.wu-wien.ac.at/
unuran/.

He:2010:SOU

[HLC^{+}10] Donghai He, Loo Hay Lee, Chun-Hung Chen, Michael C. Fu, and Segev Wasserkrug. Simulation optimization using the cross-entropy method with optimal computing budget allocation. ACM Transactions on Modeling and Computer Simulation, 20(1):4:1–4:22 + 9 (online appendix), January 2010. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Hernandez:2012:CNO

[HLC12] Alejandro S. Hernandez, Thomas W. Lucas, and Matthew Carlyle. Constructing nearly orthogonal Latin hypercubes for any nonsaturated run-variable combination. ACM Transactions on Modeling and Computer Simulation, 22(4):20:1– 20:??, November 2012. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Hormann:2007:ITD

[HLD07] Wolfgang Hörmann, Josef Leydold, and Gerhard Derflinger. Inverse transformed density rejection for unbounded monotone densities. ACM Transactions on Modeling and Computer Simulation, 17(4):18:1–18:??, September 2007. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Hung:2008:MSS

[HM08] Ying-Chao Hung and George Michailidis. Modeling, scheduling, and simulation of switched processing systems. ACM Transactions on Modeling and Computer Simulation, 18(3): 12:1–12:??, July 2008. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Hellekalek:1998:WST

[HN98] Peter Hellekalek and Harald Niederreiter. The weighted spectral test: diaphony. ACM Transactions on Modeling and Computer Simulation, 8(1): 43–60, January 1998. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Hong:2007:FLC

[HN07] L. Jeff Hong and Barry L. Nelson. A framework for locally convergent random-search algorithms for discrete optimization via simulation. ACM Transactions on Modeling and Computer Simulation, 17(4): 19:1–19:??, September 2007. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Hu:2009:ISO

[HN09] Xiaolin Hu and Lewis Ntaimo.

Integrated simulation and optimization for wildfire containment. ACM Transactions on Modeling and Computer Simulation, 19(4):19:1–19:??, October 2009. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Hofert:2011:SET

[Hof11] Marius Hofert. Sampling exponentially tilted stable distributions. ACM Transactions on Modeling and Computer Simulation, 22(1):3:1– 3:??, December 2011. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Hormann:1994:NQR

[Hör94] Wolfgang Hörmann. A note on the quality of random variates generated by the ratio of uniforms method. ACM Transactions on Modeling and Computer Simulation, 4(1):96–106, January 1994. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Hernandez:2007:DTH

[HPA07] José Alberto Hernández, Iain W. Phillips, and Javier Aracil. Discrete-time heavytailed chains, and their properties in modeling network traffic. ACM Transactions on Modeling and Computer Simulation, 17(4):17:1–17:??, September 2007. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Huls:2021:SSC

[HPS^{+}21] Jannik Hüls, Carina Pilch, Patricia Schinke, Henner Niehaus, Joanna Delicaris, and Anne Remke. Statespace construction of hybrid Petri nets with multiple stochastic firings. ACM Transactions on Modeling and Computer Simulation, 31(3):13:1–13:37, July 2021. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3449353.

Hult:2012:ISM

[HS12] Henrik Hult and Jens Svensson. On importance sampling with mixtures for random walks with heavy tails. ACM Transactions on Modeling and Computer Simulation, 22(2):8:1–8:21, March 2012.

CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Hanai:2019:EDS

[HSL^{+}19] Masatoshi Hanai, Toyotaro Suzumura, Elvis S. Liu, Georgios Theodoropoulos, and Kalyan S. Perumalla. Exactdifferential simulation: Differential processing of large-scale discrete event simulations. ACM Transactions on Modeling and Computer Simulation, 29(3):18:1–18:??, July 2019. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3301499.

Heidelberger:1994:BRE

[HSN94] Philip Heidelberger, Perwez Shahabuddin, and Victor F. Nicola. Bounded relative error in estimating transient measures of highly dependable non-Markovian systems. ACM Transactions on Modeling and Computer Simulation, 4(2):137–164, April 1994. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

He:2024:SAM

[HSS24] Linyun He, Uday V. Shanbhag, and Eunhye Song. Stochastic approximation for multiperiod simulation optimization with streaming input data. ACM Transactions on Modeling and Computer Simulation, 34(2):6:1–6:??, April 2024. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3617595.

Haraszti:1999:TDP

[HT99] Zsolt Haraszti and J. Keith Townsend. The theory of direct probability redistribution and its application to rare event simulation. ACM Transactions on Modeling and Computer Simulation, 9(2): 105–140, April 1999. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Haas:2020:ISI

[HT20] Peter J. Haas and Georgios Theodoropoulos. Introduction to the special issue for towards an ecosystem of simulation models and data. ACM Transactions on Modeling and Computer Simulation, 30(4): 20:1–20:3, December 2020. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3425907.

Heidergott:2009:GEC

[HVA09] Bernd Heidergott and Felisa J. Vázquez-Abad. Gradient estimation for a class of systems with bulk services: a problem in public transportation. ACM Transactions on Modeling and Computer Simulation, 19(3):

13:1–13:??, June 2009. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Heidergott:2010:GED

[HVAPFY10] Bernd Heidergott, Felisa J.

Vázquez-Abad, Georg Pflug, and Taoying Farenhorst-Yuan. Gradient estimation for discrete-event systems by measure-valued differentiation. ACM Transactions on Modeling and Computer Simulation, 20(1):5:1–5:28, January 2010. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Hellekalek:2003:EEC

[HW03] Peter Hellekalek and Stefan Wegenkittl. Empirical evidence concerning AES. ACM Transactions on Modeling and Computer Simulation, 13(4):322–333, October 2003. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL

http://random.mat.sbg.ac.
at/ftp/pub/publications/
peter/aes_sub.ps;   http:
//random.mat.sbg.ac.at/~
peter/slides_YACC04.pdf.

Hu:2019:DAF

[HW19] Xiaolin Hu and Peisheng Wu.

A data assimilation framework for discrete event simulations. ACM Transactions on Modeling and Computer Simulation, 29(3):17:1–17:??, July 2019. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3301502.

Hesham:2021:EMP

[HW21] Omar Hesham and Gabriel Wainer. Explicit modeling of personal space for improved local dynamics in simulated crowds. ACM Transactions on Modeling and Computer Simulation, 31(4): 23:1–23:29, October 2021. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3462202.

Hamze:2013:SAR

[HWdF13] Firas Hamze, Ziyu Wang, and Nando de Freitas. Selfavoiding random dynamics on integer complex systems. ACM Transactions on Modeling and Computer Simulation, 23(1):9:1–9:??, January 2013. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Helms:2017:SES

[HWMU17] Tobias Helms, Tom Warnke, Carsten Maus, and Adelinde M. Uhrmacher. Semantics and efficient simulation algorithms of an expressive multilevel modeling language. ACM Transactions on Modeling and Computer Simulation, 27(2): 8:1–8:??, July 2017. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Hannon:2018:CSE

[HYJ^{+}18] Christopher Hannon, Jiaqi Yan, Dong Jin, Chen Chen, and Jianhui Wang. Combining simulation and emulation systems for smart grid planning and evaluation. ACM Transactions on Modeling and Computer Simulation, 28(4): 27:1–27:??, October 2018. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Hannon:2021:DVT

[HYJ21] Christopher Hannon, Jiaqi Yan, and Dong Jin. Distributed virtual time-based synchronization for simulation of cyber-physical systems. ACM Transactions on Modeling and Computer Simulation, 31(2):10:1–10:24, April 2021. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3446237.

Hu:2014:MBA

[HZF14] Jiaqiao Hu, Enlu Zhou, and Qi Fan. Model-based annealing random search with stochastic averaging. ACM Transactions on Modeling and Computer Simulation, 24(4): 21:1–21:??, August 2014. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Inoue:1999:EES

[ICC99] Koichiro Inoue, Stephen E. Chick, and Chun-Hung Chen. An empirical evaluation of several methods to select the best system. ACM Transactions on Modeling and Computer Simulation, 9(4):381– 407, October 1999. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Inacio:2012:FSP

[IFPM12] Pedro R. M. Inácio, Mário M. Freire, Manuela Pereira, and Paulo P. Monteiro. Fast synthesis of persistent fractional Brownian motion. ACM Transactions on Modeling and Computer Simulation, 22(2): 11:1–11:21, March 2012. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Ingalls:2000:ITI

[IMW00] Ricki G. Ingalls, Douglas J. Morrice, and Andrew B. Whinston. The implementation of temporal intervals in qualitative simulation graphs. ACM Transactions on Modeling and Computer Simulation, 10(3):215–240, July 2000. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Jha:2000:SEL

[JB00] Vikas Jha and Rajive Bagrodia. Simultaneous events and lookahead in simulation protocols. ACM Transactions on Modeling and Computer Simulation, 10(3):241–267, July 2000. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Jefferson:2022:VTIa

[JB22a] David R. Jefferson and Peter Barnes. Virtual time III, Part 1: Unified virtual time synchronization for parallel discrete event simulation. ACM Transactions on Modeling and Computer Simulation, 32(4): 23:1–23:??, October 2022. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3505248.

Jefferson:2022:VTIb

[JB22b] David R. Jefferson and Peter D. Barnes. Virtual time III, Part 2: Combining conservative and optimistic synchronization. ACM Transactions on Modeling and Computer Simulation, 32(4): 24:1–24:??, October 2022. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3505249.

Joseph:2022:ROS

[JBH^{+}22] Jan Moritz Joseph, Lennart Bamberg, Imad Hajjar, Behnam Razi Perjikolaei, Alberto García-Ortiz, and Thilo Pionteck. Ratatoskr: an open-source framework for in-depth power, performance, and area analysis and optimization in 3D NoCs. ACM Transactions on Modeling and Computer Simulation, 32(1):3:1–3:21, January 2022. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3472754.

Jin:2011:SEG

[JC11] Zhanpeng Jin and Allen C. Cheng. SubsetTrio: An evolutionary, geometric, and statistical benchmark subsetting framework. ACM Transactions on Modeling and Computer Simulation, 21(3):21:1– 21:??, March 2011. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Jacyna:2024:IMW

[JFST24] Garry Jacyna, Damon Frezza, David M. Slater, and James R. Thompson. An improved model of wavelet leader covariance for estimating multifractal properties. ACM Transactions on Modeling and Computer Simulation, 34 (1):1:1–1:??, January 2024. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3631522.

Jasra:2014:AIO

[JKE14] Ajay Jasra, Nikolas Kantas, and Elena Ehrlich.

Approximate inference for observation-driven time series models with intractable likelihoods. ACM Transactions on Modeling and Computer Simulation, 24(3):13:1–13:??, June 2014. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Juneja:2007:AFS

[JKS07] S. Juneja, R. L. Karandikar, and P. Shahabuddin. Asymptotics and fast simulation for tail probabilities of maximum of sums of few random variables. ACM Transactions on Modeling and Computer Simulation, 17(2):7:1–7:35, April 2007. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Juneja:2005:ESB

[JN05] Sandeep Juneja and Victor Nicola. Efficient simulation of buffer overflow probabilities in Jackson networks with feedback. ACM Transactions on Modeling and Computer Simulation, 15(4):281–315, October 2005. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Jin:2015:PSV

[JN15] Dong Jin and David M.

Nicol. Parallel simulation and virtual-machine-based emulation of software-defined networks. ACM Transactions on Modeling and Computer Simulation, 26(1):8:1–8:??, December 2015. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Johnson:1996:RES

[Joh96] Brad C. Johnson. Radix-b extensions to some common empirical tests for pseudorandom number generators. ACM Transactions on Modeling and Computer Simulation, 6(4): 261–273, October 1996. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Juneja:2002:SHT

[JS02] Sandeep Juneja and Perwez Shahabuddin. Simulating heavy tailed processes using delayed hazard rate twisting. ACM Transactions on Modeling and Computer Simulation, 12(2):94–118, April 2002. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Jin:2001:FPS

[JSC01] Wei Jin, Xiaobai Sun, and Jeffrey S. Chase. Fast-Slim: prefetch-safe trace reduction for I/O cache simulation. ACM Transactions on Modeling and Computer Simulation, 11(2):135–160, April 2001. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Jegourel:2019:SSF

[JSD19] Cyrille Jegourel, Jun Sun, and Jin Song Dong. Sequential schemes for frequentist estimation of properties in statistical model checking. ACM Transactions on Modeling and Computer Simulation, 29(4): 25:1–25:??, December 2019. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3310226.

Jezequel:2023:UAS

[JV23] Jean-Marc Jézéquel and Antonio Vallecillo. Uncertaintyaware simulation of adaptive systems. ACM Transactions on Modeling and Computer Simulation, 33(3):8:1–8:??, July 2023. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL

https://dl.acm.org/doi/
10.1145/3589517.

Jin:2019:GET

[JW19] Kevin Jin and Philip Wilsey.

Guest editorial for the TOMACS special issue on the Principles of Advanced Discrete Simulation (PADS). ACM Transactions on Modeling and Computer Simulation, 29(2):8:1–8:??, April 2019. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3312749.

Jalilzadeh:2024:SAE

[JYE24] Afrooz Jalilzadeh, Farzad Yousefian, and Mohammadjavad Ebrahimi. Stochastic approximation for estimating the price of stability in stochastic Nash games. ACM Transactions on Modeling and Computer Simulation, 34(2):7:1–7:??, April 2024. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3632525.

Ji:2006:ISW

[JZTB06] Zhengrong Ji, Junlan Zhou, Mineo Takai, and Rajive Bagrodia. Improving scalability of wireless network simulation with bounded inaccuracies. ACM Transactions on Modeling and Computer Simulation, 16(4):329–356, October 2006. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Kawai:2010:AOA

[Kaw10] Reiichiro Kawai. Asymptotically optimal allocation of stratified sampling with adaptive variance reduction by strata. ACM Transactions on Modeling and Computer Simulation, 20(2):9:1–9:??, April 2010. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Kin:2010:GLT

[KC10] Wai Kin and Victor Chan. Generalized Lindley-type recursive representations for multiserver tandem queues with blocking. ACM Transactions on Modeling and Computer Simulation, 20(4):21:1– 21:??, October 2010. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Kim:2008:TRG

[KCK08] Chihurn Kim, Geon Ho Choe, and Dong Han Kim. Tests of randomness by the gambler’s ruin algorithm. Applied Mathematics and Computation, 199(1):195–210, May 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300307009873. See crit-

ical remarks [PJ10].

Kalayappan:2020:CCB

[KCS20] Rajshekar Kalayappan, Avantika Chhabra, and Smruti R. Sarangi. ChunkedTejas: a chunking-based approach to parallelizing a trace-driven architectural simulator. ACM Transactions on Modeling and Computer Simulation, 30(3): 15:1–15:21, July 2020. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3375397.

Knudde:2020:HGP

[KDV^{+}20] Nicolas Knudde, Vincent Dutordoir, Joachim Van Der Herten, Ivo Couckuyt, and Tom Dhaene. Hierarchical Gaussian process models for improved metamodeling. ACM Transactions on Modeling and Computer Simulation, 30(4):23:1–23:17, December 2020. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3384470.

Kapadia:2000:PUN

[KFL00] Nirav H. Kapadia, José A. B. Fortes, and Mark S. Lundstrom. The Purdue University network-computing hubs: running unmodified simulation tools via the WWW. ACM Transactions on Modeling and Computer Simulation, 10(1):39–57, January 2000. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Kuang:2018:RCR

[KH18] Xianyu Kuang and L. Jeff Hong. Replicated Computations Results (RCR) report for “Reusing Search Data in Ranking and Selection: What Could Possibly Go Wrong?”. ACM Transactions on Modeling and Computer Simulation, 28(3):19:1–19:??, August 2018. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Keller:2019:TDD

[KH19] Nicholas Keller and Xiaolin Hu. Towards data-driven simulation modeling for mobile agent-based systems. ACM Transactions on Modeling and Computer Simulation, 29 (1):1:1–1:??, February 2019. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3289229.

Kesidis:2008:MSR

[KHJ^{+}08] George Kesidis, Ihab Hamadeh, Youngmi Jin, Soranun Jiwasurat, and Milan Vojnović. A model of the spread of randomly scanning Internet worms that saturate access links. ACM Transactions on Modeling and Computer Simulation, 18(2):6:1–6:??, April 2008. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Kim:2005:CSF

[Kim05] Seong-Hee Kim. Comparison with a standard via fully sequential procedures. ACM Transactions on Modeling and Computer Simulation, 15(2): 155–174, April 2005. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Kiviat:1991:STD

[Kiv91] Philip J. Kiviat. Simulation, technology, and the decision process. ACM Transactions on Modeling and Computer Simulation, 1(2):89–98, April 1991. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Korkmaz:2000:SOT

[KK00] Turgay Korkmaz and Marwan Krunz. Source-oriented topology aggregation with multiple QoS parameters in hierarchical networks. ACM Transactions on Modeling and Computer Simulation, 10(4):295– 325, October 2000. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Kawai:2017:VWD

[KKTM17] Takaaki Kawai, Shigeru Kaneda, Mineo Takai, and Hiroshi Mineno. A virtual WLAN device model for high-fidelity wireless network emulation. ACM Transactions on Modeling and Computer Simulation, 27(3): 17:1–17:??, September 2017. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Kim:2002:TSM

[KLF02] Taewoo Kim, Jinho Lee, and Paul Fishwick. A two-stage modeling and simulation process for Web-based modeling and simulation. ACM Transactions on Modeling and Computer Simulation, 12(3): 230–248, July 2002. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Kumaran:2001:PFS

[KM01] Krishnan Kumaran and Debasis Mitra. Performance and fluid simulations of a novel shared buffer management system. ACM Transactions on Modeling and Computer Simulation, 11(1):43–75, January 2001. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Kim:2001:FSP

[KN01] Seong-Hee Kim and Barry L.

Nelson. A fully sequential procedure for indifference-zone selection in simulation. ACM Transactions on Modeling and Computer Simulation, 11(3): 251–273, July 2001. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Kroese:2002:EST

[KN02] Dirk P. Kroese and Victor F.

Nicola. Efficient simulation of a tandem Jackson network. ACM Transactions on Modeling and Computer Simulation, 12(2):119–141, April 2002. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Keane:1994:BF

[KO94] M. S. Keane and George L.

O’Brien. A Bernoulli factory. ACM Transactions on Modeling and Computer Simulation, 4(2):213–219, April 1994. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Kristiansen:2015:MME

[KPG15] Stein Kristiansen, Thomas Plagemann, and Vera Goebel. A methodology to model the execution of communication software for accurate network simulation. ACM Transactions on Modeling and Computer Simulation, 26(1):3:1– 3:??, December 2015. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Krantz:1996:AEA

[Kra96] Alan T. Krantz. Analysis of an efficient algorithm for the hard-sphere problem. ACM Transactions on Modeling and Computer Simulation, 6(3): 185–209, July 1996. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Kunz:2016:PEE

[KSL^{+}16] Georg Kunz, Mirko Stoffers, Olaf Landsiedel, Klaus Wehrle, and James Gross. Parallel expanded event simulation of tightly coupled systems. ACM Transactions on Modeling and Computer Simulation, 26(2):12:1–12:??, January 2016. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Kaplan:2003:FRT

[KSW03] Scott F. Kaplan, Yannis Smaragdakis, and Paul R. Wilson. Flexible reference trace reduction for VM simulations. ACM Transactions on Modeling and Computer Simulation, 13(1):1–38, January 2003. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Kang:2007:ERS

[KSW07] Wanmo Kang, Perwez Shahabuddin, and Ward Whitt. Exploiting regenerative structure to estimate finite time averages via simulation. ACM Transactions on Modeling and Computer Simulation, 17(2): ??, April 2007. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Kiatsupaibul:2011:AVH

[KSZ11] Seksan Kiatsupaibul, Robert L.

Smith, and Zelda B. Zabinsky. An analysis of a variation of hit-and-run for uniform sampling from general regions. ACM Transactions on Modeling and Computer Simulation, 21(3):16:1–16:??, March 2011. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Kunnumkal:2010:SAM

[KT10] Sumit Kunnumkal and Huseyin Topaloglu. A stochastic approximation method with max-norm projections and its applications to the Q-learning algorithm. ACM Transactions on Modeling and Computer Simulation, 20(3):12:1– 12:??, September 2010. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Kielanski:2023:PAW

[KV23] Grzegorz Kielanski and Benny Van Houdt. Performance analysis of work stealing strategies in large-scale multithreaded computing. ACM Transactions on Modeling and Computer Simulation, 33(4): 15:1–15:??, October 2023. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3584186.

Kesidis:1993:QSA

[KW93] G. Kesidis and J. Walrand. Quick simulation of ATM buffers with on-off multiclass Markov fluid sources. ACM Transactions on Modeling and Computer Simulation, 3(3): 269–276, July 1993. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Kim:2015:PAK

[KW15] Song-Hee Kim and Ward Whitt. The power of alternative Kolmogorov–Smirnov tests based on transformations of the data. ACM Transactions on Modeling and Computer Simulation, 25(4):24:1– 24:??, November 2015. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Koster:2022:GFS

[KWU22] Till Köster, Tom Warnke, and Adelinde M. Uhrmacher. Generating fast specialized simulators for stochastic reaction networks via partial evaluation. ACM Transactions on Modeling and Computer Simulation, 32(2):10:1–10:25, April 2022. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL

https://dl.acm.org/doi/
10.1145/3485465.

Koh:2011:MSP

[KZ11] Wee Lit Koh and Suiping Zhou. Modeling and simulation of pedestrian behaviors in crowded places. ACM Transactions on Modeling and Computer Simulation, 21(3): 20:1–20:??, March 2011. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Lopez-Ardao:2000:USS

[LALGSG^{+}00] José C. López-Ardao, Cándido López-García, Andrés Suárez-González, Manuel Fernández-Veiga, and Raúl Rodríguez-Rubio. On the use of selfsimilar processes in network simulation. ACM Transactions on Modeling and Computer Simulation, 10(2): 125–151, April 2000. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Love:2015:OBA

[LB15] David Love and Güzin Bayraksan. Overlapping batches for the assessment of solution quality in stochastic programs. ACM Transactions on Modeling and Computer Simulation, 25(3):20:1–20:??, April 2015. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

LEcuyer:1993:SGM

[LBC93] Pierre L’Ecuyer, François Blouin, and Raymond Couture. A search for good multiple recursive random number generators. ACM Transactions on Modeling and Computer Simulation, 3(2):87–98, April 1993. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Linden:2019:EIP

[LBEJ19] Jonatan Lindén, Pavol Bauer, Stefan Engblom, and Bengt Jonsson. Exposing interprocess information for efficient PDES of spatial stochastic systems on multicores. ACM Transactions on Modeling and Computer Simulation, 29(2):11:1–11:??, April 2019. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3301500.

Li:2001:APF

[LBL01] Na Li, Marissa Borrego, and San-Qi Li. Achieving perflow fair rate allocation in Diffserv. ACM Transactions on Modeling and Computer Simulation, 11(2):161–181, April 2001. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Lamps:2018:TIE

[LBN^{+}18] Jereme Lamps, Vignesh Babu, David M. Nicol, Vladimir Adam, and Rakesh Kumar. Temporal integration of emulation and network simulators on Linux multiprocessors. ACM Transactions on Modeling and Computer Simulation, 28(1):1:1–1:??, January 2018. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

LEcuyer:2010:ARE

[LBTG10] Pierre L’Ecuyer, Jose H.

Blanchet, Bruno Tuffin, and Peter W. Glynn. Asymptotic robustness of estimators in rare-event simulation. ACM Transactions on Modeling and Computer Simulation, 20(1): 6:1–6:41, January 2010. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

LEcuyer:2001:ESC

[LC01] Pierre L’Ecuyer and Yanick Champoux. Estimating small cell-loss ratios in ATM switches via importance sampling. ACM Transactions on Modeling and Computer Simulation, 11(1):76–105, January 2001. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Lathrop:2011:SPI

[LCK11] Scott Lathrop, Jim Costa, and William Kramer, editors. SC’11: Proceedings of 2011 International Conference for High Performance Computing, Networking, Storage and Analysis, Seattle, WA, November 12–18 2011. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2011. ISBN 1-4503-0771-X. LCCN ????

Lu:2016:RTC

[LCL16] Guanghui Lu, Leiting Chen, and Weiping Luo. Real-time crowd simulation integrating potential fields and agent method. ACM Transactions on Modeling and Computer Simulation, 26(4):28:1–28:??, May 2016. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Lewandowski:2007:SBE

[LCT07] Daniel Lewandowski, Roger M. Cooke, and Radboud J. Duintjer Tebbens. Sample-based estimation of correlation ratio with polynomial approximation. ACM Transactions on Modeling and Computer Simulation, 18(1):3:1–3:17, December 2007. CODEN ATMCEZ.

ISSN 1049-3301 (print), 1558-1195 (electronic).

Li:2015:ARP

[LCT^{+}15] Zengxiang Li, Wentong Cai, Stephen John Turner, Xiaorong Li, Ta Nguyen Binh Duong, and Rick Siow Mong Goh. Adaptive resource provisioning mechanism in VEEs for improving performance of HLA-based simulations. ACM Transactions on Modeling and Computer Simulation, 26(1): 1:1–1:??, December 2015. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Li:2017:CAB

[LCT17] Xiaosong Li, Wentong Cai, and Stephen J. Turner. Cloning agent-based simulation. ACM Transactions on Modeling and Computer Simulation, 27(2):15:1–15:??, July 2017. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Lomow:1991:MUI

[LDF91] Greg Lomow, Samir Ranjan Das, and Richard M. Fujimoto. Mechanisms for userinvoked retraction of events in time warp. ACM Transactions on Modeling and Computer Simulation, 1(3): 219–243, July 1991. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Lu:2004:MTM

[LDL04] Quan Lu, Maged Dessouky, and Robert C. Leachman. Modeling train movements through complex rail networks. ACM Transactions on Modeling and Computer Simulation, 14(1):48–75, January 2004. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Ledeczi:2003:MMI

[LDNA03] Akos Ledeczi, James Davis, Sandeep Neema, and Aditya Agrawal. Modeling methodology for integrated simulation of embedded systems. ACM Transactions on Modeling and Computer Simulation, 13(1): 82–103, January 2003. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

LEcuyer:2007:RES

[LDT07] Pierre L’Ecuyer, Valérie Demers, and Bruno Tuffin. Rare events, splitting, and quasi-Monte Carlo. ACM Transactions on Modeling and Computer Simulation, 17(2):??, April 2007. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

LEcuyer:2003:GI

[L’E03] Pierre L’Ecuyer. Guest introduction. ACM Transactions on Modeling and Computer Simulation, 13(4):295– 298, October 2003. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Lemire:2019:FRI

[Lem19] Daniel Lemire. Fast random integer generation in an interval. ACM Transactions on Modeling and Computer Simulation, 29(1):3:1–3:12, February 2019. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). See replication report [Qua19].

Levin:2001:SIC

[Lev01] Mordechay B. Levin. On the statistical independence of compound pseudorandom numbers over part of the period. ACM Transactions on Modeling and Computer Simulation, 11(3):294–311, July 2001. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Leydold:1998:RTS

[Ley98]  Josef Leydold.  A rejec-

tion technique for sampling from log-concave multivariate distributions. ACM Transactions on Modeling and Computer Simulation, 8(3): 254–280, July 1998. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Lee:1999:ORM

[LF99] Kangsun Lee and Paul A.

Fishwick. OOPM/RT: a multimodeling methodology for real-time simulation. ACM Transactions on Modeling and Computer Simulation, 9(2): 141–170, April 1999. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

LeCorff:2013:CPB

[LF13] Sylvain Le Corff and Gersende Fort. Convergence of a particle-based approximation of the block online expectation maximization algorithm. ACM Transactions on Modeling and Computer Simulation, 23(1):2:1–2:??, January 2013. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Lee:2003:CDF

[LG03] Shing-Hoi Lee and Peter W. Glynn. Computing the distribution function of a conditional expectation via Monte Carlo: Discrete conditioning spaces. ACM Transactions on Modeling and Computer Simulation, 13(3):238–258, July 2003. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Lan:2002:RMP

[LH02] Kun-Chan Lan and John Heidemann. Rapid model parameterization from traffic measurements. ACM Transactions on Modeling and Computer Simulation, 12(3):201–229, July 2002. CODEN ATMCEZ.

ISSN 1049-3301 (print), 1558-1195 (electronic).

Liu:2017:MSE

[LHJS17] Ning Liu, Adnan Haider, Dong Jin, and Xian-He Sun. Modeling and simulation of extreme-scale fat-tree networks for HPC systems and data centers. ACM Transactions on Modeling and Computer Simulation, 27(2):13:1– 13:??, July 2017. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Lim:2012:SAM

[Lim12] Eunji Lim. Stochastic approximation over multidimensional discrete sets with applications to inventory systems and admission control of queueing networks. ACM Transactions on Modeling and Computer Simulation, 22(4):19:1– 19:??, November 2012. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Lin:1992:PAP

[Lin92]  Yi-Bing Lin.  Parallelism

analyzers for parallel discrete event simulation. ACM Transactions on Modeling and Computer Simulation, 2(3): 239–264, July 1992. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Lee:2022:HDM

[LJS22] Seunghan Lee, Saurabh Jain, and Young-Jun Son. A hierarchical decision-making framework in social networks for efficient disaster management. ACM Transactions on Modeling and Computer Simulation, 32(1):5:1–5:26, January 2022. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3490027.

Liu:2021:ISI

[LK21] Jason Liu and Laxmikant Kale. Introduction to the special issue on PADS 2019. ACM Transactions on Modeling and Computer Simulation, 31(2):9e:1–9e:2, April 2021. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3451235.

Lin:1991:STW

[LL91a] Yi-Bing Lin and Edward D. Lazowska. A study of time warp rollback mechanisms. ACM Transactions on Modeling and Computer Simulation, 1(1):51–72, January 1991. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Lin:1991:TDA

[LL91b] Yi-Bing Lin and Edward D. Lazowska. A time-division algorithm for parallel simulation. ACM Transactions on Modeling and Computer Simulation, 1(1):73–83, January 1991. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Liu:2002:CBA

[LL02] Jie Liu and Edward A. Lee.

A component-based approach to modeling and simulating mixed-signal and hybrid systems. ACM Transactions on Modeling and Computer Simulation, 12(4):343–368, October 2002. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Li:2015:CBS

[LL15] Ting Li and Jason Liu.

Cluster-based spatiotemporal background traffic generation for network simulation. ACM Transactions on Modeling and Computer Simulation, 25(1): 4:1–4:??, January 2015. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Lam:2020:PSO

[LL20] Henry Lam and Fengpei Li.

Parametric scenario optimization under limited data: a distributionally robust optimization view. ACM Transactions on Modeling and Computer Simulation, 30(4): 21:1–21:41, December 2020. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3410152.

Lo:2013:OPB

[LLCC13] Shih-Hsiang Lo, Che-Rung Lee, I-Hsin Chung, and Yeh-Ching Chung. Optimizing pairwise box intersection checking on GPUs for large-scale simulations. ACM Transactions on Modeling and Computer Simulation, 23(3): 19:1–19:??, July 2013. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Lu:2000:SLS

[LLHL00] Tainchi Lu, Chungnan Lee, Wenyang Hsia, and Mingtang Lin. Supporting large-scale distributed simulation using HLA. ACM Transactions on Modeling and Computer Simulation, 10(3):268–294, July 2000. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Lees:2007:DSA

[LLT07] Michael Lees, Brian Logan, and Georgios Theodoropoulos. Distributed simulation of agent-based systems with HLA. ACM Transactions on Modeling and Computer Simulation, 17(3):11:1–11:??, July 2007. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Lin:1994:EBE

[LM94] Yi-Bing Lin and Victor W. Mak. Eliminating the boundary effect of a large-scale personal communication service network simulation. ACM Transactions on Modeling and Computer Simulation, 4(2): 165–190, April 1994. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Lin:2018:VDE

[LN18] Yujing Lin and Barry L. Nelson. Variance and derivative estimation of virtual performance. ACM Transactions on Modeling and Computer Simulation, 28(3):17:1–17:??, August 2018. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Loreti:2018:RCR

[Lor18] Michele Loreti. Replicated Computations Results (RCR) report for “Mesoscopic Modelling of Pedestrian Movement using Carma and its Tools”. ACM Transactions on Modeling and Computer Simulation, 28(2):12:1–12:??, April 2018. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Loreti:2019:RCR

[Lor19]  Michele Loreti.  Repli-

cated Computations Results (RCR) report for “Statistical Abstraction for Multiscale Spatio-temporal Systems”. ACM Transactions on Modeling and Computer Simulation, 29(4):23:1–23:??, December 2019. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). See [MHS19].

Lin:1991:OMM

[LP91] Yi-Bing Lin and Bruno R. Preiss. Optimal memory management for time warp parallel simulation. ACM Transactions on Modeling and Computer Simulation, 1(4):283– 307, October 1991. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Liu:2004:SFM

[LPM^{+}04] Yong Liu, Francesco L. Presti, Vishal Misra, Donald F. Towsley, and Yu Gu. Scalable fluid models and simulations for large-scale IP networks. ACM Transactions on Modeling and Computer Simulation, 14(3):305–324, July 2004. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Liao:2013:MBL

[LPPP13] Wei-Cherng Liao, Fragkiskos Papadopoulos, Konstantinos Psounis, and Constantinos Psomas. Modeling BitTorrentlike systems with many classes of users. ACM Transactions on Modeling and Computer Simulation, 23(2):13:1–13:??, May 2013. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Lee:1992:MSB

[LS92] Yuh-Jeng Lee and James F.

Stascavage. Multitasking simulation of a boiler system using qualitative model-based reasoning. ACM Transactions on Modeling and Computer Simulation, 2(4):285–306, October 1992. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Lee:2010:IHD

[LSJ10] Seungho Lee, Young-Jun Son, and Judy Jin. An integrated human decision making model for evacuation scenarios under a BDI framework. ACM Transactions on Modeling and Computer Simulation, 20(4): 23:1–23:??, October 2010. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Lubachevsky:1991:ARB

[LSW91] Boris Lubachevsky, Adam Schwartz, and Alan Weiss. An analysis of rollback-based simulation. ACM Transactions on Modeling and Computer Simulation, 1(2):154–193, April 1991. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Liu:2014:STM

[LT14] Elvis S. Liu and Georgios K.

Theodoropoulos. Space-time matching algorithms for interest management in distributed virtual environments. ACM Transactions on Modeling and Computer Simulation, 24(3): 15:1–15:??, June 2014. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Lin:2017:MSP

[LTM^{+}17] Zhongwei Lin, Carl Tropper, Robert A. McDougal, Mohammand Nazrul Ishlam Patoary, William W. Lytton, Yiping Yao, and Michael L. Hines. Multithreaded stochastic PDES for reactions and diffusions in neurons. ACM Transactions on Modeling and Computer Simulation, 27(2): 7:1–7:??, July 2017. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Luck:2016:RCR

[Lüc16]  Alexander Lück.  Repli-

cated computational results (RCR) report for “Automatic Moment-Closure Approximation of Spatially Distributed Collective Adaptive Systems”. ACM Transactions on Modeling and Computer Simulation, 26(4):27:1–27:??, May 2016. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). See [FHG16].

Lassila:2000:NOI

[LV00] Pasi E. Lassila and Jorma T. Virtamo. Nearly optimal importance sampling for Monte Carlo simulation of loss systems. ACM Transactions on Modeling and Computer Simulation, 10(4):326–347, October 2000. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Lebeck:1997:AMN

[LW97a] Alvin R. Lebeck and David A.

Wood. Active memory: a new abstraction for memory system simulation. ACM Transactions on Modeling and Computer Simulation, 7(1):42–77, January 1997. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Leeb:1997:ILC

[LW97b] Hannes Leeb and Stefan Wegenkittl. Inversive and linear congruential pseudorandom number generators in empirical tests. ACM Transactions on Modeling and Computer Simulation, 7(2): 272–286, April 1997. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Liu:2014:ESE

[LX14] Jingchen Liu and Gongjun Xu.

Efficient simulations for the exponential integrals of Hölder continuous Gaussian random fields. ACM Transactions on Modeling and Computer Simulation, 24(2):9:1–9:??, February 2014. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Li:2020:DAM

[LZ20] D. Li and J. Zhong. Dimensionally aware multiobjective genetic programming for automatic crowd behavior modeling. ACM Transactions on Modeling and Computer Simulation, 30(3): 19:1–19:24, July 2020. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3391407.

Li:2016:MUN

[LZW16] Jie Li, Jianliang Zheng, and Paula Whitlock. MaD0: an ultrafast nonlinear pseudorandom number generator. ACM Transactions on Modeling and Computer Simulation, 26(2): 13:1–13:??, January 2016. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Marsaglia:2003:XR

[Mar03] George Marsaglia. Xorshift RNGs. Journal of Statistical Software, 8(14):1–6, 2003. CODEN JSSOBK. ISSN 1548-

7660.  URL http://www.
jstatsoft.org/v08/i14;
http://www.jstatsoft.org/
v08/i14/xorshift.pdf. See

[Bre04] for corrections and the equivalence of xorshift generators and the well-understood linear feedback shift register generators. See also [SMDS11, SM12, SLF14] for the failure of Marsaglia’s xorwow() generator from this paper. See [PL05, Vig16] for detailed analysis.

Marotta:2022:RRL

[Mar22] Romolo Marotta. RCR report of “A Language for Agent-Based Discrete-Event Modeling and Simulation of Linked Lives”. ACM Transactions on Modeling and Computer Simulation, 32(1):7:1–7:4, January 2022. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL

https://dl.acm.org/doi/
10.1145/3490030.

Matsumoto:1998:SCA

[Mat98] Makoto Matsumoto. Simple cellular automata as pseudorandom m-sequence generators for built-in self-test. ACM Transactions on Modeling and Computer Simulation, 8(1): 31–42, January 1998. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Matloff:2005:EIF

[Mat05] Norman Matloff. Estimation of Internet file-access/ modification rates from indirect data. ACM Transactions on Modeling and Computer Simulation, 15(3):233–253, July 2005. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

McClean:2011:MFC

[MBGF11] Sally McClean, Maria Barton, Lalit Garg, and Ken Fullerton. A modeling framework that combines Markov models and discrete-event simulation for stroke patient care. ACM Transactions on Modeling and Computer Simulation, 21(4): 25:1–25:??, August 2011. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Mellor:2011:IHS

[MCC11] Georgina R. Mellor, Christine S. M. Currie, and Elizabeth L. Corbett. Incorporating household structure into a discreteevent simulation model of tuberculosis and HIV. ACM Transactions on Modeling and Computer Simulation, 21(4): 26:1–26:??, August 2011. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Marzolla:2020:PDD

[MD20] Moreno Marzolla and Gabriele D’Angelo. Parallel data distribution management on shared-memory multiprocessors. ACM Transactions on Modeling and Computer Simulation, 30(1):5:1–5:25, February 2020. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3369759.

Mascart:2023:ESS

[MDH^{+}23] Cyrille Mascart, David, Hill, Alexandre Muzy, and Patricia Reynaud-Bouret. Efficient simulation of sparse graphs of point processes. ACM Transactions on Modeling and Computer Simulation, 33(1–2):1:1–1:??, April 2023. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3565809.

Miller:1992:AWS

[MFFR92] David P. Miller, R. James Firby, Paul A. Fishwick, and Jeff Rothenberg. AI: what simulationists really need to know. ACM Transactions on Modeling and Computer Simulation, 2(4):269–284, October 1992. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Madisetti:1992:SMD

[MH92] Vijay K. Madisetti and David A. Hardaker. Synchronization mechanisms for distributed event-driven computation. ACM Transactions on Modeling and Computer Simulation, 2(1):12–50, January 1992. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Ma:2019:PSB

[MH19] Sijia Ma and Shane G. Henderson. Predicting the simulation budget in ranking and selection procedures. ACM Transactions on Modeling and Computer Simulation, 29(3):14:1–14:??, July 2019.

CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3323715.

Michaelides:2019:SAM

[MHS19] Michalis Michaelides, Jane Hillston, and Guido Sanguinetti. Statistical abstraction for multi-scale spatiotemporal systems. ACM Transactions on Modeling and Computer Simulation, 29(4): 22:1–22:??, December 2019. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). See replication report [Lor19].

Moka:2015:RSQ

[MJ15] Sarat Babu Moka and Sandeep Juneja. Regenerative simulation for queueing networks with exponential or heavier tail arrival distributions. ACM Transactions on Modeling and Computer Simulation, 25(4): 22:1–22:??, November 2015. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Mniszewski:2015:TDE

[MJV^{+}15] Susan M. Mniszewski, Christoph Junghans, Arthur F. Voter, Danny Perez, and Stephan J. Eidenbenz. TADSim: Discrete event-based performance prediction for temperatureaccelerated dynamics. ACM Transactions on Modeling and Computer Simulation, 25(3):

15:1–15:??, April 2015. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Matsumoto:1992:TGG

[MK92] Makoto Matsumoto and Yoshiharu Kurita. Twisted GFSR generators. ACM Transactions on Modeling and Computer Simulation, 2(3): 179–194, July 1992. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Matsumoto:1994:TGG

[MK94] Makoto Matsumoto and Yoshiharu Kurita. Twisted GFSR generators II. ACM Transactions on Modeling and Computer Simulation, 4(3): 254–266, July 1994. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Matsumoto:1996:SDR

[MK96] Makoto Matsumoto and Yoshiharu Kurita. Strong deviations from randomness in msequences based on trinomials. ACM Transactions on Modeling and Computer Simulation, 6(2):99–106, April 1996. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Malhotra:2017:PPS

[MKG^{+}17] Geetika Malhotra, Rajshekar Kalayappan, Seep Goel, Pooja Aggarwal, Abhishek Sagar, and Smruti R. Sarangi. ParTejas: a parallel simulator for multicore processors. ACM Transactions on Modeling and Computer Simulation, 27(3): 19:1–19:??, September 2017. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Mascarenhas:1998:MCA

[MKPR98] Edward Mascarenhas, Felipe Knop, Reuben Pasquini, and Vernon Rego. Minimum cost adaptive synchronization: experiments with the ParaSol system. ACM Transactions on Modeling and Computer Simulation, 8(4):401–430, October 1998. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Mustafee:2021:DAS

[MKT21] Navonil Mustafee, Korina Katsaliaki, and Simon J. E. Taylor. Distributed approaches to supply chain simulation: a review. ACM Transactions on Modeling and Computer Simulation, 31(4): 25:1–25:31, October 2021. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3466170.

Milenkovic:2007:ESP

[MM07] Aleksandar Milenković and Milena Milenković. An efficient single-pass trace compression technique utilizing instruction streams. ACM Transactions on Modeling and Computer Simulation, 17(1): ??, January 2007. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Martinez-Moyano:2008:BTI

[MMRC^{+}08] Ignacio J. Martinez-Moyano, Eliot Rich, Stephen Conrad, David F. Andersen, and Thomas R. Stewart. A behavioral theory of insider-threat risks: a system dynamics approach. ACM Transactions on Modeling and Computer Simulation, 18(2):7:1–7:??, April 2008. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Matsumoto:1998:MTD

[MN98] Makoto Matsumoto and Takuji Nishimura. Mersenne Twister: a 623-dimensionally equidistributed uniform pseudorandom number generator. ACM Transactions on Modeling and Computer Simulation, 8(1):3–30, January 1998. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL

http://www.math.keio.ac.
jp/~matsumoto/emt.html.

Martens:2006:FST

[MPK06] Jurgen Martens, Ferdi Put, and Etienne Kerre. A fuzzy set theoretic approach to validate simulation models. ACM Transactions on Modeling and Computer Simulation, 16(4):

375–398, October 2006. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Melamed:2004:HSH

[MPW04] Benjamin Melamed, Shuo Pan, and Yorai Wardi. HNS: a streamlined Hybrid Network Simulator. ACM Transactions on Modeling and Computer Simulation, 14(3):251–277, July 2004. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Mandjes:2002:LDA

[MR02] Michel Mandjes and Ad Ridder. A large deviations analysis of the transient of a queue with many Markov fluid inputs: approximations and fast simulation. ACM Transactions on Modeling and Computer Simulation, 12(1):1–26, January 2002. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Marin:2018:PFM

[MRB^{+}18] Andrea Marin, Sabina Rossi, Dario Burato, Andrea Sina, and Matteo Sottana. A product-form model for the performance evaluation of a bandwidth allocation strategy in WSNs. ACM Transactions on Modeling and Computer Simulation, 28(2):13:1–13:??, April 2018. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

McClary:2010:SAC

[MSK10] Daniel W. McClary, Violet R.

Syrotiuk, and Murat Kulahci. Steepest-ascent constrained simultaneous perturbation for multiobjective optimization. ACM Transactions on Modeling and Computer Simulation, 21(1):2:1–2:??, December 2010. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Miretskiy:2010:SDI

[MSM10] Denis Miretskiy, Werner Scheinhardt, and Michel Mandjes. State-dependent importance sampling for a Jackson tandem network. ACM Transactions on Modeling and Computer Simulation, 20(3): 15:1–15:??, September 2010. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Mustafee:2017:GET

[MST17] Navonil Mustafee, Young-Jun Son, and Simon J. E. Taylor. Guest editorial for the TOMACS special issue on the Principles of Advanced Discrete Simulation (PADS). ACM Transactions on Modeling and Computer Simulation, 27(2):7:1–7:??, July 2017. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Murdoch:2006:PSQ

[MT06] Duncan J. Murdoch and Glen Takahara. Perfect sampling for queues and network models. ACM Transactions on Modeling and Computer Simulation, 16(1):76–92, January 2006. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Mosterman:2002:GES

[MV02] Pieter J. Mosterman and Hans Vangheluwe. Guest editorial: Special issue on computer automated multi-paradigm modeling. ACM Transactions on Modeling and Computer Simulation, 12(4):249–255, October 2002. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Matsumoto:2007:CDI

[MWKA07] Makoto Matsumoto, Isaku Wada, Ai Kuramoto, and Hyo Ashihara. Common defects in initialization of pseudorandom number generators. ACM Transactions on Modeling and Computer Simulation, 17(4): 15:1–15:20, September 2007. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Madisetti:1991:AAP

[MWM91] Vijay K. Madisetti, Jean C. Walrand, and David G. Messerschmitt. Asynchronous algorithms for the parallel simulation of event-driven dynamical systems. ACM Transactions on Modeling and Computer Simulation, 1(3): 244–274, July 1991. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Mccoy:2007:MAN

[MWMD07] Aaron Mccoy, Tomas Ward, Seamus Mcloone, and Declan Delaney. Multistepahead neural-network predictors for network traffic reduction in distributed interactive applications. ACM Transactions on Modeling and Computer Simulation, 17(4):16:1– 16:??, September 2007. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Mitchell:2008:SAM

[MY08] Bradley Mitchell and Levent Yilmaz. Symbiotic adaptive multisimulation: an autonomic simulation framework for real-time decision support under uncertainty. ACM Transactions on Modeling and Computer Simulation, 19(1): 2:1–2:??, December 2008. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Marsaglia:1991:NCR

[MZ91] George Marsaglia and Arif Zaman. A new class of random number generators. Annals of Applied Probability, 1(3):462–480, August 1991. CODEN ???? ISSN 1050-5164. URL http://

projecteuclid.org/euclid.
aoap/1177005878. See pop-

ular description in [Pet91]. See remarks in [EH95, TLC93] about the extremely bad lattice structure in high dimensions of the generators proposed in this paper.

Marsaglia:1993:MTR

[MZ93] George Marsaglia and Arif Zaman. Monkey tests for random number generators. Computers and Mathematics with Applications, 26(9):1–10, November 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). See also [PW95].

Nakayama:1994:CSF

[Nak94] Marvin K. Nakayama. A characterization of the simple failure-biasing method for simulations of highly reliable Markovian Systems. ACM Transactions on Modeling and Computer Simulation, 4(1): 52–88, January 1994. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Nakayama:2014:CIQ

[Nak14] Marvin K. Nakayama. Confidence intervals for quantiles using sectioning when applying variance-reduction techniques. ACM Transactions on Modeling and Computer Simulation, 24(4):19:1–19:??, May 2014. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Nguyen:2021:TSN

[NAT^{+}21] Quang Anh Pham Nguyen, Philipp Andelfinger, Wen Jun Tan, Wentong Cai, and Alois Knoll. Transitioning spiking neural network simulators to heterogeneous hardware. ACM Transactions on Modeling and Computer Simulation, 31(2):9:1–9:26, April 2021. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3422389.

Nadoli:1993:IMS

[NB93] Gajanana Nadoli and John E.

Biegel. Intelligent Manufacturing-Simulation Agents Tool (IM-SAT). ACM Transactions on Modeling and Computer Simulation, 3(1):42–65, January 1993. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Ng:2006:RPU

[NC06] Szu Hui Ng and Stephen E.

Chick. Reducing parameter uncertainty for stochastic systems. ACM Transactions on Modeling and Computer Simulation, 16(1):26–51, January 2006. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Naing:2022:DDD

[NCN^{+}22] Htet Naing, Wentong Cai, Hu Nan, Wu Tiantian, and Yu Liang. Dynamic datadriven microscopic traffic simulation using jointly trained physics-guided long shortterm memory. ACM Transactions on Modeling and Computer Simulation, 32(4): 28:1–28:??, October 2022. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3558555.

North:2006:ECT

[NCV06] Michael J. North, Nicholson T. Collier, and Jerry R. Vos. Experiences creating three implementations of the Repast agent modeling toolkit. ACM Transactions on Modeling and Computer Simulation, 16(1):1–25, January 2006. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Nelson:1993:RMC

[Nel93] Barry L. Nelson. Robust multiple comparisons under common random numbers. ACM Transactions on Modeling and Computer Simulation, 3(3): 225–243, July 1993. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Nelson:2017:RCR

[Nel17] Barry L. Nelson. Replicated computations results (RCR) report for “Green Simulation: Reusing the Output of Repeated Experiments”. ACM Transactions on Modeling and Computer Simulation, 27(4): 24:1–24:??, December 2017. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). See [FS17].

Nicol:1995:CSP

[NH95] David M. Nicol and Philip Heidelberger. A comparative study of parallel algorithms for simulating continuous time Markov chains. ACM Transactions on Modeling and Computer Simulation, 5(4): 326–354, October 1995. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Nicol:1996:PES

[NH96] David Nicol and Philip Heidelberger. Parallel execution for serial simulators. ACM Transactions on Modeling and Computer Simulation, 6(3): 210–242, July 1996. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Ni:2015:HHS

[NH15] Eric C. Ni and Shane G. Henderson. How hard are steadystate queueing simulations? ACM Transactions on Modeling and Computer Simulation, 25(4):27:1–27:??, November 2015. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Nicol:1991:PBP

[Nic91]  David M. Nicol.  Perfor-

mance bounds on parallel selfinitiating discrete-event simulations. ACM Transactions on Modeling and Computer Simulation, 1(1):24–50, January 1991. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Nicol:1997:E

[Nic97] David Nicol. Editorial. ACM Transactions on Modeling and Computer Simulation, 7(4): 424, October 1997. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Nicol:2004:E

[Nic04] David Nicol. Editorial. ACM Transactions on Modeling and Computer Simulation, 14 (2):115, April 2004. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Nicol:2008:ESI

[Nic08] David M. Nicol. Efficient simulation of Internet worms. ACM Transactions on Modeling and Computer Simulation, 18(2):5:1–5:??, April 2008. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Niederreiter:1994:PVG

[Nie94] Harald Niederreiter. Pseudorandom vector generation by the inversive method. ACM Transactions on Modeling and Computer Simulation, 4(2): 191–212, April 1994. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Nishimura:2000:TBM

[Nis00] Takuji Nishimura. Tables of 64-bit Mersenne twisters. ACM Transactions on Modeling and Computer Simulation, 10(4):348–357, October 2000. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Nzouonta:2011:DIM

[NNB11] Josiane Nzouonta, Marvin K.

Nakayama, and Cristian Borcea. On deriving and incorporating multihop path duration estimates in VANET protocols. ACM Transactions on Modeling and Computer Simulation, 21(2):14:1– 14:??, February 2011. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Nance:1999:RMS

[NOP99] Richard E. Nance, C. Michael Overstreet, and Ernest H. Page. Redundancy in model specifications for discrete event simulation. ACM Transactions on Modeling and Computer Simulation, 9(3): 254–281, July 1999. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Nelson:2006:CVS

[NS06] Barry L. Nelson and Jeremy Staum. Control variates for screening, selection, and estimation of the best. ACM Transactions on Modeling and Computer Simulation, 16(1): 52–75, January 2006. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Nutaro:2006:DEM

[Nut06] James Nutaro. A discrete event method for wave simulation. ACM Transactions on Modeling and Computer Simulation, 16(2):174–195, April 2006. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Nutaro:2008:COS

[Nut08] James Nutaro. On constructing optimistic simulation algorithms for the discrete event system specification. ACM Transactions on Modeling and Computer Simulation, 19(1): 1:1–1:??, December 2008. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Nutaro:2020:TTS

[Nut20] James Nutaro. Toward a theory of superdense time in simulation models. ACM Transactions on Modeling and Computer Simulation, 30(3):

16:1–16:13, July 2020. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3379489.

Nicol:2004:DEF

[NY04] David M. Nicol and Guanhua Yan. Discrete event fluid modeling of background TCP traffic. ACM Transactions on Modeling and Computer Simulation, 14(3):211–250, July 2004. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Ng:2012:BKA

[NY12] Szu Hui Ng and Jun Yin.

Bayesian kriging analysis and design for stochastic simulations. ACM Transactions on Modeling and Computer Simulation, 22(3):17:1–17:??, August 2012. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Nicola:2007:EIS

[NZ07] Victor F. Nicola and Tatiana S. Zaburnenko. Efficient importance sampling heuristics for the simulation of population overflow in Jackson networks. ACM Transactions on Modeling and Computer Simulation, 17(2):??, April 2007. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Olstam:2008:FSS

[OLAM08] Johan Janson Olstam, Jan Lundgren, Mikael Adlers, and Pontus Matstoms. A framework for simulation of surrounding vehicles in driving simulators. ACM Transactions on Modeling and Computer Simulation, 18(3): 9:1–9:??, July 2008. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Osogami:2009:FPB

[Oso09] Takayuki Osogami. Finding probably best systems quickly via simulations. ACM Transactions on Modeling and Computer Simulation, 19(3): 12:1–12:??, June 2009. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Othman:2024:PSF

[OT24] Md. Shalihin Othman and Gary Tan. A prescriptive simulation framework with realistic behavioural modelling for emergency evacuations. ACM Transactions on Modeling and Computer Simulation, 34(1):4:1–4:??, January 2024. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3633330.

Owen:1998:LSS

[Owe98] Art B. Owen. Latin supercube sampling for very high-dimensional simulations. ACM Transactions on Modeling and Computer Simulation, 8(1):71–102, January 1998. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Owen:2003:VAS

[Owe03] Art B. Owen. Variance with alternative scramblings of digital nets. ACM Transactions on Modeling and Computer Simulation, 13(4):363– 378, October 2003. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Owen:2013:BES

[Owe13] Art B. Owen. Better estimation of small Sobol’ sensitivity indices. ACM Transactions on Modeling and Computer Simulation, 23(2):11:1–11:??, May 2013. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Pachoulakis:2008:RVA

[Pac08] Ioannis Pachoulakis. 3D reconstruction and visualization of astrophysical wind volumes using physical models. ACM Transactions on Modeling and Computer Simulation, 18(4): 14:1–14:??, September 2008. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Page:1993:DDE

[Pag93] Ernest H. Page. In defense of discrete-event simulation. ACM Transactions on Modeling and Computer Simulation, 3(4):281–286, October 1993. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Parker:2018:RCR

[Par18] David Parker. Replicated Computational Results (RCR) report for “ProPPA: Probabilistic Programming for Stochastic Dynamical Systems”. ACM Transactions on Modeling and Computer Simulation, 28(1):4:1–4:??, January 2018. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Park:1996:SEL

[PB96] Taeshin Park and Paul I. Barton. State event location in differential-algebraic models. ACM Transactions on Modeling and Computer Simulation, 6(2):137–165, April 1996. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Powell:2011:ERD

[PBAB^{+}11] Warren B. Powell, Belgacem Bouzaiene-Ayari, Jean Berger, Abdeslem Boukhtouta, and Abraham P. George. The effect of robust decisions on the cost of uncertainty in military airlift operations. ACM Transactions on Modeling and Computer Simulation, 22(1): 1:1–1:??, December 2011. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Prabuchandran:2016:ACA

[PBB16] K. J. Prabuchandran, Shalabh Bhatnagar, and Vivek S. Borkar. Actor–critic algorithms with online feature adaptation. ACM Transactions on Modeling and Computer Simulation, 26(4):24:1– 24:??, May 2016. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Page:2000:WBS

[PBF^{+}00] Ernest H. Page, Arnold Buss, Paul A. Fishwick, Kevin J. Healy, Richard E. Nance, and Ray J. Paul. Web-based simulation: revolution or evolution? ACM Transactions on Modeling and Computer Simulation, 10(1):3–17, January 2000. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Plagge:2018:NMP

[PCGM18] Mark Plagge, Christopher D.

Carothers, Elsa Gonsiorowski, and Neil Mcglohon. NeMo: a massively parallel discreteevent simulation model for neuromorphic architectures. ACM Transactions on Modeling and Computer Simulation, 28(4):30:1–30:??, October 2018. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Page:1997:CSV

[PCT97] Ernest H. Page, Bradford S. Canova, and John A. Tufarolo. A case study of verification, validation, and accreditation for advanced distributed simulation. ACM Transactions on Modeling and Computer Simulation, 7(3): 393–424, July 1997. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Parker:2011:DPG

[PE11] Jon Parker and Joshua M. Epstein. A distributed platform for global-scale agent-based models of disease transmission. ACM Transactions on Modeling and Computer Simulation, 22(1):2:1–2:??, December 2011. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Pellegrini:2021:RCR

[Pel21] Alessandro Pellegrini. Replication of computational results report for “Green Simulation with Database Monte Carlo”. ACM Transactions on Modeling and Computer Simulation, 31(1):5:1–5:4, February 2021. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3426823. See [FS21].

Peterson:1991:NRN

[Pet91]  Ivars Peterson.  Numbers

at random: Number theory supplies a superior randomnumber generator. Science News (Washington, DC), 140 (19):300–301, November 9, 1991. CODEN SCNEBK. ISSN 0036-8423 (print), 1943-0930 (electronic). URL http:

//www.jstor.org/stable/
3975915.

Park:2011:AQN

[PF11] Hyungwook Park and Paul A.

Fishwick. An analysis of queuing network simulation using GPU-based hardware acceleration. ACM Transactions on Modeling and Computer Simulation, 21(3):18:1–18:??, March 2011. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Phan:2014:TSS

[PG14] Dzung Phan and Soumyadip Ghosh. Two-stage stochastic optimization for optimal power flow under renewable generation uncertainty. ACM Transactions on Modeling and Computer Simulation, 24(1): 2:1–2:??, January 2014. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Piho:2021:FAB

[PH21] Paul Piho and Jane Hillston.

Fluid approximation-based analysis for mode-switching population dynamics. ACM Transactions on Modeling and Computer Simulation, 31(2):8:1–8:26, April 2021. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3441680.

Pasupathy:2015:SCR

[PHP^{+}15] Raghu Pasupathy, Susan R. Hunter, Nugroho A. Pujowidianto, Loo Hay Lee, and Chun-Hung Chen. Stochastically constrained ranking and selection via SCORE. ACM Transactions on Modeling and Computer Simulation, 25(1): 1:1–1:??, January 2015. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Piccione:2024:RRA

[Pic24]  Andrea Piccione.  Repro-

ducibility report for the article: Parallel simulation of quantum networks with distributed quantum state management. ACM Transactions on Modeling and Computer Simulation, 34(2):12:1– 12:??, April 2024. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/

3639704. See [WKC^{+}24].

Plesser:2010:RSI

[PJ10] Hans Ekkehard Plesser and Anders Grønvik Jahnsen. Re-seeding invalidates tests of random number generators. Applied Mathematics and Computation, 217 (1):339–346, September 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300310006259.  See

[KCK08].

Pasupathy:2011:SRF

[PK11] Raghu Pasupathy and Sujin Kim. The stochastic rootfinding problem: Overview, solutions, and open questions. ACM Transactions on Modeling and Computer Simulation, 21(3):19:1–19:??, March 2011. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Panneton:2005:XRN

[PL05] François Panneton and Pierre L’Ecuyer. On the xorshift random number generators. ACM Transactions on Modeling and Computer Simulation, 15(4): 346–361, October 2005. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). See [Mar03, Bre04, Vig16].

Preiss:1994:ECI

[PLM94] Bruno R. Preiss, Wayne M.

Loucks, and Ian D. Macintyre. Effects of the checkpoint interval on time and space in time warp. ACM Transactions on Modeling and Computer Simulation, 4(3): 223–253, July 1994. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Pichitlamken:2003:CPO

[PN03] Juta Pichitlamken and Barry L. Nelson. A combined procedure for optimization via simulation. ACM Transactions on Modeling and Computer Simulation, 13(2):155–179, April 2003. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Perumalla:2013:RSE

[PP13] Kalyan S. Perumalla and Vladimir A. Protopopescu. Reversible simulations of elastic collisions. ACM Transactions on Modeling and Computer Simulation, 23(2):12:1– 12:??, May 2013. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Perumalla:2014:DEE

[PPT14] Kalyan S. Perumalla, Alfred J. Park, and Vinod Tipparaju. Discrete event execution with one-sided and two-sided GVT algorithms on 216,000 processor cores. ACM Transactions on Modeling and Computer Simulation, 24(3):16:1–16:??, June 2014. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Pellegrini:2017:FGT

[PQ17] Alessandro Pellegrini and Francesco Quaglia. A finegrain time-sharing time warp system. ACM Transactions on Modeling and Computer Simulation, 27(2):10:1–10:??, July 2017. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Petkov:2013:CPA

[PRO13] Vladislav Petkov, Ram Rajagopal, and Katia Obraczka. Characterizing per-application network traffic using entropy. ACM Transactions on Modeling and Computer Simulation, 23(2):14:1–14:??, May 2013. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Pasupathy:2009:RAA

[PS09] Raghu Pasupathy and Bruce W.

Schmeiser. Retrospectiveapproximation algorithms for the multidimensional stochastic root-finding problem. ACM Transactions on Modeling and Computer Simulation, 19(2): 5:1–5:??, March 2009. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Park:2000:PEM

[PT00] Kihong Park and Tsunyi Tuan. Performance evaluation of multiple time scale TCP under self-similar traffic conditions. ACM Transactions on Modeling and Computer Simulation, 10(2):152–177, April 2000. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Pan:2011:DSB

[PTCL11] Ke Pan, Stephen John Turner, Wentong Cai, and Zengxiang Li. A dynamic sortbased DDM matching algorithm for HLA applications. ACM Transactions on Modeling and Computer Simulation, 21(3):17:1–17:??, March 2011. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Principe:2020:DSM

[PTD^{+}20] Matteo Principe, Tommaso Tocci, Pierangelo Di Sanzo, Francesco Quaglia, and Alessandro Pellegrini. A distributed shared memory middleware for speculative parallel discrete event simulation. ACM Transactions on Modeling and Computer Simulation, 30(2): 11:1–11:26, April 2020. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3373335.

Puzis:2011:DSS

[PTE^{+}11] Rami Puzis, Meytal Tubi, Yuval Elovici, Chanan Glezer, and Shlomi Dolev. A decision support system for placement of intrusion detection and prevention devices in large-scale networks. ACM Transactions on Modeling and Computer Simulation, 22(1):5:1– 5:??, December 2011. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Percus:1995:TAM

[PW95] Ora E. Percus and Paula A.

Whitlock. Theory and application of Marsaglia’s monkey test for pseudorandom number generators. ACM Transactions on Modeling and Computer Simulation, 5(2):87–100, April 1995. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). See [MZ93].

Parker:2021:ISI

[PW21] David Parker and Verena Wolf. Introduction to the special issue on QEST 2019. ACM Transactions on Modeling and Computer Simulation, 31(3):12:1, July 2021. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3463764.

Quaglia:2002:PSP

[QC02] Francesco Quaglia and Vittorio Cortellessa. On the processor scheduling problem in time warp synchronization. ACM Transactions on Modeling and Computer Simulation, 12(3): 143–175, July 2002. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Qu:2021:RVG

[QDZ21] Yan Qu, Angelos Dassios, and Hongbiao Zhao. Random variate generation for exponential and gamma tilted stable distributions. ACM Transactions on Modeling and Computer Simulation, 31(4): 19:1–19:21, October 2021. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3449357.

Qu:2014:GES

[QF14] Huashuai Qu and Michael C. Fu. Gradient extrapolated stochastic kriging. ACM Transactions on Modeling and Computer Simulation, 24(4): 23:1–23:??, August 2014. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Quarles:2010:MRA

[QFL^{+}10] John Quarles, Paul Fishwick, Samsun Lampotang, Ira Fischler, and Benjamin Lok. A mixed reality approach for interactively blending dynamic models with corresponding physical phenomena. ACM Transactions on Modeling and Computer Simulation, 20(4): 22:1–22:??, October 2010. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Quaglia:2020:ESI

[QTP20] Francesco Quaglia, Georgios Theodoropoulos, and Alessandro Pellegrini. Editorial to the special issue on the Principles of Advanced Discrete Simulation (PADS). ACM Transactions on Modeling and Computer Simulation, 30(2): 8:1–8:2, April 2020. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3381903.

Quaglia:2019:RCR

[Qua19] Francesco Quaglia. Replicated computational results (RCR) report for “Fast Random Integer Generation in an Interval”. ACM Transactions on Modeling and Computer Simulation, 29(1):4:1–4:3, February 2019. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). See [Lem19].

Quaglia:2020:ENE

[Qua20] Francesco Quaglia. Editorial from the new Editor-in-Chief. ACM Transactions on Modeling and Computer Simulation, 30(1):1e:1, February 2020. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3377148.

Ronngren:1997:CSP

[RA97] Robert Rönngren and Rassul Ayani. A comparative study of parallel and sequential priority queue algorithms. ACM Transactions on Modeling and Computer Simulation, 7(2):157–209, April 1997. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Raatikainen:1993:SPS

[Raa93] Kimmo E. E. Raatikainen. A sequential procedure for simultaneous estimation of several means. ACM Transactions on Modeling and Computer Simulation, 3(2): 108–133, April 1993. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Riley:2004:FAD

[RAF^{+}04] George F. Riley, Mostafa H. Ammar, Richard M. Fujimoto, Alfred Park, Kalyan Perumalla, and Donghua Xu. A federated approach to distributed network simulation. ACM Transactions on Modeling and Computer Simulation, 14(2):116–148, April 2004. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Rahman:2019:PAP

[RAGN19] Shafiur Rahman, Nael Abu-Ghazaleh, and Walid Najjar. PDES-A: Accelerators for parallel discrete event simulation implemented on FPGAs. ACM Transactions on Modeling and Computer Simulation, 29(2):12:1–12:??, April 2019. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3302259.

Rohloff:2008:DSM

[RB08] Kurt R. Rohloff and Tamer Bacşar. Deterministic and stochastic models for the detection of random constant scanning worms. ACM Transactions on Modeling and Computer Simulation, 18(2): 8:1–8:??, April 2008. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Rosenblum:1997:USM

[RBDH97] Mendel Rosenblum, Edouard Bugnion, Scott Devine, and Stephen A. Herrod. Using the SimOS machine simulator to study complex computer systems. ACM Transactions on Modeling and Computer Simulation, 7(1):78–103, January 1997. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Robinson:2010:SCS

[RD10] William N. Robinson and Yi Ding. A survey of customization support in agentbased business process simulation tools. ACM Transactions on Modeling and Computer Simulation, 20(3):14:1– 14:??, September 2010. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Reijsbergen:2018:PZG

[RDSJ18] Daniël Reijsbergen, Pieter-Tjerk De Boer, Werner Scheinhardt, and Sandeep Juneja. Path-ZVA: General, efficient, and automated importance sampling for highly reliable Markovian systems. ACM Transactions on Modeling and Computer Simulation, 28(3):22:1–22:??, August 2018. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Reed:2000:IAD

[RFA00] John A. Reed, Gregory J. Follen, and Abdollah A. Afjeh. Improving the aircraft design process using Web-based modeling and simulation. ACM Transactions on Modeling and Computer Simulation, 10(1): 58–83, January 2000. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Rainville:2012:EOL

[RGTL12] François-Michel D. Rainville, Christian Gagné, Olivier Teytaud, and Denis Laurendeau. Evolutionary optimization of low-discrepancy sequences. ACM Transactions on Modeling and Computer Simulation, 22(2):9:1–9:25, March 2012. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Rao:2019:MPE

[RH19] Dhananjai M. Rao and Julius D. Higiro. Managing pending events in sequential and parallel simulations using three-tier heap and two-tier ladder queue. ACM Transactions on Modeling and Computer Simulation, 29(2):9:1–9:??, April 2019. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3265750.

Randhawa:2004:CIS

[RJ04] R. S. Randhawa and S. Juneja.

Combining importance sampling and temporal difference control variates to simulate Markov Chains. ACM Transactions on Modeling and Computer Simulation, 14(1):1–30, January 2004. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Rahman:2020:SSI

[RK20] Atiqur Rahman and Peter Kemper. Simulation study to identify the characteristics of Markov chain properties. ACM Transactions on Modeling and Computer Simulation, 30(2):9:1–9:26, April 2020. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3361744.

Ruess:2015:MBM

[RL15] Jakob Ruess and John Lygeros. Moment-based methods for parameter inference and experiment design for stochastic biochemical reaction networks. ACM Transactions on Modeling and Computer Simulation, 25(2):8:1–8:??, February 2015. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Rodriguez:2020:GPB

[RL20] Sergio Rodriguez and Michael Ludkovski. Generalized probabilistic bisection for stochastic root finding. ACM Transactions on Modeling and Computer Simulation, 30(1):2:1– 2:27, February 2020. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3355607.

Reppas:2016:ENE

[RLDH16] Andreas I. Reppas, Georgios Lolas, Andreas Deutsch, and Haralampos Hatzikirou. The extrinsic noise effect on lateral inhibition differentiation waves. ACM Transactions on Modeling and Computer Simulation, 26(3):19:1– 19:??, February 2016. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Ruiz-Martin:2021:DEM

[RMWLP21] Cristina Ruiz-Martin, Gabriel Wainer, and Adolfo Lopez-Paredes. Discrete-event modeling and simulation of diffusion processes in multiplex networks. ACM Transactions on Modeling and Computer Simulation, 31(1):6:1– 6:32, February 2021. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3434490.

Reynolds:1997:CMM

[RNS97] Paul F. Reynolds, Jr., Anand Natrajan, and Sudhir Srinivasan. Consistency maintenance in multiresolution simulation. ACM Transactions on Modeling and Computer Simulation, 7(3):368–392, July 1997. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Rosenfeld:2008:ABG

[Ros08] Simon Rosenfeld. Approximate bivariate gamma generator with prespecified correlation and different marginal shapes. ACM Transactions on Modeling and Computer Simulation, 18(4):16:1– 16:??, September 2008. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Rajasekaran:1993:FAG

[RR93] Sanguthevar Rajasekaran and Keith W. Ross. Fast algorithms for generating discrete random variates with changing distributions. ACM Transactions on Modeling and Computer Simulation, 3(1):1–19, January 1993. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Ramesh:2000:CBP

[RRP00] Sridhar Ramesh, George N. Rouskas, and Harry G. Perros. Computing blocking probabilities in multiclass wavelength routing networks. ACM Transactions on Modeling and Computer Simulation, 10(2): 87–103, April 2000. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Rao:2000:WBN

[RRW00] Dhananjai Madhava Rao, Radharamanan Radhakrishnan, and Philip A. Wilsey. Web-based network analysis and design. ACM Transactions on Modeling and Computer Simulation, 10(1):18–38, January 2000. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Radiya:1994:LBF

[RS94] Ashvin Radiya and Robert G. Sargent. A logic-based foundation of discrete event modeling and simulation. ACM Transactions on Modeling and Computer Simulation, 4(1):3–51, January 1994. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Roeder:2010:IMQ

[RS10] Theresa M. Roeder and Lee W. Schruben. Information models for queueing system simulation. ACM Transactions on Modeling and Computer Simulation, 20(2): 8:1–8:??, April 2010. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Roy:2021:ASH

[RSG21] Sudipendra Nath Roy, Bhavin J.

Shah, and Hasmukh Gajjar. Application of simulation in healthcare service operations: a review and research agenda. ACM Transactions on Modeling and Computer Simulation, 31(1):3:1–3:23, February 2021. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3427753.

Raczy:2005:SBD

[RTY05] C. Raczy, G. Tan, and J. Yu.

A sort-based DDM matching algorithm for HLA. ACM Transactions on Modeling and Computer Simulation, 15(1): 14–38, January 2005. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Rubinstein:2002:CER

[Rub02] Reuven Y. Rubinstein. Crossentropy and rare events for maximal cut and partition problems. ACM Transactions on Modeling and Computer Simulation, 12(1):27–53, January 2002. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Ross:1993:AOI

[RW93] Keith W. Ross and Jie Wang. Asymptotically optimal importance sampling for product-form queuing networks. ACM Transactions on Modeling and Computer Simulation, 3(3):244–268, July 1993. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Roberts:2007:DSM

[RWK^{+}07] Stephen Roberts, Lijun Wang, Robert Klein, Reid Ness, and Robert Dittus. Development of a simulation model of colorectal cancer. ACM Transactions on Modeling and Computer Simulation, 18(1):4:1– 4:30, December 2007. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Reinhardt:2022:LAB

[RWU22] Oliver Reinhardt, Tom Warnke, and Adelinde M. Uhrmacher. A language for agent-based discrete-event modeling and simulation of linked lives. ACM Transactions on Modeling and Computer Simulation, 32(1):6:1–6:26, January 2022. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3486634.

Sengul:2015:SSM

[SABF15] Cigdem Sengul, Mustafa Al-Bado, and Anja Feldmann. Site-specific models for realistic wireless network simulation. ACM Transactions on Modeling and Computer Simulation, 25(1):3:1–3:??, January 2015. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Sebastio:2018:HAC

[SALS18] Stefano Sebastio, Michele Amoretti, Alberto Lluch Lafuente, and Antonio Scala. A holistic approach for collaborative workload execution in volunteer clouds. ACM Transactions on Modeling and Computer Simulation, 28(2): 14:1–14:??, April 2018. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Sanchez:2020:DFM

[San20] Susan M. Sanchez. Data farming: Methods for the present, opportunities for the future. ACM Transactions on Modeling and Computer Simulation, 30(4):22:1–22:30, December 2020. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL

https://dl.acm.org/doi/
10.1145/3425398.

Stytz:2001:DMT

[SB01] Martin R. Stytz and Sheila B. Banks. The distributed mission training integrated threat environment system architecture and design. ACM Transactions on Modeling and Computer Simulation, 11(1):106– 133, January 2001. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Stopford:2008:FSS

[SC08] Benjamin Stopford and Steve Counsell. A framework for the simulation of structural software evolution. ACM Transactions on Modeling and Computer Simulation, 18(4):17:1– 17:??, September 2008. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Schruben:2010:SMA

[Sch10] Lee Schruben. Simulation modeling for analysis. ACM Transactions on Modeling and Computer Simulation, 20(1): 2:1–2:22 + 17 (online appendix), January 2010. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Schafer:2013:PAO

[Sch13] Christian Schäfer. Particle algorithms for optimization on binary spaces. ACM Transactions on Modeling and Computer Simulation, 23(1):8:1– 8:??, January 2013. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Singh:2013:BLN

[SCW13] Sumeetpal S. Singh, Nicolas Chopin, and Nick Whiteley. Bayesian learning of noisy Markov decision processes. ACM Transactions on Modeling and Computer Simulation, 23(1):4:1–4:??, January 2013. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Saltzman:2012:SMN

[SDLH12] Evan A. Saltzman, John H.

Drew, Lawrence M. Leemis, and Shane G. Henderson. Simulating multivariate nonhomogeneous Poisson processes using projections. ACM Transactions on Modeling and Computer Simulation, 22(3):15:1– 15:??, August 2012. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Sottile:2015:SAT

[SDZ^{+}15] Matthew Sottile, Jason Dagit, Deli Zhang, Gilbert Hendry, and Damian Dechev. Static analysis techniques for semiautomatic synthesis of message passing software skeletons. ACM Transactions on Modeling and Computer Simulation, 26(1):4:1–4:??, December 2015. CODEN ATMCEZ.

ISSN 1049-3301 (print), 1558-1195 (electronic).

Song:2010:CLI

[SF10] Yang Song and Yuguang Fang. Cross-layer interactions in multihop wireless sensor networks: a constrained queueing model. ACM Transactions on Modeling and Computer Simulation, 21(1):4:1–4:??, December 2010. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Schreck:2013:AEE

[SFM13] Amandine Schreck, Gersende Fort, and Eric Moulines. Adaptive equi-energy sampler: Convergence and illustration. ACM Transactions on Modeling and Computer Simulation, 23(1):5:1–5:??, January 2013. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Soule:1991:ECM

[SG91] Larry Soulé and Anoop Gupta. An evaluation of the Chandy-Misra-Bryant algorithm for digital logic simulation. ACM Transactions on Modeling and Computer Simulation, 1(4):308–347, October 1991. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Singham:2014:SSR

[Sin14] Dashi I. Singham. Selecting stopping rules for confidence interval procedures. ACM Transactions on Modeling and Computer Simulation, 24(3): 18:1–18:??, May 2014. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Sharma:2010:JCC

[SJSM10] Gaurav Sharma, Changhee Joo, Ness B. Shroff, and Ravi R. Mazumdar. Joint congestion control and distributed scheduling for throughput guarantees in wireless networks. ACM Transactions on Modeling and Computer Simulation, 21(1):5:1–5:??, December 2010. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Swisher:2003:DES

[SJY03] James R. Swisher, Sheldon H.

Jacobson, and Enver Yücesan. Discrete-event simulation optimization using ranking, selection, and multiple comparison procedures: a survey. ACM Transactions on Modeling and Computer Simulation, 13(2):134–154, April 2003. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Song:2023:BAV

[SK23] Chenxiao Song and Reiichiro Kawai. Batching adaptive variance reduction. ACM Transactions on Modeling and Computer Simulation, 33 (1–2):3:1–3:??, April 2023. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3573386.

Shorey:1997:IPL

[SKR97] Rajeev Shorey, Anurag Kumar, and Kiran M. Rege. Instability and performance limits of distributed simulators of feedforward queueing networks. ACM Transactions on Modeling and Computer Simulation, 7(2):210–238, April 1997. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Schormans:2001:HTA

[SLCP01] John Schormans, Enjie Liu, Laurie Cuthbert, and Jonathan Pitts. A hybrid technique for accelerated simulation of ATM networks and network elements. ACM Transactions on Modeling and Computer Simulation, 11(2):182–205, April 2001. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Steele:2014:FSP

[SLF14] Guy L. Steele, Jr., Doug Lea, and Christine H. Flood. Fast splittable pseudorandom number generators. ACM SIG-PLAN Notices, 49(10):453– 472, October 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Steiger:2005:ABM

[SLW^{+}05] Natalie M. Steiger, Emily K.

Lada, James R. Wilson, Jeffrey A. Joines, Christos Alexopoulos, and David Goldsman. ASAP3: a batch means procedure for steady-state simulation analysis. ACM Transactions on Modeling and Computer Simulation, 15(1): 39–73, January 2005. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Saito:2012:DCS

[SM12] Mutsuo Saito and Makoto Matsumoto. A deviation of CURAND: Standard pseudorandom number generator in CUDA for GPGPU. Slides presented at the Tenth International Conference on Monte Carlo and Quasi-Monte Carlo Methods in Scientific Computing, February 2012. URL http:/

/www.mcqmc2012.unsw.edu.
au/slides/MCQMC2012_Matsumoto.
pdf.

Salmon:2011:PRN

[SMDS11] John K. Salmon, Mark A.

Moraes, Ron O. Dror, and David E. Shaw. Parallel random numbers: as easy as 1, 2, 3. In Lathrop et al. [LCK11], pages 16:1–16:12. ISBN 1-4503-0771-X. LCCN ????

Shortle:2009:RCQ

[SMG09] John F. Shortle, Brian L. Mark, and Donald Gross. Reduction of closed queueing networks for efficient simulation. ACM Transactions on Modeling and Computer Simulation, 19(3):10:1–10:??, June 2009. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Schwaninger:2015:SOA

[SMI15] Clemens Arthur Schwaninger, Denis Menshykau, and Dagmar Iber. Simulating organogenesis: Algorithms for the image-based determination of displacement fields. ACM Transactions on Modeling and Computer Simulation, 25(2): 10:1–10:??, February 2015. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Salemi:2016:MLS

[SNS16] Peter Salemi, Barry L. Nelson, and Jeremy Staum. Moving least squares regression for high-dimensional stochastic simulation metamodeling. ACM Transactions on Modeling and Computer Simulation, 26(3):16:1–16:??, February 2016. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Seal:2011:RPD

[SP11] Sudip K. Seal and Kalyan S. Perumalla. Reversible parallel discrete event formulation of a TLM-based radio signal propagation model. ACM Transactions on Modeling and Computer Simulation, 22(1): 4:1–4:??, December 2011. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Stamos:2010:CST

[SPV^{+}10] Konstantinos Stamos, George Pallis, Athena Vakali, Dimitrios Katsaros, Antonis Sidiropoulos, and Yannis Manolopoulos. CDNsim: a simulation tool for content distribution networks. ACM Transactions on Modeling and Computer Simulation, 20(2):10:1–10:??, April 2010. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Su:2024:OBC

[SPYG24] Ziwei Su, Raghu Pasupathy, Yingchieh Yeh, and Peter Glynn. Overlapping batch confidence intervals on statistical functionals constructed from time series: Application to quantiles, optimization, and estimation. ACM Transactions on Modeling and Computer Simulation, 34 (2):10:1–10:??, April 2024. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3649437.

Santoro:2012:TOS

[SQ12] Andrea Santoro and Francesco Quaglia. Transparent optimistic synchronization in the high-level architecture via time-management conversion. ACM Transactions on Modeling and Computer Simulation, 22(4):21:1–21:??, November 2012. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Srinivasan:1998:ET

[SR98] Sudhir Srinivasan and Paul F. Reynolds, Jr. Elastic time. ACM Transactions on Modeling and Computer Simulation, 8(2):103–139, April 1998. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Seznec:2003:HUL

[SS03] André Seznec and Nicolas Sendrier. HAVEGE: a userlevel software heuristic for generating empirically strong random numbers. ACM Transactions on Modeling and Computer Simulation, 13(4): 334–346, October 2003. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Sanchez:2005:VLF

[SS05] Susan M. Sanchez and Paul J. Sanchez. Very large fractional factorial and central composite designs. ACM Transactions on Modeling and Computer Simulation, 15(4):362– 377, October 2005. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Strunz:2008:SFS

[SS08] Kai Strunz and Qianli Su.

Stochastic formulation of SPICE-type electronic circuit simulation with polynomial chaos. ACM Transactions on Modeling and Computer Simulation, 18(4):15:1– 15:??, September 2008. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Schruben:2014:DDS

[SS14] Lee W. Schruben and Dashi I.

Singham. Data-driven simulation of complex multidimensional time series. ACM Transactions on Modeling and Computer Simulation, 24(1): 5:1–5:??, January 2014. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Stoffers:2018:AMF

[SSDW18] Mirko Stoffers, Daniel Schemmel, Oscar Soria Dustmann, and Klaus Wehrle. On automated memoization in the field of simulation parameter studies. ACM Transactions on Modeling and Computer Simulation, 28(4):26:1–26:??, October 2018. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Schiavone:1997:TDI

[SSH97] Guy A. Schiavone, S. Sureshchandran, and Kenneth C. Hardis. Terrain database interoperability issues in training with distributed interactive simulation. ACM Transactions on Modeling and Computer Simulation, 7(3):332–367, July 1997. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Sanderson:1991:HSL

[SSRT91] D. P. Sanderson, R. Sharma, R. Rozin, and S. Treu. The hierarchical simulation language HSL: a versatile tool for process-oriented simulation. ACM Transactions on Modeling and Computer Simulation, 1(2):113–153, April 1991. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Solow:2021:NAF

[SSY21] Daniel Solow, Roberto Szechtman, and Enver Yücesan. Novel approaches to feasibility determination. ACM Transactions on Modeling and Computer Simulation, 31 (1):1:1–1:25, February 2021. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3426359.

Suchard:2013:MPS

[SSZ^{+}13] Marc A. Suchard, Shawn E. Simpson, Ivan Zorych, Patrick Ryan, and David Madigan. Massive parallelization of serial inference algorithms for a complex generalized linear model. ACM Transactions on Modeling and Computer Simulation, 23(1):10:1–10:??, January 2013. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Suryanarayanan:2013:SRQ

[ST13] Vinoth Suryanarayanan and Georgios Theodoropoulos. Synchronised range queries in distributed simulations of multiagent systems. ACM Transactions on Modeling and Computer Simulation, 23(4): 25:1–25:??, October 2013. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Szabo:2015:FWE

[ST15] Claudia Szabo and Yong Meng Teo. Formalization of weak emergence in multiagent systems. ACM Transactions on Modeling and Computer Simulation, 26(1):6:1–6:??, December 2015. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Sainudiin:2013:PER

[STHL13] Raazesh Sainudiin, Gloria Teng, Jennifer Harlow, and Dominic Lee. Posterior expectation of regularly paved random histograms. ACM Transactions on Modeling and Computer Simulation, 23(1):

6:1–6:??, January 2013. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Steiniger:2016:ICV

[SU16] Alexander Steiniger and Adelinde M. Uhrmacher. Intensional couplings in variable-structure models: an exploration based on Multilevel-DEVS. ACM Transactions on Modeling and Computer Simulation, 26(2): 9:1–9:??, January 2016. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Stiliadis:1997:RHA

[SV97] Dimitrios Stiliadis and Anujan Varma. A reconfigurable hardware approach to network simulation. ACM Transactions on Modeling and Computer Simulation, 7(1):131–156, January 1997. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Srikant:1996:SRL

[SW96] Rayadurgam Srikant and Ward Whitt. Simulation run lengths to estimate blocking probabilities. ACM Transactions on Modeling and Computer Simulation, 6(1):7–52, January 1996. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Setayeshgar:2013:EIS

[SW13] Leila Setayeshgar and Hui Wang. Efficient importance sampling schemes for a feedforward network. ACM Transactions on Modeling and Computer Simulation, 23(4):21:1– 21:??, October 2013. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Sanchez:2009:TPS

[SWL09] Susan M. Sanchez, Hong Wan, and Thomas W. Lucas. Twophase screening procedure for simulation experiments. ACM Transactions on Modeling and Computer Simulation, 19(2): 7:1–7:??, March 2009. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Sellami:1995:PSM

[SY95] Hatem Sellami and Sudhakar Yalamanchili. Parallelism in sequential multiprocessor simulation models: a case study. ACM Transactions on Modeling and Computer Simulation, 5(2):101–128, April 1995. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Stadlober:1999:PRT

[SZ99] Ernst Stadlober and Heinz Zechner. The patchwork rejection technique for sampling from unimodal distributions. ACM Transactions on Modeling and Computer Simulation, 9(1):59–80, January 1999. CODEN ATMCEZ.

ISSN 1049-3301 (print), 1558-1195 (electronic).

Topcu:2008:MFA

[TAO08] Okan Topçu, Mehmet Adak, and Halit Oğuztüzün. A metamodel for federation architectures. ACM Transactions on Modeling and Computer Simulation, 18(3):10:1–10:??, July 2008. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Tofts:1998:DSP

[TB98] Chris Tofts and Graham Birtwistle. A denotational semantics for a process-based simulation language. ACM Transactions on Modeling and Computer Simulation, 8(3): 281–305, July 1998. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Trunfio:2011:NAS

[TDR^{+}11] Giuseppe A. Trunfio, Donato D’Ambrosio, Rocco Rongo, William Spataro, and Salvatore Di Gregorio. A new algorithm for simulating wildfire spread through cellular automata. ACM Transactions on Modeling and Computer Simulation, 22(1):6:1–6:??, December 2011. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Tezuka:1993:PAA

[Tez93] Shu Tezuka. Polynomial arithmetic analogue of Halton sequences. ACM Transactions on Modeling and Computer Simulation, 3(2):99–107, April 1993. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Talby:2007:CPA

[TFR07] David Talby, Dror G. Feitelson, and Adi Raveh. A Co-Plot analysis of logs and models of parallel workloads. ACM Transactions on Modeling and Computer Simulation, 17(3): 12:1–12:??, July 2007. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Tang:2005:LQP

[TGT05] Wai Teng Tang, Rick Siow Mong Goh, and Ian Li-Jin Thng. Ladder queue: an O(1) priority queue structure for largescale discrete event simulation. ACM Transactions on Modeling and Computer Simulation, 15(3):175–204, July 2005. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Tsompanas:2016:MCM

[TKS16] Michail-Antisthenis I. Tsompanas, Christoforos Kachris, and Georgios Ch. Sirakoulis. Modeling cache memory utilization on multicore using common pool resource game on cellular automata. ACM Transactions on Modeling and Computer Simulation, 26(3): 21:1–21:??, February 2016.

CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Tezuka:1991:EPC

[TL91] Shu Tezuka and Pierre L’Ecuyer. Efficient and portable combined Tausworthe random number generators. ACM Transactions on Modeling and Computer Simulation, 1(2): 99–112, April 1991. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

TerBeek:2018:GES

[TL18] Maurice H. Ter Beek and Michele Loreti. Guest editorial for the special issue on FORmal methods for the quantitative evaluation of collective adaptive Sys-Tems (FORECAST). ACM Transactions on Modeling and Computer Simulation, 28(2): 8:1–8:??, April 2018. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Tezuka:1993:LSA

[TLC93] Shu Tezuka, Pierre L’Ecuyer, and Raymond Couture. On the lattice structure of the add-with-carry and subtractwith-borrow random number generators. ACM Transactions on Modeling and Computer Simulation, 3(4):315– 331, October 1993. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). See remark in [EH95, page 248], and [MZ91] for the original work analyzed in this paper.

Taylor:2008:GEI

[TR08] Simon J. E. Taylor and George Riley. Guest editors’ introduction to special issue on successes in modeling and simulation methodologies. ACM Transactions on Modeling and Computer Simulation, 18(4): 13:1–13:??, September 2008. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Tafazzoli:2009:PCE

[TRK^{+}09] Ali Tafazzoli, Stephen Roberts, Robert Klein, Reid Ness, and Robert Dittus. Probabilistic cost-effectiveness comparison of screening strategies for colorectal cancer. ACM Transactions on Modeling and Computer Simulation, 19(2): 6:1–6:??, March 2009. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Tezuka:1994:NPA

[TT94] Shu Tezuka and Takeshi Tokuyama. A note on polynomial arithmetic analogue of Halton sequences. ACM Transactions on Modeling and Computer Simulation, 4(3): 279–284, July 1994. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Taylor:2012:BGS

[TTSM12] Simon J. E. Taylor, Stephen J.

Turner, Steffen Strassburger, and Navonil Mustafee. Bridging the gap: a standardsbased approach to OR/MS distributed simulation. ACM Transactions on Modeling and Computer Simulation, 22(4): 18:1–18:??, November 2012. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Tuffin:1997:VRA

[Tuf97] Bruno Tuffin. Variance reduction applied to product form multiclass queuing networks. ACM Transactions on Modeling and Computer Simulation, 7(4):478–500, October 1997. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Turkyilmazoglu:2017:PAD

[Tur17] Mustafa Turkyilmazoglu. Parametrized Adomian decomposition method with optimum convergence. ACM Transactions on Modeling and Computer Simulation, 27(4):21:1–21:??, December 2017. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Tuzhilin:1995:ETL

[Tuz95] Alexander Tuzhilin. Extending temporal logic to support high-level simulations. ACM Transactions on Modeling and Computer Simulation, 5(2):

129–155, April 1995. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Ungredda:2024:BOC

[UB24] Juan Ungredda and Juergen Branke. Bayesian optimisation for constrained problems. ACM Transactions on Modeling and Computer Simulation, 34(2):9:1–9:??, April 2024. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3641544.

Uhrmacher:2001:DSM

[Uhr01] A. M. Uhrmacher. Dynamic structures in modeling and simulation: a reflective approach. ACM Transactions on Modeling and Computer Simulation, 11(2):206–232, April 2001. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Uhlig:1997:TDM

[UNMS97] Richard Uhlig, David Nagle, Trevor Mudge, and Stuart Sechrest. Trap-driven memory simulation with Tapeworm II. ACM Transactions on Modeling and Computer Simulation, 7(1):7–41, January 1997. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Ungredda:2022:BOV

[UPB22] Juan Ungredda, Michael Pearce, and Juergen Branke.

Bayesian optimisation vs. input uncertainty reduction. ACM Transactions on Modeling and Computer Simulation, 32(3):17:1–17:26, July 2022. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3510380.

Unger:2000:PSM

[UXC^{+}00] Brian Unger, Zhonge Xiao, John Cleary, Jya-Jang Tsai, and Carey Williamson. Parallel shared-memory simulator performance for large ATM networks. ACM Transactions on Modeling and Computer Simulation, 10(4):358– 391, October 2000. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Vazquez-abad:2002:EBP

[VaAE02] Felisa J. Vázquez-abad, Lachlan L. H. Andrew, and David Everitt. Estimation of blocking probabilities in cellular networks with dynamic channel assignment. ACM Transactions on Modeling and Computer Simulation, 12(1):54–81, January 2002. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Viroli:2018:ERC

[VAB^{+}18] Mirko Viroli, Giorgio Audrito, Jacob Beal, Ferruccio Damiani, and Danilo Pianini. Engineering resilient collective adaptive systems by self-stabilisation. ACM Transactions on Modeling and Computer Simulation, 28(2):16:1– 16:??, April 2018. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Vakili:1992:MPD

[Vak92] Pirooz Vakili. Massively parallel and distributed simulation of a class of discrete event systems: a different perspective. ACM Transactions on Modeling and Computer Simulation, 2(3):214–238, July 1992. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Vandin:2018:RCR

[Van18]  Andrea Vandin.  Repli-

cated Computations Results (RCR) report for “A Holistic Approach for Collaborative Workload Execution in Volunteer Clouds”. ACM Transactions on Modeling and Computer Simulation, 28(2): 15:1–15:??, April 2018. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Vandin:2019:RRA

[Van19] Andrea Vandin. RCR report for analysis of spatiotemporal properties of stochastic systems using TSTL. ACM Transactions on Modeling and Computer Simulation, 29(4): 21:1–21:??, December 2019.

CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3341093.

Villen-Altamirano:2006:ERM

[VAVA06]  Manuel  Villén-Altamirano

and José Villén-Altamirano. On the efficiency of RESTART for multidimensional state systems. ACM Transactions on Modeling and Computer Simulation, 16(3):251–279, July 2006. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

vanBeek:2003:DUD

[vBBR03] D. A. van Beek, V. Bos, and J. E. Rooda. Declaration of unknowns in DAEbased hybrid system specification. ACM Transactions on Modeling and Computer Simulation, 13(1):39–61, January 2003. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Vigna:2016:EEM

[Vig16] Sebastiano Vigna. An experimental exploration of Marsaglia’s xorshift generators, scrambled. ACM Transactions on Mathematical Software, 42(4):30:1– 30:23, July 2016. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://dl.

                                        acm.org/citation.cfm?id=
                                        2845077.

Vissat:2019:AST

[VLN^{+}19] Ludovica Luisa Vissat, Michele Loreti, Laura Nenzi, Jane Hillston, and Glenn Marion. Analysis of spatio-temporal properties of stochastic systems using TSTL. ACM Transactions on Modeling and Computer Simulation, 29(4):20:1–20:??, December 2019. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Vorobeychik:2010:PAS

[Vor10] Yevgeniy Vorobeychik. Probabilistic analysis of simulationbased games. ACM Transactions on Modeling and Computer Simulation, 20(3):16:1– 16:??, September 2010. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Velho:2013:VFL

[VSCL13] Pedro Velho, Lucas Mello Schnorr, Henri Casanova, and Arnaud Legrand. On the validity of flow-level TCP network models for grid and cloud simulations. ACM Transactions on Modeling and Computer Simulation, 23(4):23:1– 23:??, October 2013. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Vieira:2014:RMH

[VSS^{+}14] Hélcio Vieira, Jr., Susan M.

Sanchez, Paul J. Sanchez, Karl Heinz Kienitz, and Mischel Carmen Neyra Belderrain. A restricted multinomial hybrid selection procedure. ACM Transactions on Modeling and Computer Simulation, 24(2):10:1–10:??, February 2014. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

VanMierlo:2020:EEM

[VVB^{+}20] Simon Van Mierlo, Hans Vangheluwe, Simon Breslav, Rhys Goldstein, and Azam Khan. Extending explicitly modelled simulation debugging environments with dynamic structure. ACM Transactions on Modeling and Computer Simulation, 30(1):3:1– 3:25, February 2020. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3338530.

Vicino:2022:UDE

[VWD22] Damian Vicino, Gabriel A. Wainer, and Olivier Dalle. Uncertainty on discrete-event system simulation. ACM Transactions on Modeling and Computer Simulation, 32 (1):2:1–2:27, January 2022. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3466169.

Vu:2022:EPT

[VXE^{+}22] Minh Vu, Lisong Xu, Sebastian Elbaum, Wei Sun, and Kevin Qiao. Efficient protocol testing under temporal uncertain event using discreteevent network simulations. ACM Transactions on Modeling and Computer Simulation, 32(2):13:1–13:30, April 2022. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3490028.

Wang:2015:AAL

[WAGP15] Jingjing Wang, Nael Abu-Ghazaleh, and Dmitry Ponomarev. AIR: Applicationlevel interference resilience for PDES on multicore systems. ACM Transactions on Modeling and Computer Simulation, 25(3):19:1–19:??, April 2015. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Wainer:2015:EPA

[Wai15] Gabriel A. Wainer. Editorial for principles of advanced discrete simulation. ACM Transactions on Modeling and Computer Simulation, 26(1):1:1– 1:??, December 2015. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Wang:2019:VIB

[WCCY19] Wenjing Wang, Nan Chen, Xi Chen, and Linchang Yang. A variational inference-based heteroscedastic Gaussian process approach for simulation metamodeling. ACM Transactions on Modeling and Computer Simulation, 29(1):6:1– 6:??, February 2019. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Wang:2023:EMS

[WCF23] Runan Wang, Giuliano Casale, and Antonio Filieri. Estimating multiclass service demand distributions using Markovian arrival processes. ACM Transactions on Modeling and Computer Simulation, 33(1–2):2:1–2:??, April 2023. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3570924.

Wu:2019:EPS

[WCL^{+}19] Yulin Wu, Wentong Cai, Zengxiang Li, Wen Jun Tan, and Xiangting Hou. Efficient parallel simulation over largescale social contact networks. ACM Transactions on Modeling and Computer Simulation, 29(2):10:1–10:??, April 2019. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3265749.

Wu:2010:TFI

[WCLG10] Dalei Wu, Song Ci, Haiyan Luo, and Hai-Feng Guo. A theoretical framework for interaction measure and sensitivity analysis in cross-layer design. ACM Transactions on Modeling and Computer Simulation, 21(1):6:1–6:??, December 2010. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Wang:2016:BAP

[WCS16] Weikun Wang, Giuliano Casale, and Charles Sutton. A Bayesian approach to parameter inference in queueing networks. ACM Transactions on Modeling and Computer Simulation, 27(1):2:1–2:??, November 2016. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Wadman:2016:LDB

[WCZ16] Wander S. Wadman, Daan T.

Crommelin, and Bert P. Zwart. A large-deviation-based splitting estimation of power flow reliability. ACM Transactions on Modeling and Computer Simulation, 26(4): 23:1–23:??, May 2016. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Wang:2016:FEN

[WDYR16] Jun Wang, Zhenjiang Dong, Sudhakar Yalamanchili, and George Riley. FNM: an enhanced null-message algorithm for parallel simulation of multicore systems. ACM Transactions on Modeling and Computer Simulation, 26(2): 11:1–11:??, January 2016. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Waeber:2012:FSS

[WFH12] Rolf Waeber, Peter I. Frazier, and Shane G. Henderson. A framework for selecting a selection procedure. ACM Transactions on Modeling and Computer Simulation, 22(3): 16:1–16:??, August 2012. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Wu:2004:EAB

[WG04] Yujing Wu and Weibo Gong. Error analysis of burst level modeling of active-idle sources. ACM Transactions on Modeling and Computer Simulation, 14(3):278–304, July 2004. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Wang:2016:MSE

[WG16] Rob J. Wang and Peter W. Glynn. On the marginal standard error rule and the testing of initial transient deletion methods. ACM Transactions on Modeling and Computer Simulation, 27(1):1:1– 1:??, November 2016. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Wang:2020:ERP

[WhN20] Songhao Wang and Szu hui Ng. Enhancing response predictions with a joint Gaussian process model for stochastic simulation models. ACM Transactions on Modeling and Computer Simulation, 30(1): 4:1–4:25, February 2020. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3364219.

Wilson:2007:EIS

[Wil07] James R. Wilson. Editor’s introduction: Special issue honoring Perwez Shahabuddin. ACM Transactions on Modeling and Computer Simulation, 17(2):??, April 2007. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Wu:2022:RCR

[WJ22] Xiaoliang Wu and Dong Jin.

Replicated computational results (RCR) report for “A New Test for Hamming-Weight Dependencies”. ACM Transactions on Modeling and Computer Simulation, 32(3):20:1–20:3, July 2022. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3527583.

Wu:2024:PSQ

[WKC^{+}24] Xiaoliang Wu, Alexander Kolar, Joaquin Chung, Dong Jin, Martin Suchara, and Rajkumar Kettimuthu. Parallel simulation of quantum networks with distributed quantum state management. ACM Transactions on Modeling and Computer Simulation, 34(2):11:1–11:??, April 2024. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/

3634701. See reproducibility report [Pic24].

Wegenkittl:1999:GRC

[WM99] Stefan Wegenkittl and Makoto Matsumoto. Getting rid of correlations among pseudorandom numbers: discarding versus tempering. ACM Transactions on Modeling and Computer Simulation, 9(3): 282–294, July 1999. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Wolfe:2018:MLS

[WMC^{+}18] Noah Wolfe, Misbah Mubarak, Christopher D. Carothers, Robert B. Ross, and Philip H. Carns. Modeling large-scale slim fly networks using parallel discrete-event simulation. ACM Transactions on Modeling and Computer Simulation, 28(4):29:1–29:??, October 2018. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Warren:2004:NSE

[WNFM04] Gary Warren, Ronald Nolte, Ken Funk, and Brian Merrell. Network simulation enhancing network management in realtime. ACM Transactions on Modeling and Computer Simulation, 14(2):196–210, April 2004. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Ware:1998:AMF

[WPN98] Peter P. Ware, Thomas W.

Page, Jr., and Barry L. Nelson. Automatic modeling of file system workloads using two-level arrival processes. ACM Transactions on Modeling and Computer Simulation, 8(3):305–330, July 1998. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Wang:2013:IOS

[WPS13] Honggang Wang, Raghu Pasupathy, and Bruce W. Schmeiser. Integer-ordered simulation optimization using R-SPLINE: Retrospective search with piecewise-linear interpolation and neighborhood enumeration. ACM Transactions on Modeling and Computer Simulation, 23(3): 17:1–17:??, July 2013. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Wu:2009:OSI

[WPW09] Tongqiang Tony Wu, Warren B. Powell, and Alan Whisman. The optimizingsimulator: an illustration using the military airlift problem. ACM Transactions on Modeling and Computer Simulation, 19(3):14:1–14:??, June 2009. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Walsh:2004:SSG

[WS04] Kevin Walsh and Emin Gün Sirer. Staged simulation: a general technique for improving simulation scale and performance. ACM Transactions on Modeling and Computer Simulation, 14(2):170–195, April 2004. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Wu:2001:RNG

[Wu01] Pei-Chi Wu. Random number generation with primitive pentanomials. ACM Transactions on Modeling and Computer Simulation, 11(4):346– 351, October 2001. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Wagner:1995:GIS

[WW95] Mary Ann Flanigan Wagner and James R. Wilson. Graphical interactive simulation input modeling with bivariate Bézier distributions. ACM Transactions on Modeling and Computer Simulation, 5(3): 163–189, July 1995. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Wang:2003:ESQ

[WW03] Chia-Li Wang and Ronald W.

Wolff. Efficient simulation of queues in heavy traffic. ACM Transactions on Modeling and Computer Simulation, 13(1): 62–81, January 2003. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Wunderlich:2006:SSM

[WWFH06] Roland E. Wunderlich, Thomas F.

Wenisch, Babak Falsafi, and James C. Hoe. Statistical sampling of microarchitecture simulation. ACM Transactions on Modeling and Computer Simulation, 16(3):197–224, July 2006. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Wilsdorf:2023:ARA

[WWH^{+}23] Pia Wilsdorf, Anja Wolpers, Jason Hilton, Fiete Haack, and Adelinde M. Uhrmacher. Automatic reuse, adaption, and execution of simulation experiments via provenance patterns. ACM Transactions on Modeling and Computer Simulation, 33(1–2):4:1–4:??, April 2023. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL

https://dl.acm.org/doi/
10.1145/3564928.

Wang:2023:TDC

[WXC^{+}23] Ruihang Wang, Deneng Xia, Zhiwei Cao, Yonggang Wen, Rui Tan, and Xin Zhou. Toward data center digital twins via knowledge-based model calibration and reduction. ACM Transactions on Modeling and Computer Simulation, 33(4):11:1–11:??, October 2023. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL

https://dl.acm.org/doi/
10.1145/3604283.

Wenjie:2020:APW

[WYT^{+}20] Tang Wenjie, Yao Yiping, Li Tianlin, Song Xiao, and Zhu Feng. An adaptive persistence and work-stealing combined algorithm for load balancing on parallel discrete event simulation. ACM Transactions on Modeling and Computer Simulation, 30(2):12:1– 12:26, April 2020. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3364218.

Wang:2015:CES

[WZ15] Hui Wang and Xiang Zhou. A cross-entropy scheme for mixtures. ACM Transactions on Modeling and Computer Simulation, 25(1):6:1–6:??, January 2015. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Wu:2022:SQK

[WZCJ22] Xiaoliang Wu, Bo Zhang, Gong Chen, and Dong Jin.

A scalable quantum key distribution network testbed using parallel discrete-event simulation. ACM Transactions on Modeling and Computer Simulation, 32(2):11:1–11:22, April 2022. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL

https://dl.acm.org/doi/
10.1145/3490029.

Xu:2017:RSP

[XCA^{+}17] Yadong Xu, Wentong Cai, Heiko Aydt, Michael Lees, and Daniel Zehe. Relaxing synchronization in parallel agentbased road traffic simulation. ACM Transactions on Modeling and Computer Simulation, 27(2):14:1–14:??, July 2017. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Xue:2012:DAU

[XGH12] Haidong Xue, Feng Gu, and Xiaolin Hu. Data assimilation using sequential Monte Carlo methods in wildfire spread simulation. ACM Transactions on Modeling and Computer Simulation, 22(4):23:1– 23:??, November 2012. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Xie:2017:FBB

[XLZ17] Wei Xie, Cheng Li, and Pu Zhang. A factor-based Bayesian framework for risk analysis in stochastic simulations. ACM Transactions on Modeling and Computer Simulation, 27(4):27:1–27:??, December 2017. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Xie:2016:MIU

[XNB16] Wei Xie, Barry L. Nelson, and Russell R. Barton. Multivariate input uncertainty in output analysis for stochastic simulation. ACM Transactions on Modeling and Computer Simulation, 27(1):5:1–5:??, November 2016. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Xu:2010:ISC

[XNH10] Jie Xu, Barry L. Nelson, and Jeff L. Hong. Industrial strength COMPASS: a comprehensive algorithm and software for optimization via simulation. ACM Transactions on Modeling and Computer Simulation, 20(1):3:1– 3:29 + 14 (online appendix), January 2010. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Xu:2014:DRR

[XVN14] Jie Xu, Anand Vidyashankar, and Martin K. Nielsen. Drug resistance or re-emergence? Simulating equine parasites. ACM Transactions on Modeling and Computer Simulation, 24(4):20:1–20:??, August 2014. CODEN ATMCEZ.

ISSN 1049-3301 (print), 1558-1195 (electronic).

Xie:2021:GLM

[XYZ21] Wei Xie, Yuan Yi, and Hua Zheng. Global-local metamodel-assisted stochastic programming via simulation. ACM Transactions on Modeling and Computer Simulation, 31(1):2:1–2:34, February 2021. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3411080.

Xue:2023:PBM

[XZY23] Junxiao Xue, Mingchuang Zhang, and Hui Yin. A personality-based model of emotional contagion and control in crowd queuing simulations. ACM Transactions on Modeling and Computer Simulation, 33(1–2):6:1–6:??, April 2023. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL

https://dl.acm.org/doi/
10.1145/3577589.

Yau:1999:APS

[Yau99] Victor Yau. Automating parallel simulation using parallel time streams. ACM Transactions on Modeling and Computer Simulation, 9(2): 171–201, April 1999. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Yucesan:1996:CIA

[YJ96] Enver Yücesan and Sheldon H. Jacobson. Computational issues for accessibility in discrete event simulation. ACM Transactions on Modeling and Computer Simulation, 6(1):53–75, January 1996. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Yang:2021:BCE

[YKA^{+}21] Ran Yang, David Kent, Daniel W. Apley, Jeremy Staum, and David Ruppert. Bias-corrected estimation of the density of a conditional expectation in nested simulation problems. ACM Transactions on Modeling and Computer Simulation, 31(4): 22:1–22:36, October 2021. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3462201.

Yang:1996:CAV

[YL96] Wei-Ning Yang and Wei-Win Liou. Combining antithetic variates and control variates in simulation experiments. ACM Transactions on Modeling and Computer Simulation, 6(4): 243–260, October 1996. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Yuan:1993:MCB

[YN93] Mingjian Yuan and Barry L.

Nelson. Multiple comparisons with the best for steady-state simulation. ACM Transactions on Modeling and Computer Simulation, 3(1):66–79, January 1993. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Yuan:2015:CVP

[YN15] Jun Yuan and Szu Hui Ng.

Calibration, validation, and prediction in random simulation models: Gaussian process metamodels and a Bayesian integrated solution. ACM Transactions on Modeling and Computer Simulation, 25(3): 18:1–18:??, April 2015. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Yuan:2020:IMS

[YN20] Jun Yuan and Szu Hui Ng.

An integrated method for simultaneous calibration and parameter selection in computer models. ACM Transactions on Modeling and Computer Simulation, 30(1):7:1– 7:23, February 2020. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3364217.

Yoginath:2015:EPD

[YP15] Srikanth B. Yoginath and Kalyan S. Perumalla. Efficient parallel discrete event simulation on cloud/virtual machine platforms. ACM Transactions on Modeling and Computer Simulation, 26(1):5:1– 5:??, December 2015. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Yoginath:2018:SCL

[YP18] Srikanth B. Yoginath and Kalyan S. Perumalla. Scalable cloning on large-scale GPU platforms with application to time-stepped simulations on grids. ACM Transactions on Modeling and Computer Simulation, 28(1):5:1–5:??, January 2018. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Yucesan:1992:SBE

[YS92] Enver Yücesan and Lee Schruben. Structural and behavioral equivalence of simulation models. ACM Transactions on Modeling and Computer Simulation, 2(1):82–103, January 1992. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Yi:2017:EBA

[YX17] Yuan Yi and Wei Xie. An efficient budget allocation approach for quantifying the impact of input uncertainty in stochastic simulation. ACM Transactions on Modeling and Computer Simulation, 27(4): 25:1–25:??, December 2017. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Zhang:2024:KED

[ZBTT24] Nan Zhang, Rami Bahsoon, Nikos Tziritas, and Georgios Theodoropoulos. Knowledge equivalence in digital twins of intelligent systems. ACM Transactions on Modeling and Computer Simulation, 34(1):3:1–3:??, January 2024. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3635306.

Zhang:2018:SAS

[ZC18] Chen Zhang and Nan Chen.

Statistical analysis of simulation output from parallel computing. ACM Transactions on Modeling and Computer Simulation, 28(3):21:1–21:??, August 2018. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Zhou:2010:CMS

[ZCC^{+}10] Suiping Zhou, Dan Chen, Wentong Cai, Linbo Luo, Malcolm Yoke Hean Low, Feng Tian, Victor Su-Han Tay, Darren Wee Sze Ong, and Benjamin D. Hamilton. Crowd modeling and simulation technologies. ACM Transactions on Modeling and Computer Simulation, 20(4):20:1–20:??, October 2010. CODEN ATM-CEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Zhou:2004:TSC

[ZCLT04] Suiping Zhou, Wentong Cai, Bu-Sung Lee, and Stephen J. Turner. Time-space consistency in large-scale distributed virtual environments. ACM Transactions on Modeling and Computer Simulation, 14(1):31–47, January 2004. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Zorn:1994:EMM

[ZG94] Benjamin Zorn and Dirk Grunwald. Evaluating models of memory allocation. ACM Transactions on Modeling and Computer Simulation, 4(1): 107–131, January 1994. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Zimmermann:2019:ISN

[ZH19] Armin Zimmermann and Thomas Hotz. Integrating simulation and numerical analysis in the evaluation of generalized stochastic Petri nets. ACM Transactions on Modeling and Computer Simulation, 29(4): 24:1–24:??, December 2019. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3321518.

Zhang:2006:ACT

[ZIC06] Jianlong Zhang, Petros A. Ioannou, and Anastasios Chassiakos. Automated container transport system between inland port and terminals. ACM Transactions on Modeling and Computer Simulation, 16(2):95–118, April 2006. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Zhou:2004:MIH

[ZJTB04] Junlan Zhou, Zhengrong Ji, Mineo Takai, and Rajive Bagrodia. MAYA: Integrating hybrid network modeling to the physical world. ACM Transactions on Modeling and Computer Simulation, 14(2): 149–169, April 2004. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Zikos:2010:ISD

[ZK10] Stylianos Zikos and Helen D.

Karatza. The impact of service demand variability on resource allocation strategies in a grid system. ACM Transactions on Modeling and Computer Simulation, 20(4):19:1– 19:??, October 2010. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Zhao:2017:TXB

[ZL17] Xueqian Zhao and Zhonghai Lu. A tool for xMASbased modeling and analysis of communication fabrics in Simulink. ACM Transactions on Modeling and Computer Simulation, 27(3):16:1– 16:??, September 2017. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Zhong:2022:DDC

[ZLH^{+}22] Jinghui Zhong, Dongrui Li, Zhixing Huang, Chengyu Lu, and Wentong Cai. Datadriven crowd modeling techniques: a survey. ACM Transactions on Modeling and Computer Simulation, 32 (1):4:1–4:33, January 2022. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3481299.

Zeigler:1991:MBM

[ZLK91] Bernard P. Zeigler, Cheng-Jye Luh, and Tag-Gon Kim. Model base management for multifacetted systems. ACM Transactions on Modeling and Computer Simulation, 1(3): 195–218, July 1991. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Zhu:2020:RQS

[ZLZ20] Helin Zhu, Tianyi Liu, and Enlu Zhou. Risk quantification in stochastic simulation under input uncertainty. ACM Transactions on Modeling and Computer Simulation, 30(1): 1:1–1:24, February 2020. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3329117.

Zhu:2023:LSS

[ZLZ23] Tingyu Zhu, Haoyu Liu, and Zeyu Zheng. Learning to simulate sequentially generated data via neural networks and Wasserstein training. ACM Transactions on Modeling and Computer Simulation, 33(3):9:1–9:??, July 2023. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3583070.

Zeltyn:2011:SBM

[ZMM^{+}11] Sergey Zeltyn, Yariv N.

Marmor, Avishai Mandelbaum, Boaz Carmeli, Ohad Greenshpan, Yossi Mesika, Sergev Wasserkrug, Pnina Vortman, Avraham Shtub, Tirza Lauterman, Dagan Schwartz, Kobi Moskovitch, Sara Tzafrir, and Fuad Basis. Simulation-based models of emergency departments:: Operational, tactical, and strategic staffing. ACM Transactions on Modeling and Computer Simulation, 21(4):24:1– 24:??, August 2011. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Zhang:2017:MMB

[ZS17] Qiong Zhang and Yongjia Song. Moment-matching-based conjugacy approximation for Bayesian ranking and selection. ACM Transactions on Modeling and Computer Simulation, 27(4):26:1– 26:??, December 2017. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Zhao:2018:RDD

[ZZC18] Mingbi Zhao, Jinghui Zhong, and Wentong Cai. A roledependent data-driven approach for high-density crowd behavior modeling. ACM Transactions on Modeling and Computer Simulation, 28(4): 28:1–28:??, October 2018. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).